A Complete Bibliography of Publications in Annual Review in Automatic Programming

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

16 June 2026 Version 1.01

Title word cross-reference

-2 [Dub85]. -3 [Dub85].

1 [Dub85, MMO94].

2 [Ahr85, GLTD88, PF61]. 2122 [And85].

3D [VPM94].

60 [BBG^{+}61, Dij63c, Dij63d, Eva64, HH63, Nau64a, Rut63, Woo63]. 67 [IM72]. 68 [BL73, McG78].

A.I [Rus92]. abstraction [BM85, VHR92]. access [Elz80]. access-rights [Elz80]. accountancy [How92]. Achieving [Puc94]. ACM [Ano60e, Ano60f]. ACM-GAMM [Ano60e, Ano60f]. acquisition [CH85, GK77, KGN94, LN92b, NSY92, NK92, YJ89, vdL94]. actions [KHP92]. activated [AJT94]. active [RKC85]. active-set [RKC85]. ADA [FHK88, PEGR80, DG80, Sav81]. Adaptation [RE85]. Adaptive [HH92a, SFVvL92, Ber92, CA94, HR94, Kar92, LK85, PCY92, Sim85, SV92]. adaptively [Bor85]. Additional [HH88]. address [Boo60]. advanced [HJ94]. advances [Hal81, KM81]. adventures [KSŽ92]. aerospace [SG85]. agents [HR94, Jen92, SRRP^{+}94]. aggregate [Lun85]. agricultural [Rei85a]. ai [BJT^{+}94, HR94, Hoo92, KJ92, SM89, WTI92]. aided [Fle85, GRW88, Ono89, Pop85, SG85, Wel85]. aids [Wel85]. aircraft [Pie85, Shi89]. alarm [Lar92]. Alexip [CBGG92]. algebraic [Ano60e, Ano60f]. ALGOL [Ano63a, Ano63e, BBG^{+}61, BL73, Dij63a, Dij63c, Dij63d, Eva64, HH63, Hig63, McG78, Nau64a, Nau64b, Rut61, Rut63, Woo63, vW63, Bem69]. ALGOL-60 [Dij63c, Dij63d]. ALGOL-procedure [Rut61]. algorithm [BCSM94, Che85a, DG94, Dui85, Kiw85, KMEK85, LK85, MNV85]. algorithmic [BBG^{+}61, Nau64a]. algorithms [Dov85, Dub92, FLMS92, GKKP85, KTF92, Kar92, LN92b, Nev85, RNB92, Sim85]. allocation [Elz80, JKW85, JS85]. alternative [Elz80]. alternatives [Boe85]. AMI [DLP92]. Analysis [SP94, Sav80, Tza85, Ahr85, GCOA94, IM85, KTF92, KMP^{+}94, Lar92, Lun85, NK92, Pop85, Ron60, Sta85b, Sta85c, TAT81, TZ89, TAR92, Vor85, vS92]. analyzing [WFR92]. any [BFH85]. Appendix [Ano60a, Ano60c, Ano60b, Ano60d, Ano60e, Ano60f, BBG^{+}61, Dij63a, Nau64a]. applicability [Fan85]. Application [EhGhSl89, FPCN94, GSGPAT94, GGG85, HvALT92, NLÁR94, Stă85a, VK85, ZEZ89, And85, Ano60c, Ano60b, BFH85, BJJ^{+}92, Cle60, Cor94, FCWM92, GK80, Gre85, How92, KM81, Tie85, VK94, Woł85, dJM85]. Applications [ASM88, HGM92, Ano88d, Arz92, BL89, Cus60, Dum94, Est85, FLMS92, GLR81, LN92a, ML85, Mot94, OR88, Rei85a, SNG94, VBRD94, Wel85]. applicative [ČŠ88]. applied [GP94, McK92]. Applying [Mit88, BJT^{+}94, DLP92]. approach [AA85, AFGvdS80, BJNR77, Boe85, FJ92, Fra92a, GRSA94, HKM77, Har92, HK85, Kap77, LNYz89, LN92a, MMA92, PKS88, Pie85, SS70, Sta85b, Tam85, TKD85, TFMM81, Wit85]. Approximate [Vál85, LN92a]. Approximation [ML85, Che85a]. APT [Ric61]. architecture [AAG^{+}94, Han94, HGH^{+}85, KM94, LM89, Mes94, Mos94, SRRP^{+}94, SNG94, VBRD94]. Architectures [Efs92]. area [Kel85]. arising [GM85]. Arithmetic [Maz61]. arm [WB85]. ARTIE [TW94]. Artificial [JST^{+}92, LS94, Mot92, VGQ94, VÅ89, Efs92, GMWK92, JX92, MRSS94, Mor92, Mot94, RLM94, SKKG94]. aspect [BK92, MBK92]. aspects [Tie85]. Assembly [Fel60, Cou69, LCM94, MS92a]. assessment [BBC^{+}92b, Kie92, SKL85]. assignment [MSL^{+}94, Tos85]. assistance [GRSA94]. assistant [PB94]. assisted [HD85]. assurance [Bas92]. Asynchronous [BL89, VK85]. ATMS [FLMS92]. attempt [CW85]. attitude [SKKG94]. augmentation [PCY92]. Augmented [Rus92]. Author [Ano85b, Ano85a, Ano88a, Ano88b, Ano89a, Ano90a, Ano92a, Ano92b, Ano94a, Ano94b]. authors [Ano85c]. Auto [Meg60, HLH92]. Auto-programming [Meg60]. auto-tuner [HLH92]. Autocode [PF61, Red60, Bro60, BM61, Pay60]. autocoded [Ron60]. automata [Sab77, Sav81]. Automated [Mel88, NIYY81]. Automatic [Ano60d, Cus60, Pir73, PV85, Rob60, Cle60, Gle60, Łuk61, PHW60, Sav80, Tay60, Woo63]. automating [Zha92]. automation [TBR^{+}94]. automaton [Woł85]. autonomous [VPM94]. Average [Vor85]. avoid [KvD92].

B [Ano60c]. back [FKKK81]. balance [HS85]. bargaining [Fan85]. base [Ave92, Eng72, LŚT90, TS90]. based [ARA^{+}92, AH92, Arz92, BCD92, BBG^{+}88, BSMS94, Bro94, CA94, CBGG92, Dek85, DS88, DCA94, EB94, FS94, FQVM88, Fin92, Fra92b, GL80, HLH92, How92, Hüb92, JVB92, JX92, JA94, KTF92, KM94, Lar92, Lei92, LwYz89, LH92, LS92, MMA92, Met64, Nev85, NAP92, NIYY81, NK92, Ono89, PA88, PvWD92, PPE^{+}94, Pir73, Puc94, RW89, SJ92, Son90, SGZ89, SWR85, SNG94, TBR^{+}94, TM85, VPM94, VHR92, WHS77, vS92]. bases [Rod90]. basic [Eng85]. batch [Car85, Per92]. be [Bri85, Rei85b]. beam [HvALT92]. BED [BLA94]. Behaviour [Ahr85]. Belgrade [BBS^{+}88]. best [ABDW94]. between [BM85]. bibliography [Ano60d]. bilinear [DHD85]. biomedical [MSTZ77]. bioprocess [FCWM92, IGM^{+}94, SL92b]. Biotechnology [VD92]. blackboard [Cor94]. blast [JX92]. block [Cas94]. Board [Ano60g, Ano64b, Ano69, NIYY81]. boiler [OHSH92]. Book [Bem69]. Bootstrap [Ano92c]. Boundary [Lio85]. bounds [Che85a]. Box [Pop85]. Box-Jenkins [Pop85]. BOXES [Rus92]. build [Jen92, Wit92]. building [Dah85, SR85]. business [Cli61, Cus60, Kop92].

C [BP94]. CAD [OGF85]. calculation [BFH85, Ili61]. CAM [WI81]. capabilities [Ano92c, PPMM94]. care [BBS^{+}88]. Case [RLM94, BK85, KMP^{+}94, KLM^{+}92, Kuv92, LOKK89]. causal [BTM92, LS92]. cell [MS92a, Vor85]. centers [RTC92]. Centralised [IMF^{+}81]. certification [BBC^{+}92b]. chains [PMH85]. chair [MRL^{+}94]. challenges [SAM85]. changeover [UOH^{+}81]. channels [BFH85]. chaotic [JK85]. characterization [UW85]. charts [Arz92]. checkers [Sam69]. checking [RP92]. chemical [Dov85, RK89]. Chronology [Bem69]. CIM [KLM^{+}92]. circuit [Grz85]. circuit-oriented [Grz85]. civil [Sav81]. class [DHD85, JST^{+}92, Mar64]. Clinical [BBS^{+}88]. clock [RP92]. closed [ACD^{+}85, DI92]. closed-loop [ACD^{+}85]. clustering [YJ89]. Cn [Fan85]. co [OTTK81]. co-operating [OTTK81]. COBOL [Ano61, Cli61, Sam61a, Sam61b]. code [Hal77, OS60]. codesign [Tem94]. coding [Cle60, Łuk61, PHW60]. coefficient [CW85]. coherence [SP94]. cold [JHJ90]. Collaborative [RWMC92]. collisions [KvD92]. collocation [AA85]. column [CSDM94]. Combi [Ger85]. Combined [SM94]. command [Sha63]. commands [MRL^{+}94]. Commercial [BJT^{+}94, DBG63, Åré92, Bae61, BG64, Cli61, Hum60]. committee [Ano60e, Ano60f, Ano77g]. common [KHM81, Ric61]. Communication [NK94, Fan94, Grz85, Kap77, KK88, NY94, Sch80, Ste90]. communities [Sar85]. company [Kuv92, Sch92]. compare [MR94]. Comparison [Bro94, Cli61, Pie85]. compensation [PC92]. compilation [BG73, KSK94]. compiler [BMMR63, Cli61, Eva64, Hum60, Iro63, Met64, Nau64b, Wil64]. Compiling [lCMM77, Wil64]. complementarity [APPZ85, NSY92]. complex [BCSM94, Bor85, Eng85, Hüb92, MSL^{+}94, rnhWS85, Sch85a, Stă85a, Syd85a, VK85]. complexity [Bos77, KZ88]. complicated [Dem85]. computable [Ano60c, Ano60b]. Computation [Mor85, BSF85, MdCMC88, Tza85]. computational [BTM92, Pao94, RKC85]. Computer [Fle85, Sab77, Wel85, And85, BI85, Fed85, Fis85, Fri64, GK77, Grz85, GRW88, HD85, IMF^{+}81, Kel85, MSL^{+}94, NKIT81, NHIK77, Ono89, Par85, Pop85, Ric61, SG85, Ush81, Wal80, WI81]. computer-aided [Ono89, SG85]. computers [HJLT77, KM81]. computing [Dui85, Mes94, Woo63]. concave [SWR85]. Concept [FP94, Eng85, GRW88, HJ73, NK92, Sow81]. concepts [Fan85, IM72]. Conceptual [Mes94]. concerning [Loc85]. concerns [Mit88]. concurrency [BM85, GG90]. Concurrent [Kuz85, And85, KSK94, VS88]. condition [RLM94]. conditional [Sch80]. conditions [EB94]. configuration [KKK92, TBR^{+}94]. CONRAD [KMP^{+}94]. consideration [BK92, LH92]. constrained [AA85, Kiw85, dJM85]. constraint [NAP92, RP92]. constraints [DR94]. Construction [KK88, HK85, PV85]. constructs [BP94]. consulting [BBG^{+}88]. content [Hug92]. contention [GJ85]. Contents [Ano63b, Ano63c, Ano64a]. Continuation [KSK94]. continuous [APPZ85, CBGG92, GRSA94, Spe85, Woo92]. contribution [IM85]. Control [GCOA94, ÅHP89, AA85, ARA^{+}92, AAG^{+}94, AFGvdS80, AH92, BBC^{+}92a, BJ92, Ber92, BSMS94, Bri85, Bro94, Buc85, CA94, Car85, CSDM94, CVB85, CTW94, CH85, CC85, CNV^{+}92a, DISN77, DSD^{+}88, Dem85, DCA94, Dov85, ES77, Efs92, FQ90, FS94, Fin92, Fis85, FJ92, GKKP85, GL80, GCOA92, GLTD88, GGG85, GK80, GRSA94, GK77, Grz85, GM85, Hal81, HPP85, Har89, HGM92, HH92a, HvALT92, HUS92, HGH^{+}85, HH92b, HD85, JVB92, JST^{+}92, JH92, Jen92, JX92, KTF92, Kar92, KHP92, Kra85, KJVB92, LL94, LNYz89, LS94, Lei92, LŚT90, LI92, LA92, LN92a, Lio85, LD85, Lud80, LOKK89, LM89, MS94, MP85, MRL^{+}94, McK92, MS92a, Mel88, MBK92, MAU^{+}81, MMW92, MSTZ77, ML85, Mor92, Mot94, NSY92, NS92, Nev85, NLÁR94, NIYY81, Obe85, Ono89, OGF85]. control [OHSH92, PCY92, Pao92a, Pao92b, PvWD92, PPMM94, Per92, Put85, PEGR80, Rei85a, Rei85b, RNB92, RK89, RPF^{+}92, RW89, Rod90, RHJ^{+}94, RTC92, RK85, Rus85, Rus92, Sab77, SP94, SRRP^{+}94, Sav80, rnhWS85, SKKG94, SG85, Séd90, SL92a, SM89, SGZ89, SL92b, SWR85, SFVvL92, SV92, Tho77, Tie85, Tos85, Ule85, UW85, VM92, VÅ89, VHR92, VK94, WTI92, WDM92, WI81, WB85, Wei85, WHS77, Woł85, XL85, YKI92, YJ89, ZEZ89, Zha92, dJM85, vdL94]. control/monitoring [LŚT90]. Controllability [DHD85]. controlled [Bor85, HGP92, Meg60, Wal80]. controller [AJT94, EhGhSl89, Han94, HJ94, KRK92, Poh85, TBR^{+}94, YLZ89]. controllers [Chr85, EB94, FCWM92, KMO^{+}92, PA88, PBMR94]. Controlling [KZ88, Fri92]. controls [BSF85]. conventional [Bro94]. Convergence [Sim85]. conversational [BG73, Lec73]. conversations [CCMM90]. conversion [Fel60]. convex [Kiw85]. cooperating [Jen92]. cooperative [BK85]. core [Tza85]. Corporation [IMF^{+}81]. correct [KK88]. correction [Ano60c]. Correctness [Szm88, Szl77, SSSS94]. cost [GLTD88, KTF92, KZ88, MS92a, VG94]. costs [LMK77]. coupled [Tza85]. course [Hug92]. CRDB [TS90]. creative [Tem88]. criteria [RE85, SWR85]. criterion [Spi85]. critic [SV92]. critical [Ano61, Sch80]. CSS [Spe85]. CSS-PASCAL [Spe85]. CSTR [ZMM94]. cucumber [SR85]. current [DR94]. cyclical [SFVvL92].

DAI [SRRP^{+}94]. DAI-DEPUR [SRRP^{+}94]. dangerous [BBG^{+}88]. Data [D’I69, vdL94, BBO94, BJ92, BM85, CH85, DG94, Eng72, GK77, Har89, LŚT90, NKIT81, Nev85, Pfe94, Rod90, Séd90, Sjö92, Sow81, TS90]. data-base [Eng72]. data-based [Nev85]. data-driven [Sow81]. database [BZ90, FQ90, FP94, GG90, JHJ90, OR88, Son90, Ste90]. deadlines [VHR92]. debugging [DISN77, TFMM81]. Decentralized [LL94, ACD^{+}85]. decentrally [UW85]. decision [BJJ^{+}92, BK85, Est85, Fan85, Osy85, SGZ89, Sta85b, Syd85a, Tam85, Wel85, Wit85, ZBZ85]. decisions [SS70]. Decomplexing [Bos77]. decomposition [Bak85, Put85, Syd85a]. decoupling [FJ92]. dedicated [Sch77]. deep [MS94]. Definition [McG78, HJ73]. delay [BFH85, LNYz89, Tza85, ZEZ89]. demonstrating [FHK88]. department [BBG^{+}88]. dependent [Ben85a]. deposition [LS94]. depth [VPM94]. DEPUR [SRRP^{+}94]. Derivation [NLÁR94, FQVM88]. Description [Cli61, BM61, LW69, Rus85, Sam61a, Sav80, Woo63]. Design [ASM88, Dij63b, MC92, Smi70, BJNR77, BJJ^{+}92, Bri85, EB94, FQ90, Fle85, GLTD88, GM85, Hoo92, HD85, KTF92, KSŽ92, MEN77, Mes94, Moo90, Nau64b, NLÁR94, Ono89, PA88, PBMR94, Put85, SG85, SSSS94, TM85, Tos85, Ush81]. designed [Bri85]. Designing [SJ92, SNG94, DSD^{+}88, SM89, Tem88, Tho77]. designs [Sch92]. Detailed [Sam61a]. Detection [MČ85, CSF92, DI92, Fra92b, LH92, PPMM94, Shi89, SG92, VM92, vS92]. DEUCE [Den60, Rob60]. developers [HBS92]. Developing [BSMS94]. Development [ACDM85, BBG^{+}88, Kie92, Kop92, KJ92, PF92, ČŠ88, Fri92, GLR81, GMWK92, JL88, KLM^{+}92, Kuv92, LMK77, Mar92, Mit88, PA88, RS92a, RPF^{+}92, Spe85, VGQ94]. device [ES77]. devoted [LTK92]. diagnose [LCM94]. diagnosis [Ave92, BH92, BCD92, CBGG92, GMS94, How92, Hüb92, LS92, NAP92, PC92, ZMM94]. diagnosis-state [NAP92]. diagnostic [PvWD92]. diagnostician [YY92]. diagrams [Cas94]. dialogue [Sch77]. Dice [KJVB92]. Dictionary [OR88]. Dictionary-supported [OR88]. different [Sch80]. differential [Cle60, Mor85]. Digital [Moo90, Dek85, Wos85]. dim [SC94]. Dimensions [Fin92]. directed [HGP92, Iro63]. discrete [DHD85, SM89]. Discussion [Ano77a, Ano77b, Ano77c, Ano77d, Ano77e, Ano77f, Ano61]. distillation [SGZ89]. Distributed [Har89, LM89, NHIK77, SRRP^{+}94, Tos85, ADLD94, Ber92, BJJ^{+}92, BZ90, Bra85, CCMM90, DG94, Dem85, DG80, FQ90, GLR81, Hal81, HJLT77, HGH^{+}85, JH92, JHJ90, KHM81, LL94, LŚT90, ML92, MS94, Moo90, Rod90, RHJ^{+}94, Sch94a, Sch94b, Séd90, Son90, TS90, Tho77, WHS77]. Distribution [Yas85, Rei85a, Sab77, YY92]. documentation [BGS80]. down [Hoo92]. draft [Dah85]. drainage [Kie92]. drinking [Wit92]. driven [Sow81]. dual [Eng85]. Dynamic [GCOA94, HJ73, Bak85, PC92, Pie85, RK85, Woo92]. dynamical [Ben85a, LS92, SAM85]. dynamics [BA92, MdCMC88]. dynamize [CW85].

E.S [BSMS94]. ecology [Stă85a, SKL85]. ecosystems [SKL85]. Editorial [Ano60g, Ano63d, Ano64b, Ano69]. education [Sau92]. Educational [MBK92]. Effects [Kuv92]. efficient [GRW88]. effort [ABDW94]. electricity [Sab77]. Electronic [MRL^{+}94]. elementary [Dah85, KGN94]. Elliott [PF61]. ellipsoid [KMEK85]. ellipsometry [PPE^{+}94]. embedded [Bes92, BLA94, Hoo92]. emergency [VM92]. EMG [HUS92]. empirical [Fra92a]. employing [NKIT81]. empty [ZBZ85]. end [GTW94, JA94]. end-user [JA94]. engine [MC92]. engineering [DSD^{+}88, GRW88, Sav81, TM85]. Enhancing [GMWK92]. Entscheidungsproblem [Ano60c, Ano60b]. Environment [PF92, ADLD94, CNV^{+}92a, JL88, KTF92, Kie92, Moo90, NK94, PvWD92, PFTD85]. environments [Åré92, CNV^{+}92b, Pao92b]. EPOS [BGS80]. equations [Cle60, NAP92]. error [MS92a, PPMM94]. ESPRIT [DISN77]. estimation [Har89, LMK77, MMO94, SG92, Unb85, VPM94, WB92]. estimation/a [Unb85]. estimator [vS92]. Europe [Ano92c]. European [PF92]. Evaluation [KHM81, BGS80, CSF92, KGN94, PKS88]. Events [Kry85, NAP92]. evidence [Sch85a]. Evolution [LOKK89, Voi85]. example [Bri85, FHK88, Ron60]. Examples [DI92]. execution [KGN94, PvWD92]. executive [Gaf88]. executor [Hal77]. exhibiting [LI92]. exothermic [Per92]. expansions [Tza85]. Experience [HW77, Par85, Wei88, CBGG92, Pay60]. Experiences [Bas92, MEN77, Pop85, Koc88, Sau92]. experiment [Wil64]. experimental [IM85, Spi85]. experiments [Sch85c, TW94]. Expert [Shi89, Åré92, BL89, CNV^{+}92b, CNV^{+}92a, EhGhSl89, GSGPAT94, GRSA94, LN92a, LOKK89, MBK92, NK94, RK89, TZ89, WFR92, YLZ89]. Explicit [RP92]. exploration [TM85]. Exponential [PMH85]. expressing [Tay80]. extended [Ono89]. extensible [SY74]. extensive [Koc88]. extraction [RPF^{+}92]. fact [Ste61, Cli61]. Factors [BJJ^{+}92]. failure [CSF92, Shi89]. failures [BFH85, LCM94]. fall [FKKK81]. fall-back [FKKK81]. Fault [VM92, ZMM94, Ave92, BH92, DI92, Fra92b, LH92, PC92, PPMM94, SG92, YY92, vS92]. feasibility [Har92]. features [BGS80]. fee [Kie92]. feedback [BA92, BSF85, DS88, KMEK85, Ono89, VÅ89, Yas85]. fermentation [EhGhSl89, GSGPAT94, GMWK92, JST^{+}92]. fes [SFVvL92]. fes-induced [SFVvL92]. field [Che85b, DSD^{+}88, ES77]. Fieldbus [DR94]. File [Cla63, Ush81, McG69]. films [PPE^{+}94]. finding [LD85]. findings [Ano92c]. fingers [HUS92]. fire [BBG^{+}88]. FIREX [BBG^{+}88]. First [Ano92c, BCSM94]. fixed [ABDW94]. Flanex [WFR92]. flexible [Ave92, HvALT92, Kap77, MS92a, Wei88, Wel85]. flight [McK92, VK94]. flotation [CSDM94]. Flow [Grz85, DG94, RWMC92, WFR92, Tay60]. FLOW-MATIC [Tay60]. FMS [PPMM94]. follow [XL85]. following [MRSS94, Puc94]. force [FS94]. Forecast [SKL85]. forecasting [Pop85, Ron60]. Foreword [BW85, Cre94]. Form [Bem69]. Formal [Ano88c, McG78, Woł85, LW69, NK92]. formation [Voi85]. formula [Cle60]. formulae [Maz61]. Formulation [BGS80]. FORTRAN [Sch85b]. four [Ano60a]. framework [KJVB92, MS92b]. friendly [Fed85, NS92]. front [GTW94]. full [Sau92]. full-time [Sau92]. function [Arz92, MS92a, SM94]. function-oriented [SM94]. functional [Tay80]. functions [MSL^{+}94, NHIK77]. Functor [OTTK81]. funds [JKW85]. furnace [JX92, NLÁR94, WTI92]. Further [Den60, Rus92]. fusion [Har89]. Future [Gle60, Hal81]. Fuzzy [GCOA94, JVB92, WDM92, AJT94, Boe85, Bro94, CSDM94, Dum94, EhGhSl89, EB94, GCOA92, HJ94, KMO^{+}92, LN92b, NLÁR94, Pfe94, SP94, VG94, YLZ89, ZMM94].

GAELIC [lCMM77]. gain [EB94, KMEK85]. game [BK85, Sam69]. games [Mor85]. GAMM [Ano60e, Ano60f]. gas [LH92]. gases [Gre85]. GEC [HJLT77]. General [IM72, Sam61b, BCD92, Lec73, MČ85]. Generalized [McG69, vW63]. Generation [Sch77, MMA92]. Genetic [Dub92, RNB92, Kar92, LN92b]. genie [Ili61]. Geometric [MMA92]. GIER [Nau64b]. GKS [FHK88]. glass [Wei85]. Global [HS85, JS85, Ben85a, lCMM77]. goal [MSL^{+}94]. goals [Sta85b]. goals-a [Sta85b]. goods [BBG^{+}88]. GPSS [Sch85b]. GPSS-FORTRAN [Sch85b]. GRAFCET [LQM94, BDRS94]. graph [Dui85, UW85]. graph-theoretic [Dui85]. graphic [Ben85b]. Graphical [Chr85, Lec73]. graphics [Buc85]. graphs [Rei85b]. grasp [FS94]. grate [Jen92]. greenhouse [SR85]. gripper [KHP92]. group [Fan85]. growth [Bae61, PPE^{+}94, PMH85, SR85]. guidance [VG94]. guided [MRL^{+}94]. guideway [NIYY81].

hand [FS94]. hard [DG94, Gaf88, Mes94, Mot92]. hard-real-time [DG94]. Hardware [RK89, Tem94]. HDDBMS [ZZY90]. health [BBS^{+}88]. heuristic [MNV85]. Hierarchical [Rei85a, SSSS94, Buc85]. High [DISN77, BP94, GL80, Kra85, Puc94, TS90, WHS77]. high-level [BP94]. higher [OTTK81]. higher-level [OTTK81]. highly [VS88]. hint [AAG^{+}94, SNG94]. hints [MNV85]. History [Bem69]. homogenization [Lio85]. Human [JA94, HH92b, NS92]. human-friendly [NS92]. Hybrid [BDRS94, NLÁR94, WB92]. hyperbolic [Lio85].

I/O [KPR^{+}80]. IBM [Fri64, Han60, Hum60]. ideal [Fri64]. Identification [Pfe94, Ben85a, GP94, HS85, JK85, LK85, Unb85]. II [BW61, Sam69, Smi70]. image [LMS94]. impacts [SKL85]. implement [CCMM90]. Implementation [HLH92, NS92, Sch80, GLTD88, GG90, HH88, Hal77, Kuv92, MEN77, Sab77, VS88]. implementations [Mos94]. implemented [Han94]. Implementing [CC85, Har92]. improved [IGM^{+}94]. improvement [MS92b]. improving [SOJ92]. Incidence [BZ90]. Incorporating [KSŽ92]. Increasing [PPMM94]. Incremental [BG73, KHP92]. Independent [Dij63b, Cou69]. index [Ano85b, Ano85a, Ano85d, Ano88a, Ano88b, Ano89a, Ano89b, Ano90a, Ano90b, Ano92a, Ano92b, Ano92d, Ano92e, Ano94a, Ano94b, Ano94c]. indoor [VPM94]. induced [SFVvL92]. inducing [KMO^{+}92]. Induction [HH92b]. industrial [CR92, GLTD88, Jen92, Kuv92, KM81, LwYz89, LA92, MAU^{+}81]. industries [HPP85]. inexact [APPZ85]. inference [JVB92]. inferencing [Har89]. influence [BJ92]. Influences [HBS92]. influencing [BJJ^{+}92]. influential [MČ85]. informatics [Sch85c]. information [BBS^{+}88, Dab̧85, How92, Kop92, PKS88, PS85]. infrared [MRL^{+}94]. ingredient [Wei85]. inherently [HH88]. inheritance [KKK92]. inputs [SV92]. insensitive [SG85]. inspection [CR92]. instrument [vS92]. Integrated [JL88, KGN94, Bri85, PKS88, RWMC92, RHJ^{+}94]. Integrating [ABDW94, HR94, ZZY90]. Integration [Ben85b, MS92a, AAG^{+}94, FHK88, IGM^{+}94, Kun85, WHS77]. Intelligence [VGQ94, Efs92, JST^{+}92, JX92, LS94, ML92, Mot92, Mot94, Pao94, VÅ89]. Intelligent [FCWM92, KHP92, ÅHP89, AAG^{+}94, Ave92, HR94, HGM92, HH92a, JH92, KGN94, KJVB92, LL94, LTK92, LM89, MS94, MC92, NS92, Ono89, OHSH92, Pao92b, RPF^{+}92, ZEZ89]. Interaction [TM85, HS85, JA94]. Interactive [Fed85, JKW85, MSTZ77, AFGvdS80, Osy85, PFTD85, Spi85, SWR85, Wit85]. interconnected [Eng85, HS85, Lun85]. interconnection [BK92]. interdependent [Sta85b]. interface [Buc85, Fed85, LTK92, RTC92]. interfaces [SNG94]. Interference [Rut61]. internal [KRK92]. International [Ano77g, Ano60e, Ano60f]. interpretation [BG73, OS60]. interpretative [Den60]. interpreted [AK90]. interpretive [AFGvdS80, Fel60, WHS77]. Intertask [Ste90]. Intertask-communication [Ste90]. introduced [Sow81]. Introducing [Sch92, Bas92]. Introduction [McG78, BL73]. invariant [Dui85]. invariants [GK80]. inverse [MdCMC88]. investigated [Rei85b]. investment [CW85]. involved [LD85]. IPAS [HBS92]. IPS [PFTD85]. iPSC [And85]. isolation [CSF92]. issues [Dum94, Rod90, Sch85a, SAM85].

Japan [KM81]. Jenkins [Pop85]. job [Lei92]. jobs [Yas85]. JOSS [Smi70]. JOVIAL [Mar64, Sha63]. Jump [LI92].

Kernel [Voj77, BLA94, GLR81, Gór80]. Keyword [Ano89b, Ano90b, Ano92d, Ano92e, Ano94c]. Kimitsu [IMF^{+}81]. Knowledge [ARA^{+}92, AH92, CR92, How92, IGM^{+}94, LTK92, Lei92, LH92, RW89, VHR92, Arz92, BBG^{+}88, Fin92, HLH92, JA94, KTF92, KM94, NSY92, NK92, PPMM94, SJ92, YJ89]. Knowledge-based [ARA^{+}92, AH92, How92, RW89, VHR92, Arz92, BBG^{+}88, Fin92, HLH92, KTF92]. Kodak [BL85]. KRTM [Gór80].

ladder [GMS94]. landing [Pie85]. landing-approach [Pie85]. language [Ano60e, Ano60f, BBG^{+}61, Bae61, BM61, Car85, Cou69, DISN77, DG80, Dub85, FP94, GL80, GTW94, HH88, HW77, IM72, Kry85, LMS94, LQM94, Lud81, Nau64a, Ono89, PS77, Ric61, RTC92, Sha63, Spe85, Wal80, Wil64, lCMM77]. Languages [Dij63b, Ano88d, Ben85b, Bri85, Chr85, DBG63, HJ73, SY74]. large [Dui85, HGH^{+}85, Kuv92, Mar92, MHS^{+}88, OGF85, Stă85a, Tie85, XL85]. laser [LS94]. learned [MHS^{+}88]. Learning [CVB85, GP94, KvD92, LCM94, RBSG94, Ber92, Dub92, FS94, Fin92, KHP92, Pfe94, Sam69, YY92, YLZ89]. Least [BCSM94, Sim85]. Least-space-time-first [BCSM94]. leg [SFVvL92]. Leontieff [CW85]. less [Séd90]. lessons [MHS^{+}88]. level [BP94, DISN77, GL80, Kuv92, OTTK81, WHS77]. library [BP94, Bro60]. lights [HJ94]. limited [Boe85]. line [CR92, CC85, PC92, Rei85a, Rus85, UOH^{+}81, Wei85, WFR92]. linear [Ben85a, BJ92, Dov85, GKKP85, Gut85, Rei85b, TAR92, Tos85]. linearly [Kiw85]. linguistic [Sta85b]. linguistics [Met64]. link [BM85, WB85]. liquids [Gre85]. List [Ano60a, Ano77h, Ano80, Ano85c, Wil64]. list-processing [Wil64]. load [NHIK77, XL85]. local [Ben85a, JS85, Kel85, VG94]. Log [Bem69]. Logic [LS92, CSF92, Dum94, GMS94, KSK94, RP92, VG94, ZZY90]. Logic-based [LS92]. Logical [PBMR94]. logically [KK88]. logics [RKC85]. loop [ACD^{+}85, DS88, Mor85]. loops [DI92]. loss [BFH85]. LOTOS [FQVM88]. Low [KTF92, GLTD88, VG94]. Low-cost [KTF92, VG94]. lower [SFVvL92]. LTPL [KPR^{+}80]. LTPL-E [KPR^{+}80].

Machine [Dij63b, YY92, Cou69, Meg60, RBSG94, Sam69, UOH^{+}81, WI81]. machine-independent [Cou69]. macro [Bro69]. MADCAP [BW61]. magnetic [ES77]. Making [Ano63e, Dij63d, BK85, Est85, Fan85, Osy85, Sta85b, Tam85, Wit85, ZBZ85]. management [Ano92c, BJJ^{+}92, DS88, Fed85, Fri92, HK92, HBS92, Hug92, JL88, KKK92, Koc88, LOKK89, MS92a, PB94, Sau92, Wei88]. Managing [DLP92]. manipulator [MdCMC88, MSTZ77]. manufacturing [DCA94, NK94, Wel85]. map [VBRD94]. Mapping [BP94, ES77]. marine [GP94]. Mark [PF61, PHW60]. Markovian [PV85, vS92]. Maschine [Sta85c]. master [Bos77]. materials [How92]. MATH [Tay60]. MATH-MATIC [Tay60]. Mathematical [Sar85, SKL85, AA85, Ger85, Loc85]. MATIC [Tay60]. matrix [CW85]. maturity [Kuv92]. means [Bor85, Elz80, Kuz85, Lun85, Rei85b]. measure [Vor85]. measuring [Ano92c, Wos85]. mechanical [Met64]. mechanism [Ber92, VS88]. medicine [Wit92]. memory [GJ85]. MERCURY [Bro60, BM61]. metallurgical [MP85]. Metaparadigm [KSŽ92]. method [APPZ85, BFH85, FKKK81, HS85, KMO^{+}92, KMEK85, SG92, Tay80, UW85, VG94, YJ89]. Methodical [Koc88]. methodological [BJNR77]. methodologies [MMW92]. Methodology [Dek85, AAG^{+}94, BCD92, FQVM88, FPCN94, Kal85, NS92, Rus92]. Methods [ČŠ88, BL89, Eng85, Est85, GGG85, HBS92, Hin94, LQM94, LH92, MTK^{+}94, SAM85, Unb85, Wei88]. MFM [Lar92]. Microcomputer [NIYY81]. Microprocessor [PA88, Sow81, WHS77]. microprocessors [Kun85]. Microsoft [Fri92]. mill [JHJ90]. mind [Moo90]. mine [MS94]. mineral [MP85]. minimax [Obe85]. minimization [Kiw85]. Minimum [WB85]. Minimum-time [WB85]. MIP [ARA^{+}92]. mix [ZEZ89]. mixed [YKI92]. MMOSS [Fan94]. mobile [KvD92, MMO94, VG94, VPM94]. mode [HvALT92]. mode-switch [HvALT92]. Model [CBGG92, Hüb92, Lar92, NAP92, SR85, BCD92, BTM92, Dah85, FQ90, Fra92b, Gaf88, GJ85, HK92, HK85, KPR^{+}80, KRK92, LH92, OTTK81, PvWD92, Sch85c, Woł85, ZBZ85]. Model-based [CBGG92, Hüb92, Lar92, NAP92, Fra92b, PvWD92]. modeling [BI85, Loc85, Par85]. Modelling [AK90, Dab̧85, LNYz89, BDRS94, Ger85, Gre85, Kal85, LI92, MMW92, NS92, RLM94, Stă85a, Tie85, TAR92, Wos85]. Models [TKD85, Ave92, Ben85a, CW85, Lun85, NLÁR94, PMH85, Pie85, PV85, Sar85, SOJ92, Sch85c]. modern [Wei88]. modified [Kuz85]. MODULA [Ahr85, HW77, GLTD88]. MODULA-2 [Ahr85, GLTD88]. Modulas [Dub85]. Modulas-1 [Dub85]. moisture [ZEZ89]. Monitoring [BH92, Mor92, ČŠ88, Cor94, LTK92, LŚT90, LM89, Pao94, PPE^{+}94, RPF^{+}92, RLM94]. monitors [RBSG94]. MORSAF [PB94]. motor [LOKK89]. movable [Woł85]. movements [SFVvL92]. multi [BCD92, HKM77, Har89, HH63, Mar92, NKIT81, NHIK77, RE85, Sow81, Spi85, SWR85]. multi-computer [NHIK77]. multi-criteria [RE85, SWR85]. multi-criterion [Spi85]. multi-microprocessor [Sow81]. multi-model-based [BCD92]. multi-national [Mar92]. multi-pass [HH63]. multi-processor [HKM77, NKIT81]. multi-sensor [Har89]. multicriteria [Est85]. Multicriterial [ZBZ85, BK85]. multicriterion [Osy85]. multidimensional [MS92b]. multifingered [KHP92]. multilevel [Wal80]. multimedia [Fan94, NY94]. Multiobjective [SG85, Tam85, Fle85, Gut85, Kiw85, Wit85]. multiphase [SSSS94].

Multiple [MSL^{+}94, BCSM94, Obe85]. Multiple-goal [MSL^{+}94]. Multiprocessor [TAT81, GJ85, VK85]. multiprograms [Gór80]. multirecursive [Sim85]. multithreaded [BLA94]. Multivalued [CSF92]. myth [Ste61].

N [SC94]. N-dim [SC94]. NARMAX [FJ92]. national [Mar92]. Natural [GTW94, RTC92, Gre85, SKL85]. Navigation [MMO94, KvD92, VPM94]. navigator [Zha92]. NECTAR [VK94]. NECTAR-project [VK94]. needed [Mot92]. negotiations [JKW85]. net [GK77, PCY92, Pao92a, PPE^{+}94]. nets [AK90, Ger85, Kuz85, Sta85c, Szl77]. network [ACDM85, AFGvdS80, Ave92, CA94, DCA94, Grz85, KRK92, LwYz89, LI92, MRSS94, Mor92, Sjö92, SFVvL92, YKI92, YY92]. networks [ACD^{+}85, BJ92, CTW94, GMWK92, HUS92, Kel85, McK92, MMW92, Par85, RLM94, SKKG94, VK94, ZMM94]. Netz [Sta85c]. Neural [Ave92, CA94, CTW94, FS94, LwYz89, LI92, McK92, MMW92, PCY92, Pao92a, PPE^{+}94, BJ92, DCA94, GMWK92, Han94, HUS92, KRK92, MRSS94, Mor92, PBMR94, RLM94, SKKG94, Sjö92, SFVvL92, VK94, YKI92, ZMM94]. Neural-based [FS94]. Neural-net [PCY92, PPE^{+}94]. neurally [HGP92]. neurofuzzy [MBH94]. Nippon [IMF^{+}81]. No [NKIT81]. non [Ben85a, BJ92, Dov85, SWR85, TFMM81]. non-linear [Ben85a, BJ92]. nondominated [Boe85]. nonlinear [BSF85, FJ92, GKKP85, GCOA92, GGG85, HPP85, LNYz89, LK85, LD85, MČ85, RKC85, Zha92]. nonsmooth [Kiw85]. note [Ano63d, Tem94]. nuclear [Tza85, XL85]. number [BFH85, Boe85, Dui85, Yas85]. numbers [Ano60c, Ano60b]. Numerical [Kun85, LD85, Obe85, Ili61, PV85, WI81]. numerically [Meg60].

O [KPR^{+}80]. Object [ASM88, BJJ^{+}92, GSGPAT94, Ono89, Woł85, ZZY90]. object-oriented [GSGPAT94, Ono89]. Objective [Syd85a, MSL^{+}94]. Objective-oriented [Syd85a]. objects [CCMM90, Dem85, LM89]. observations [Woo63]. observer [MMA92]. observer-based [MMA92]. on-board [NIYY81]. On-line [PC92, UOH^{+}81, Wei85, CR92, CC85, Rei85a, Rus85]. one [Ano60c, Ano60b, Wit92]. op [Kuz85]. open [Mor85]. open-loop [Mor85]. Opening [Boo60]. operating [BJNR77, Fan94, GLR81, Han60, MR94, MEN77, OTTK81, Sch77]. Operational [Pay60]. optical [PPE^{+}94]. Optimal [Dem85, GKKP85, Kra85, MP85, Pie85, rnhWS85, SKKG94, XL85, AA85, ACDM85, BSF85, Dov85, Gut85, JS85, LD85, McK92, Nev85, RK85, YKI92, dJM85]. Optimization [GM85, Eng85, Fle85, GGG85, Gut85, KMEK85, MSL^{+}94, Pao94, RE85, SR85, SS70, Spi85, SWR85, SKL85, TM85, Vál85, Voi85, Wit85]. optimization-based [SWR85, TM85]. Optimizing [Bor85]. ORACLE [FHK88]. oral [MRL^{+}94]. ordinary [Cle60]. Organisation [Mar92, BK92]. organization [Eng72, IM85, RS92a]. Oriented [ASM88, GK80, ES77, GSGPAT94, Grz85, Kry85, Ono89, Sch94b, SM94, Syd85a]. orthogonal [AA85]. oscillations [XL85]. output [GCOA94]. over-specification [KK88]. overlapping [Bak85]. own [Wit92].

P [FCWM92]. package [BL85, MAU^{+}81, Rus85]. packet [Grz85]. packet-switched [Grz85]. papers [Dij63a, VR92]. parabolic [Poh85]. paradigm [KvD92, Mot92, RHJ^{+}94, ZZY90]. paradigms [BBC^{+}92a]. Parallel [FJ92, SAM85, Dek85, Hal77, Han94, Kun85, Szl77]. Parameter [JK85, MTK^{+}94, SG92, Tza85, Unb85]. parameterized [Met64]. parameters [BZ90, Bra85, Dem85]. parametric [Woł85]. parametrization [Mor85]. Pareto [Gut85, ZBZ85]. participants [Ano60a, Ano77h, Ano80, Ano94d]. participation [JA94]. Particle [Gre85]. particularity [Dah85]. PASCAL [Spe85]. Pascalism [Kry85]. pass [HH63]. Pattern [SL92b, HUS92, LNYz89]. PCSL [Lud80]. pearl [HH88, TBR^{+}94]. pearl-based [TBR^{+}94]. PEARL’s [BP94]. PEGASUS [Fel60, Pay60, Ron60]. penicillin [EhGhSl89]. perfect [SJ92, SOJ92]. perfect-models [SOJ92]. Performance [BI85, MTK^{+}94, BJ92, BZ90, Kra85, MR94, Puc94]. perspective [BJT^{+}94, Pao94]. Petri [AK90, Kuz85, Sta85c, Szl77]. Petri-Netz-Maschine [Sta85c]. petrochemical [PB94]. Phase [Zha92]. Phenomena [LI92, WB92]. philosophy [Gil60, Pop85, Smi70]. phrase [BM61]. physical [OGF85]. PID [ÅHP89, HLH92]. piecewise [TAR92]. pilot [SKKG94]. pipelined [KHM81]. PL [LW69]. PL/I [LW69]. Planning [Bes92, BBS^{+}88, Fed85, Fri92, Mel88, SC94]. plant [Dov85, LI92, NKIT81, PB94, SR85]. plants [AJT94, How92, LŚT90, RK89]. Plenary [VR92]. points [MČ85]. pole [Tos85]. policies [ACDM85, MP85]. policy [BCSM94, Sch85a]. Politico [Bem69]. Politico-Social [Bem69]. poor [How92]. populations [Sar85]. position [MMO94]. power [HH92a, LŚT90, Mor92, TAT81, VM92, YY92]. practitioner [BK92, Hug92]. pre [Sjö92]. pre-processing [Sjö92]. Predicting [SOJ92]. predictive [PCY92]. Preface [Ano81, Ano89c, Di 85, EH92, Goo60, Goo61, Goo64, Haa80, Hal94, Knu90, RS92b, Sme77, Syd85b]. Preliminary [Ano60e, Ano60f]. Preprocessing [SV92]. presence [GJ85]. press [BK92]. principle [Mit88]. Principles [Bro60, Fra92b, Stă85a]. priority [ABDW94, Son90]. priority-based [Son90]. problem [ACD^{+}85, ML85]. Problems [Kal85, AA85, APPZ85, GKKP85, GGG85, GM85, Gut85, Hug92, LD85, MČ85, NK94, Obe85, RE85, Red60, Sch85c, SWR85, Syd85a, Vál85, Wit85, Wos85, dJM85]. procedural [Har92]. procedure [LMK77, Pop85, Rut61, Spi85]. procedures [GMWK92, Rut63]. Process [JX92, Lud80, NSY92, Séd90, SL92a, ARA^{+}92, AAG^{+}94, BSMS94, BTM92, Bro94, Car85, CSDM94, CBGG92, Cor94, CNV^{+}92a, DISN77, Efs92, Fis85, FJ92, GL80, GSGPAT94, GRSA94, HPP85, HK92, HGH^{+}85, IMF^{+}81, JA94, Kap77, Kar92, KSŽ92, Kuz85, LS94, LTK92, LOKK89, LS92, MS92b, MMW92, NKIT81, Pao94, Per92, PEGR80, RBSG94, RNB92, RPF^{+}92, RTC92, Rus92, RLM94, SR85, Sch80, TBR^{+}94, Tho77, VK85, Wei85, vdL94]. process-automation [TBR^{+}94]. processes [EhGhSl89, HvALT92, JST^{+}92, KCP85, LwYz89, LA92, MBK92, PMH85, PS77, PV85, Wel85, Woo63, ZEZ89]. processing [Cla63, HJLT77, KHM81, LMS94, MP85, McG69, NK92, Sjö92, Wil64]. processor [BCSM94, HKM77, KHM81, NKIT81, Yas85]. processor-sharing [Yas85]. processors [Bro69, Dek85]. PRODAT [FQVM88]. production [LwYz89]. productivity [ČŠ88, Kuv92]. professionals [Sau92]. Program [FKKK81, Ano77g, Bro60, Fri64, OTTK81, Ron60]. programmable [Chr85, TBR^{+}94]. Programming [Ano88d, Dij63b, AA85, Ano60d, Bae61, Ben85b, Buc85, Car85, Chr85, Cus60, Dov85, DG80, Elz80, FHK88, GKKP85, Gil60, Gle60, GGG85, HJ73, HPP85, Hin94, HW77, IM72, IMF^{+}81, KPR^{+}80, KSK94, Lec73, LD85, Meg60, Mos94, PFTD85, RKC85, Rob60, Sch94b, SM94, Sha63, SY74, Spe85, Tay60, WI81, Woo63, dJM85]. programs [Ahr85, Cou69, DG94, Den60, Fel60, Szl77, Szm88]. Progress [DBG63, Sam69]. Project [KKK92, ARA^{+}92, AAG^{+}94, BJJ^{+}92, DS88, FQVM88, Fri92, HK92, HBS92, Hug92, JL88, Koc88, Mel88, MHS^{+}88, Sau92, VK94, Wei88, Fri92]. projects [DLP92, Fri92, JS85, KZ88, Sto88]. proof [LQM94]. Properties [KRK92, KMP^{+}94, Rei85b, Szl77]. prospective [Bes92]. prosthetic [HUS92]. PROSUS [MAU^{+}81]. protocol [NY94]. protocols [KK88, Son90]. prototype [BJJ^{+}92]. prototyping [ADLD94, MTK^{+}94, OR88]. proving [Pir73, Szl77]. pulsed [LS94]. purpose [Lec73].

quadratic [dJM85]. Qualitative [TAR92, UJFB92, MTK^{+}94, NSY92]. quality [Bas92, BL85, CC85, DCA94, HBS92, PKS88, Rus85, Ule85]. quantitative [NSY92]. quasi [Che85b]. quasi-field [Che85b]. query [FP94]. queue [Yas85]. queueing [ACDM85].

RAM [MMO94]. RAM-1 [MMO94]. random [Bor85]. randomly [Che85a]. rapid [NIYY81]. Rapidwrite [Hum63]. reactor [Tza85, XL85]. REAKT [BBO94, KM94, VGQ94]. REAKT1 [FPCN94]. Real [ASM88, And85, BGS80, CSDM94, GL80, GK77, Han94, LN92b, LA92, NIYY81, Pao92b, Rod90, SC94, Sow81, VPM94, WI81, Åré92, ARA^{+}92, ADLD94, AFGvdS80, Ano88d, AK90, Arz92, BBC^{+}92a, BL89, BP94, Ben85b, Bes92, BJNR77, BCSM94, CR92, CNV^{+}92b, CNV^{+}92a, DG94, Dub85, Dum94, Elz80, Fan94, FQ90, Fin92, FP94, FLMS92, FHK88, FKKK81, Gaf88, GLR81, GK80, GRSA94, Gór80, HH88, HPP85, Har89, HR94, Hin94, Hoo92, HJLT77, HD85, JVB92, JHJ90, KPR^{+}80, KM94, KHP92, KJVB92, KJ92, LMS94, LŚT90, LwYz89, LN92a, LOKK89, LM89, ML92, MdCMC88, MEN77, Mes94, MTK^{+}94, Mos94, Mot92, Mot94, NY94, NLÁR94, PvWD92, PB94, PS85, RP92, RK89, RW89, RHJ^{+}94, SRRP^{+}94, SJ92, Sch94a, Sch94b, SM94, Sha63, SM89, Son90, Ste90, Szm88, SSSS94, TS90]. real [Tay80, Tem88, TZ89, TFMM81, Ush81, VGQ94, VBRD94, VHR92, Voj77, YJ89, lCMM77, vdL94]. Real-Time [Rod90, And85, BGS80, GL80, GK77, Han94, LN92b, LA92, NIYY81, Pao92b, VPM94, LMS94, Åré92, ARA^{+}92, AFGvdS80, Ano88d, AK90, Arz92, BBC^{+}92a, BL89, BP94, Ben85b, Bes92, BJNR77, BCSM94, CR92, CNV^{+}92b, Dub85, Elz80, Fan94, FQ90, Fin92, FP94, FLMS92, GLR81, HH88, HR94, Hin94, Hoo92, HJLT77, HD85, JVB92, JHJ90, KPR^{+}80, KHP92, KJVB92, KJ92, LŚT90, LwYz89, LN92a, LOKK89, LM89, ML92, MdCMC88, MEN77, Mos94, Mot92, Mot94, NY94, NLÁR94, PvWD92, PB94, RP92, RK89, RW89, RHJ^{+}94, SRRP^{+}94, SJ92, Sch94a, Sch94b, SM94, Sha63, Son90, Ste90, Szm88, SSSS94, Tay80, Tem88, Ush81, VGQ94, Voj77, YJ89, vdL94]. real-world [SM89]. reality [Dab̧85]. realization [PA88]. realizing [Tho77]. realtime [MR94, Sch77]. reasoning [BBO94, Hüb92, LN92a, RC92, UJFB92]. recognition [CVB85, HUS92, LNYz89, SL92b]. Recovery [OHSH92, PPMM94, Sto88]. recruitment [Ber92]. recursive [Cas94, Rut63]. refinement [TM85]. region [BBS^{+}88]. regional [Rei85a]. regions [Sch80]. Regularization [Sjö92]. regulators [Fle85]. reheating [WTI92]. Reinforcement [Ber92, KvD92]. Relay [GMS94]. reliability [TS90, vS92]. reliable [HH88]. reorganization [Gre85]. repair [BFH85]. repeated [TFMM81]. Report [BBG^{+}61, Ano60e, Ano60f, Nau64a]. representation [BBO94, CR92, D’I69, LTK92, PPMM94]. representations [Woo92]. represented [PV85]. Requirements [BCD^{+}80, DR94, LMS94, HBS92, Tay80, Ule85]. research [JKW85, VK94]. residual [CSF92, MMA92, PC92]. resilient [CCMM90]. resolution [Pir73]. Resource [Elz80, JS85]. resources [Fed85]. response [SOJ92]. response-times [SOJ92]. Restarting [Dov85]. Results [HBS92, Par85]. reusable [Kop92]. reuse [BK92, Fra92a, RWMC92, Sch92, Wit92]. Reusing [Wit92]. Revised [Nau64a]. Rigas [CNV^{+}92b]. right [Lei92]. rights [Elz80]. Road [MRSS94]. robot [CTW94, FS94, KHP92, KvD92, MdCMC88, MMO94, VPM94, WB85]. robots [KGN94, LL94, TKD85, VG94]. Robust [EB94, PC92, Poh85]. room [MC92]. routing [ACD^{+}85, ACDM85]. RT [FP94]. RT-SQL [FP94]. rule [Bro94, JVB92, SGZ89]. rule-based [Bro94, JVB92, SGZ89]. rulebase [SP94]. rules [HH92b, KMO^{+}92, LN92b, Pfe94]. run [BP94]. run-Time [BP94]. runs [TFMM81].

safe [HH88, VS88]. SAKO [Łuk61, Maz61]. sales [Ron60]. satellites [SKKG94]. satisfy [VHR92]. Scalarizing [Wit85]. scale [Dui85, OGF85, Stă85a, Tie85, Wei85]. schedulability [DG94]. scheduling [ABDW94, BCSM94, LwYz89, TW94]. schema [PPMM94]. scheme [CSF92, HH63]. schemes [vS92]. Scientific [Sch85a]. SEAL [Cla63]. search [Bor85, MNV85]. second [Séd90]. Selecting [Lei92]. selection [BJ92, Boe85]. self [Gre85, IM85, Poh85, Wil64, YLZ89]. self-compiling [Wil64]. self-learning [YLZ89]. self-organization [IM85]. self-reorganization [Gre85]. self-tuning [Poh85]. semiconductor [PPE^{+}94]. semilinear [Poh85]. sensor [Har89, LH92, Shi89]. sensors [MRL^{+}94]. separation [Mit88]. sequence [JS85]. sequencing [SS70]. Sequential [Arz92, PS77, SM94, dJM85]. series [HJLT77, Pop85, Tza85]. server [CNV^{+}92b]. set [Gut85, RKC85, Tho77, ZBZ85]. shaft [MS94]. shape [Hug92]. share [Han60]. sharing [NHIK77, Yas85]. shell [GSGPAT94, TZ89]. ship [MC92]. shooting [Obe85]. short [Ano60d]. should [Bri85, Rei85b]. signals [TZ89]. simple [GG90, OS60, Wil64]. Simula [IM72]. Simulation [Kel85, Ahr85, Fed85, HK85, Kal85, Kuz85, MS94, NS92, Par85, PFTD85, SAM85, Spe85, Stă85a, Tie85, TAR92, TKD85, VK85, Vor85]. sinterplants [ZEZ89]. SISO [PA88]. situation [HBS92]. situations [CVB85, KSŽ92]. skill [HH92b]. skills [BJT^{+}94, KGN94]. Slab [WTI92]. sludge [AJT94]. small [EB94, Ush81]. Smoothing [Gut85]. Social [Bem69, Dah85]. Soft [Fan94]. Software [CH85, DSD^{+}88, Fra92a, HKM77, HGH^{+}85, JH92, KLM^{+}92, PS77, PF92, RK85, Sau92, Ule85, Ano92c, AK90, Bas92, BBC^{+}92b, ČŠ88, DLP92, FHK88, Fri92, GRW88, HK92, HBS92, Hug92, JL88, Koc88, KZ88, KSŽ92, Kop92, Kuv92, LMK77, Lud80, Mel88, MHS^{+}88, Mit88, MS92b, MAU^{+}81, RWMC92, RS92a, RK89, Sch92, Sta85c, Sto88, SWR85, SSSS94, Tem88, Tem94, Tho77, TFMM81, Wei88, WHS77, Wit92]. software-process [MS92b]. SOLDA [DISN77]. solids [Gre85]. solution [ACD^{+}85]. solutions [DR94, Gut85, Mor85, Vál85]. solvent [RPF^{+}92]. solving [GGG85, Sch85c]. Some [DBG63, Loc85, Red60, Sam69, Tie85, TW94, Woo63, Wos85, Den60, Est85, Szl77]. sonar [Puc94]. sonar-based [Puc94]. source [DBG63, DISN77]. Space [PF92, Ben85a, BCSM94, SC94, Zha92]. spaces [YY92, Zha92]. Spatio [DR94]. Spatio-temporal [DR94]. special [GGG85]. specific [Fri64, RBSG94]. Specification [GLTD88, GG90, KMP^{+}94, Lud81, Ano88c, GK80, KK88, LTK92, Lud80, PS77, Sav81, TM85]. specifications [BCD^{+}80]. speculative [Mot94]. spreadsheet [KZ88]. spy [VS88]. SQL [FP94]. squares [Sim85]. Stability [BA92, GCOA92, Lun85, Mor92]. stabilizable [UW85]. stabilizers [HH92a]. stack [VBRD94]. Stackelberg [ML85]. stacks [BG64]. standard [FHK88]. STANTEC [OS60]. STANTEC-ZEBRA [OS60]. state [AA85, Ben85a, GL80, HS85, NAP92, SC94, SG92, dJM85, vS92]. state-estimator-based [vS92]. state-space [Ben85a]. state-space-models [Ben85a]. Statecharts [BDRS94]. station [HD85]. statistical [Ave92, CC85, Fis85, Rus85, Ule85, Wei85]. steady [HS85]. steady-state [HS85]. Steel [IMF^{+}81, DCA94]. steelmaking [NKIT81]. step [Tem88]. Steps [Hal77]. Stepwise [MS92b]. Stochastic [Che85a, GJ85, MP85]. storage [D’I69]. stormwater [Kie92]. strategies [LD85]. strategy [NK92]. strip [DCA94]. Structural [Rei85b, WB92]. Structure [Voi85, BM61, Iro63, LS92, Ono89, OGF85, RK89, YKI92]. structured [FP94, Loc85, MEN77]. structures [D’I69, HH88]. Structuring [Hüb92, Wal80]. studies [Rus92, RLM94, Sam69]. study [Har92, KLM^{+}92, Kuv92, LOKK89, MS94, Mot94, RKC85, SKKG94]. Subject [Ano85d, Dab̧85]. subroutines [Maz61]. substantially [Bra85]. substitute [Sjö92]. subsystem [Fan94, Lun85]. subsystems [LL94]. successful [GRSA94]. sum [Mor85]. Supervision [Per92, BTM92, CBGG92, Cor94, IGM^{+}94, JA94, LTK92, RBSG94, SRRP^{+}94]. supervisor [MC92, SM89]. Supervisory [HvALT92, Har89, LA92, PvWD92]. supplementary [HH92a]. support [BJJ^{+}92, BK92, DS88, Fan94, Fri64]. supported [OR88]. Supporting [ASM88, BK85, Mos94]. supports [Gór80]. survey [Åré92, Bro69, RNB92, SY74, Unb85]. susceptible [BFH85]. switch [HvALT92]. switched [Grz85]. synchronization [Son90]. Synchronized [Sch94a]. syntax [HJ73, Iro63]. Synthesis [YKI92, Zha92]. System [Sta85b, Unb85, Åré92, BBG^{+}88, BK92, Buc85, CNV^{+}92a, Dab̧85, DISN77, DCA94, DG80, ES77, Fan94, Fin92, Fri64, Gaf88, GCOA92, GLTD88, GLR81, GJ85, GSGPAT94, GG90, GRSA94, GK77, Han60, HH92a, HJLT77, Ili61, IMF^{+}81, IM85, JKW85, JHJ90, JX92, Kar92, KHM81, LNYz89, Lec73, LN92a, Łuk61, MdCMC88, MEN77, MTK^{+}94, Mor92, NKIT81, NS92, NHIK77, NIYY81, OGF85, Osy85, PHW60, RK89, RPF^{+}92, Sav80, Sch94b, Shi89, SGZ89, Sow81, Spe85, SWR85, TAT81, TS90, Tie85, TZ89, UOH^{+}81, Ush81, Voj77, WFR92, YJ89, ZEZ89]. systematic [HK85]. Systems [Bra85, Che85b, GCOA94, ARA^{+}92, AAG^{+}94, ADLD94, BCD92, Bak85, BFH85, BA92, BL89, BJ92, BI85, Bes92, BJNR77, BCSM94, BDRS94, Bor85, BBS^{+}88, BLA94, CSF92, CVB85, CCMM90, CR92, Cou69, Dah85, DHD85, Dui85, Eng85, FPCN94, FKKK81, GTW94, GCOA92, GGG85, GK80, GM85, HKM77, HH88, Hal81, HS85, Hoo92, HGH^{+}85, Hüb92, HD85, JH92, JK85, Kal85, KM94, Kop92, KJ92, Kun85, Lio85, Loc85, LOKK89, Lun85, LS92, MR94, MSL^{+}94, Mar92, MBH94, NK94, NS92, Ono89, OGF85, PvWD92, PC92, PKS88, Poh85, PV85, Put85, RP92, Rei85a, Rei85b, RW89, RHJ^{+}94, RK85, SP94, SJ92, rnhWS85, Sch94a, Sch80, Sch77, SM94, SG85, Sha63, SAM85, Son90, Stă85a, Tay60, Tay80, Tho77, TM85, Tos85, UW85, VM92, VK85, VS88, Wal80]. systems [WHS77, Woo92, Wos85, Zha92, vdL94].

tandem [JHJ90]. tank [Wei85]. target [HGP92]. target-directed [HGP92]. Task [TBR^{+}94, CNV^{+}92b, GP94, MSL^{+}94, Sch94b]. Task-configuration [TBR^{+}94]. task-oriented [Sch94b]. tasks [BCSM94, LCM94]. team [JL88]. technically [Sch85a]. technique [DS88]. Techniques [BBC^{+}92b, BJT^{+}94, CC85, Dov85, Efs92, HR94, LD85, Sch80, SM89]. technological [MBK92]. technology [Kal85, KM81]. telerobotic [HGM92]. temperature [WTI92]. Temporal [BBO94, RC92, DR94, Hüb92, RP92]. terms [BM61]. test [GTW94]. tests [Ave92]. textual [Ben85b]. their [D’I69]. theorem [Pir73]. theoretic [Dui85, Tam85, UW85]. theory [Che85b, Dah85, GCOA92, Lio85, Loc85, ML85, WB92]. three [Ano60d]. TIDE [Hum60]. Time [ASM88, LMS94, Rod90, Åré92, ARA^{+}92, ADLD94, AFGvdS80, And85, Ano88d, AK90, Arz92, BBC^{+}92a, BL89, BP94, Ben85b, Bes92, BJNR77, BGS80, BCSM94, CSDM94, Che85a, CR92, CNV^{+}92b, CNV^{+}92a, DG94, Dub85, Elz80, Fan94, FQ90, Fin92, FP94, FLMS92, FHK88, FKKK81, Gaf88, GL80, GLR81, GK80, GRSA94, GK77, Gór80, HH88, HPP85, Han94, Har89, HR94, Hin94, Hoo92, HJLT77, HD85, JVB92, JS85, JHJ90, KPR^{+}80, KM94, KHP92, KJVB92, KJ92, LNYz89, LŚT90, LwYz89, LN92b, LA92, LN92a, LOKK89, LM89, ML92, MdCMC88, MEN77, Mes94, MTK^{+}94, Mos94, Mot92, Mot94, NY94, NLÁR94, NIYY81, Pao92b, PvWD92, PB94, PS85, Pop85, RP92, RK89, RW89, RHJ^{+}94, SRRP^{+}94, SJ92, Sau92, Sch94a, Sch94b, SM94, Sha63, SC94, Son90, Sow81, Ste90, Szm88, SSSS94, TS90]. time [Tay80, Tem88, TZ89, TFMM81, Ush81, VPM94, VGQ94, VBRD94, VHR92, Voj77, WI81, WB85, YJ89, ZEZ89, lCMM77, vdL94]. time-delay [LNYz89, ZEZ89]. time-optimal [JS85]. time-varying [Che85a]. times [SOJ92]. timing [KMP^{+}94]. Today [BCD^{+}80]. tomorrow [BCD^{+}80]. tool [ASM88, BK92, DS88, Fis85, Fri92, GRW88, KMP^{+}94, Lei92, PA88, RBSG94, Sta85c, TFMM81, WI81]. tool-to-organisation [BK92]. toolkit [VGQ94]. tools [BBC^{+}92b, BSMS94, HBS92, JH92, MMA92, Meg60, Tho77, Wit92]. Top [Hoo92]. Top-down [Hoo92]. tradeoff [TM85]. traditional [BJT^{+}94]. traffic [HJ94]. Training [MBH94, BJ92, Sau92, Sjö92]. trajectories [Pie85]. Trajectory [Put85]. transit [NIYY81]. transition [NAP92]. translating [Den60]. translation [Cle60, HH63]. Translator [Ano63a, Ano63e, Cli61, Dij63a, Dij63c, Dij63d, Hig63]. Transport [NY94, BBG^{+}88, KCP85]. TREAC [PHW60]. treated [GKKP85]. treatment [Obe85]. Trends [Mot94, Gle60, Hal81, Loc85]. trilogy [PS85]. truncated [Che85a]. tundra [Sar85]. tuner [HLH92]. tuning [FCWM92, KGN94, MTK^{+}94, Poh85]. tunnel [Kra85]. turbines [LH92]. tutorial [Eng72]. Two [Ano60e, Ano60f, Dij63a, LQM94, Pie85, WB85]. two-link [WB85].

ultrasonic [MRL^{+}94]. uncertain [BA92, MMO94]. UNCOL [Ste61]. unifield [LN92a]. Universal [KCP85, KPR^{+}80, Red60]. UNIX [BP94, BSMS94]. Use [Sch85c, BG64, DG80, FHK88, GK80, HPP85, Ili61, Iro63, Maz61, Rut63, Sav81, VHR92]. used [NKIT81, Put85]. user [ES77, Fed85, JA94, SNG94]. user-interface [Fed85]. uses [MNV85]. Using [CCMM90, FLMS92, Fri92, Hug92, Jen92, SM89, AK90, Ben85a, BSMS94, BDRS94, Fle85, GCOA92, HJLT77, Kar92, Lar92, LN92b, MRSS94, MS92a, Mor92, Pop85, Rus92, RLM94, Sam69, Sch80, SG92, Tza85, UOH^{+}81, Unb85, VG94, WTI92, YKI92, YY92, ZMM94]. UTC [Sch94a]. utilising [LS92]. utility [Tam85].

validating [BDRS94]. validation [BBG^{+}88, Sta85c]. varying [Che85a]. vector [Vál85, Voi85]. vehicle [GP94, HGP92]. verification [AK90, Szm88, SSSS94]. version [Sch85b, YY92]. via [GMWK92, LNYz89]. view [Elz80]. views [Sam61b]. virtual [Grz85, UOH^{+}81]. vision [MRSS94, RW89, VPM94]. vision-based [VPM94]. Visual [Hin94]. voltage [Mor92]. volume [Ano63b, Ano63c, Ano64a]. voting [CSF92].

wage [Fan85]. wall [Puc94]. wall-following [Puc94]. Walsh [Tza85]. waste [AJT94]. water [AJT94, Fed85, Put85, Rei85a]. way [NKIT81]. weekly [Loc85]. weighted [GCOA94]. weighted-output [GCOA94]. wheel [MRL^{+}94]. wheel-chair [MRL^{+}94]. which [Gór80, MNV85]. wind [Kra85]. winder [MS94]. without [KK88]. Work [RS92a, HBS92, HD85, RWMC92]. works [IMF^{+}81]. Workshop [Ano94d]. workstation [Kie92]. world [SM89]. wwtp [SRRP^{+}94].

X1 [Ano63a, Dij63a, Dij63c]. xenon [XL85].

Yawata [NKIT81].

ZEBRA [OS60]. zero [Mor85]. zero-sum [Mor85]. zeros [Dui85].

References

Abdelrahman:1985:OCM

[AA85] O. E. Abdelrahman and B. M. Abuelnasr. On the orthogonal collocation and mathematical programming approach for state constrained optimal control problems. Annual Review in Automatic Programming, 13 (part 2):1–7, 1985. CODEN ????

            ISSN ???? URL https://www.sciencedirect.com/science/
            article/pii/0066413885904562.

Alarcon:1994:IMA

[AAG^{+}94] I. Alarcón, X. Alamán, P. Gómez, S. Romero, J. A. Aguilar-Crespo, M. Campos, P. Serrahima, J. M. Domínguez, P. Rodríguez-Marín, and E. De Pablo. An integration methodology and architecture for intelligent systems in process control: The hint project. Annual Review in Automatic Programming, 19:353–358, 1994.

            CODEN ???? ISSN ???? URL https://www.sciencedirect.
            com/science/article/pii/0066413894900906.

Audsley:1994:IBE

[ABDW94] N. C. Audsley, A. Burns, R. I. Davis, and A. J. Wellings. Integrating best effort and fixed priority scheduling. Annual Review in Automatic Programming, 18:45–50, 1994. CODEN ????

            ISSN ???? URL https://www.sciencedirect.com/science/
            article/pii/0066413894900094.

Aicardi:1985:DCL

[ACD^{+}85] M. Aicardi, G. Casalino, F. Davoli, R. Minciardi, and R. Zoppoli. A decentralized closed-loop solution to the routing problem in networks. Annual Review in Automatic Programming, 13 (part 2):9–17, 1985. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413885904574.

Aicardi:1985:DOR

[ACDM85] M. Aicardi, G. Casalino, F. Davoli, and R. Minciardi. Development of optimal routing policies in a queueing network. Annual Review in Automatic Programming, 12 (part 1):404–407, 1985.

CODEN ???? ISSN ???? URL https://www.sciencedirect.
com/science/article/pii/0066413885900722.

Alonso:1994:EDP

[ADLD94] A. Alonso, J. C. Dueñas, G. León, and J. A. De La Puente. An environment for distributed prototyping of real time systems. Annual Review in Automatic Programming, 18:21–26, 1994. CO-

DEN ???? ISSN ???? URL https://www.sciencedirect.com/
science/article/pii/0066413894900051.

Altaber:1980:IAI

[AFGvdS80] J. Altaber, V. Frammery, C. Gareyte, and P. D. V. van der Stok. An interpretive approach for interactive real-time control through a network. Annual Review in Automatic Programming, 10:1–7, 1980. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/0066413882900027.

An:1992:KBC

[AH92] P. E. An and C. J. Harris. Knowledge-based control. Annual Review in Automatic Programming, 17:87–96, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/S0066413809910163.

Aastrom:1989:TIP

[ÅHP89] K. J. Åström, C. C. Hang, and P. Persson. Towards intelligent PID control. Annual Review in Automatic Programming, 15 (part 2):53–58, 1989. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413889900104.

Ahrens:1985:BAM

[Ahr85] Klaus Ahrens. Behaviour analysis of MODULA-2 programs by simulation. Annual Review in Automatic Programming, 12 (part 1):479–482, 1985. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/0066413885900916.

Alex:1994:FCA

[AJT94] J. Alex, U. Jumar, and R. Tschepetzki. A fuzzy controller for activated sludge waste water plants. Annual Review in Automatic Programming, 19:61–66, 1994. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413894900434.

Arendt:1990:MVR

[AK90] F. Arendt and B. Klühe. Modelling and verification of real-time software using interpreted Petri nets. Annual Review in Automatic Programming, 15 (part 1):35–40, 1990. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/006641389090024L.

Anderson:1985:RTA

[And85] T. Anderson. Real-time application of the iPSC ˘2122 concurrent computer. Annual Review in Automatic Programming, 13 (part 1):55–61, 1985. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413885904446.

Anonymous:1960:AFL

[Ano60a] Anonymous. Appendix four: List of participants. Annual Review in Automatic Programming, 1:295–300, 1960. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413860900495.

Anonymous:1960:AOC

[Ano60b] Anonymous. Appendix one (a): On computable numbers, with an application to the Entscheidungsproblem. Annual Review in Automatic Programming, 1:230–264, 1960. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413860900458.

Anonymous:1960:AOB

[Ano60c] Anonymous. Appendix one (b): On computable numbers, with an application to the Entscheidungsproblem. A correction. Annual Review in Automatic Programming, 1:265–267, 1960. CO-

DEN ???? ISSN ???? URL https://www.sciencedirect.com/
science/article/pii/006641386090046X.

Anonymous:1960:ATA

[Ano60d] Anonymous. Appendix three: Automatic programming — a short bibliography. Annual Review in Automatic Programming, 1:291–294, 1960. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/0066413860900483.

Anonymous:1960:ATP

[Ano60e] Anonymous. Appendix two: Preliminary report of ACM-GAMM committee on an international algebraic language. Annual Review in Automatic Programming, 1:268–290, 1960. CO-DEN ARVAAM. ISSN 0066-4138 (print), 1878-545X (elec-

tronic).  URL https://www.sciencedirect.com/science/
article/pii/0066413860900471.

Anonymous:1960:LPR

[Ano60f] Anonymous. Appendix two: Preliminary report of ACM-GAMM committee on an international algebraic language. Annual Review in Automatic Programming, 1:268–290, 1960. CODEN ARVAAM. ISSN 0066-4138 (print), 1878-545X (electronic).

Anonymous:1960:EB

[Ano60g] Anonymous. Editorial board. Annual Review in Automatic Programming, 1:i, 1960. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413860900252.

Anonymous:1961:CDC

[Ano61] Anonymous. A critical discussion of COBOL. Annual Review in Automatic Programming, 2:293–304, 1961. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/S0066413861800128.

Anonymous:1963:ATX

[Ano63a] Anonymous. An ALGOL 60 translator for the X1. Annual Review in Automatic Programming, 3:329–345, 1963. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/S0066413863800154.

Anonymous:1963:CVa

[Ano63b] Anonymous. Contents of volume 1. Annual Review in Automatic Programming, 3:357–358, 1963. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/S0066413863800178.

Anonymous:1963:CVb

[Ano63c] Anonymous. Contents of volume 2. Annual Review in Automatic Programming, 3:359–360, 1963. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/S006641386380018X.

Anonymous:1963:EN

[Ano63d]  Anonymous.  Editorial note.  Annual Review in Auto-

matic Programming, 3:325, 1963. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/S0066413863800130.

Anonymous:1963:MTA

[Ano63e] Anonymous. Making a translator for ALGOL 60. Annual Review in Automatic Programming, 3:347–356, 1963. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/S0066413863800166.

Anonymous:1964:CV

[Ano64a] Anonymous. Contents of volume 1. Annual Review in Automatic Programming, 4:259–263, 1964. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413864900217.

Anonymous:1964:EB

[Ano64b] Anonymous. Editorial board. Annual Review in Automatic Programming, 4:i, 1964. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413864900114.

Anonymous:1969:EB

[Ano69] Anonymous. Editorial board. Annual Review in Automatic Programming, 5 (part 1):i, 1969. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413869900354.

Anonymous:1977:Da

[Ano77a] Anonymous. Discussion. Annual Review in Automatic Programming, 8:29–33, 1977. CODEN ???? ISSN ???? URL https://

www.sciencedirect.com/science/article/pii/0066413877900064.

Anonymous:1977:Db

[Ano77b] Anonymous. Discussion. Annual Review in Automatic Programming, 8:85–90, 1977. CODEN ???? ISSN ???? URL https://

www.sciencedirect.com/science/article/pii/006641387790012X.

Anonymous:1977:Dc

[Ano77c] Anonymous. Discussion. Annual Review in Automatic Programming, 8:125–129, 1977. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413877900179.

Anonymous:1977:Dd

[Ano77d] Anonymous. Discussion. Annual Review in Automatic Programming, 8:163–168, 1977. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413877900222.

Anonymous:1977:De

[Ano77e] Anonymous. Discussion. Annual Review in Automatic Programming, 8:207–210, 1977. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413877900271.

Anonymous:1977:Df

[Ano77f] Anonymous. Discussion. Annual Review in Automatic Programming, 8:229–232, 1977. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413877900295.

Anonymous:1977:IPC

[Ano77g] Anonymous. International program committee. Annual Review in Automatic Programming, 8:233, 1977. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413877900301.

Anonymous:1977:LP

[Ano77h] Anonymous. List of participants. Annual Review in Automatic Programming, 8:235–237, 1977. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413877900313.

Anonymous:1980:LP

[Ano80] Anonymous. List of participants. Annual Review in Automatic Programming, 10:129–131, 1980. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413882900179.

Anonymous:1981:P

[Ano81] Anonymous. Preface. Annual Review in Automatic Programming, 11:vi, 1981. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/006641388190001X.

Anonymous:1985:AIb

[Ano85a] Anonymous. Author index. Annual Review in Automatic Programming, 13 (part 1):107, 1985. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413885904537.

Anonymous:1985:AIa

[Ano85b] Anonymous. Author index. Annual Review in Automatic Programming, 13 (part 2):209, 1985. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413885904835.

Anonymous:1985:LA

[Ano85c] Anonymous. List of authors. Annual Review in Automatic Programming, 12 (part 1):493–495, 1985. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413885900953.

Anonymous:1985:SI

[Ano85d] Anonymous. Subject index. Annual Review in Automatic Programming, 13 (part 1):109, 1985. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413885904549.

Anonymous:1988:AIa

[Ano88a] Anonymous. Author index. Annual Review in Automatic Programming, 14 (part 1):87, 1988. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413888900158.

Anonymous:1988:AIb

[Ano88b] Anonymous. Author index. Annual Review in Automatic Programming, 14 (part 2):95, 1988. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/006641389090017L.

Anonymous:1988:FS

[Ano88c] Anonymous. Formal specification. Annual Review in Automatic Programming, 14 (part 1):83, 1988. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413888900134.

Anonymous:1988:PLR

[Ano88d] Anonymous. Programming languages for real-time applications. Annual Review in Automatic Programming, 14 (part 1):85, 1988.

CODEN ???? ISSN ???? URL https://www.sciencedirect.
com/science/article/pii/0066413888900146.

Anonymous:1989:AI

[Ano89a] Anonymous. Author index. Annual Review in Automatic Programming, 15 (part 2):111, 1989. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413889900219.

Anonymous:1989:KI

[Ano89b] Anonymous. Keyword index. Annual Review in Automatic Programming, 15 (part 2):113, 1989. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413889900220.

Anonymous:1989:P

[Ano89c] Anonymous. Preface. Annual Review in Automatic Programming, 15 (part 2):vii, 1989. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413889900013.

Anonymous:1990:AI

[Ano90a] Anonymous. Author index. Annual Review in Automatic Programming, 15 (part 1):87, 1990. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/006641389090033N.

Anonymous:1990:KI

[Ano90b] Anonymous. Keyword index. Annual Review in Automatic Programming, 15 (part 1):89, 1990. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/006641389090034O.

Anonymous:1992:AIa

[Ano92a] Anonymous. Author index. Annual Review in Automatic Programming, 16 (part 1):165, 1992. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/006641389190028A.

Anonymous:1992:AIb

[Ano92b] Anonymous. Author index. Annual Review in Automatic Programming, 17:527–528, 1992. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/S0066413809910850.

Anonymous:1992:BMS

[Ano92c] Anonymous. Bootstrap — measuring software management capabilities. first findings in Europe. Annual Review in Automatic Programming, 167 (part 2):171–177, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/006641389290026L.

Anonymous:1992:KIa

[Ano92d] Anonymous. Keyword index. Annual Review in Automatic Programming, 16 (part 1):167–168, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/006641389190029B.

Anonymous:1992:KIb

[Ano92e] Anonymous. Keyword index. Annual Review in Automatic Programming, 17:529–531, 1992. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/S0066413809910862.

Anonymous:1994:AIa

[Ano94a] Anonymous. Author index. Annual Review in Automatic Programming, 18:187, 1994. CODEN ???? ISSN ???? URL https://

www.sciencedirect.com/science/article/pii/0066413894900329.

Anonymous:1994:AIb

[Ano94b] Anonymous. Author index. Annual Review in Automatic Programming, 19:389–390, 1994. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413894900965.

Anonymous:1994:KI

[Ano94c] Anonymous. Keyword index. Annual Review in Automatic Programming, 18:189–190, 1994. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413894900337.

Anonymous:1994:WP

[Ano94d] Anonymous. Workshop participants. Annual Review in Automatic Programming, 18:191–192, 1994. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413894900345.

Aluffi-Pentini:1985:ICM

[APPZ85] F. Aluffi-Pentini, V. Parisi, and F. Zirilli. An inexact continuous method in complementarity problems. Annual Review in Automatic Programming, 13 (part 2):19–26, 1985. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413885904586.

Alaman:1992:KBS

[ARA^{+}92] X. Alamán, S. Romero, C. Aguirre, P. Serrahima, R. Muñoz, V. López, J. Dorronsoro, and E. de Pablo. Knowledge-based systems for real-time process control: The MIP project. Annual Review in Automatic Programming, 17:391–396, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/S0066413809910643.

Aaren:1992:SCR

[Åré92] K.-E. Årén. A survey of commercial real-time expert system environments. Annual Review in Automatic Programming, 17:

483–490, 1992. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/S0066413809910783.

Arzen:1992:SFC

[Arz92] K.-E Arzén. Sequential function charts for knowledge-based, real-time applications. Annual Review in Automatic Programming, 16 (part 1):91–96, 1992. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413891900154.

Alrebdawi:1988:STO

[ASM88] G. Alrebdawi, J. J. Skubich, and Y. Martinez. Supporting tool for object oriented design of real time applications. Annual Review in Automatic Programming, 14 (part 1):29–33, ???? 1988. CODEN ARVAAM. ISSN 0066-4138 (print), 1878-545X

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0066413888900055.

Averbukh:1992:NNM

[Ave92] E. A. Averbukh. Neural network models and statistical tests as flexible base for intelligent fault diagnosis. Annual Review in Automatic Programming, 17:259–266, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/S0066413809910436.

Barreiro:1992:SFS

[BA92] A. Barreiro and J. Aracil. Stability of feedback systems with uncertain dynamics. Annual Review in Automatic Programming, 17:127–132, 1992. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/S0066413809910229.

Baecker:1961:GCP

[Bae61] H. D. Baecker. The growth of a commercial programming language. Annual Review in Automatic Programming, 2:305– 323, 1961. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/S006641386180013X.

Bakule:1985:ODD

[Bak85] L. Bakule. On the overlapping decomposition of dynamic systems. Annual Review in Automatic Programming, 12 (part 1): 263–266, 1985. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/0066413885900394.

Bassler:1992:EIS

[Bas92] T. Bassler. Experiences in introducing software quality assurance. Annual Review in Automatic Programming, 167 (part 2): 91–94, 1992. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/006641389290014G.

Barber:1992:NPR

[BBC^{+}92a] F. Barber, V. Botti, A. Crespo, D. Gallardo, and E. Onaindía. New paradigms for real-time control. Annual Review in Automatic Programming, 17:459–465, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/S0066413809910758.

Bazzana:1992:TTS

[BBC^{+}92b] G. Bazzana, R. Brigliadori, R. Cole, K. Kirkwood, and F. Seigneur. Techniques and tools for software assessment and certification. Annual Review in Automatic Programming, 167 (part 2):153–160, 1992. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/006641389290024J.

Backus:1961:ARA

[BBG^{+}61] J. W. Backus, F. L. Bauer, J. Green, C. Katz, J. Mc-Carthy, P. Naur, A. J. Perlis, H. Rutishauser, J. H. Wegstein, A. van Wijngaarden, and M. Woodger. Appendix: Report on the algorithmic language ALGOL 60. Annual Review in Automatic Programming, 2:351–390, January 1961. CO-DEN ARVAAM. ISSN 0066-4138 (print), 1878-545x (elec-

tronic).  URL https://www.sciencedirect.com/science/
article/pii/S0066413861800165.

Belli:1988:DVF

[BBG^{+}88] F. Belli, H. Bonin, H. Gerdes, W. Filipowicz, and P. Jedrzejowicz. Development and validation of FIREX — a knowledge-based system for the transport of dangerous goods and fire department consulting. Annual Review in Automatic Programming, 14 (part 2):69–74, 1988. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/006641389090013H.

Barber:1994:TDR

[BBO94] Federico Barber, Vicente Botti, and Eva Onaindía. Temporal data representation and reasoning in REAKT. Annual Review in Automatic Programming, 19:347–352, 1994. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413894900892.

Brusic:1988:CIS

[BBS^{+}88] V. Brusić, Z. Boz̈ović, Z. Stojiljković, D. Djaković, and M. Trajković. Clinical information systems planning in Belgrade health care region. Annual Review in Automatic Programming, 14 (part 2):27–29, 1988. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/006641389090005C.

Bianchi:1980:RST

[BCD^{+}80] Mme G. Bianchi, B. T. Cronhjort, W.-M Dehnert, J. Ludewig, and B. Taylor. Requirements specifications — today and tomorrow. Annual Review in Automatic Programming, 10:121– 127, 1980. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/0066413882900167.

Badle:1992:TGM

[BCD92] C. Badle, C. Castel, and O. Duffaut. Towards a general multimodel-based methodology for diagnosis systems. Annual Review in Automatic Programming, 17:293–297, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/S0066413809910485.

Bo-Chao:1994:LST

[BCSM94] Cheng Bo-Chao, Alexander D. Stoyenko, and Thomas J. Marlowe. Least-space-time-first scheduling algorithm: a policy for complex real-time tasks in multiple processor systems. Annual Review in Automatic Programming, 18:33–38, 1994. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413894900078.

Boissier:1994:HSM

[BDRS94] R. Boissier, B. Dima, D. Razafindramary, and T. Soriano. Hybrid systems modelling and validating using Statecharts and Grafcet. Annual Review in Automatic Programming, 18:73–79, 1994. CO-

DEN ???? ISSN ???? URL https://www.sciencedirect.com/
science/article/pii/0066413894900140.

Bemer:1969:PSH

[Bem69] Robert W. Bemer. A politico-social history of Algol? (with a chronology in the form of a log book). Annual Review in Automatic Programming, 5:151–238, 1969. CODEN ARVAAM. ISSN 0066-4138 (print), 1878-545X (electronic).

Bendel:1985:GNL

[Ben85a] Udo Bendel. A global non-linear and dependent dynamical state-space identification by using local linear dynamical state-space-models. Annual Review in Automatic Programming, 12 (part 1):344–351, 1985. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/006641388590059X.

Benedick:1985:ITG

[Ben85b] D. R. Benedick. Integration of textual and graphic languages for real-time programming. Annual Review in Automatic Programming, 13 (part 1):63–69, 1985. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413885904458.

Bersini:1992:RLR

[Ber92] H. Bersini. Reinforcement learning and recruitment mechanism for adaptive distributed control. Annual Review in Automatic Programming, 17:467–473, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/S006641380991076X.

Bestavros:1992:PES

[Bes92] A. Bestavros. Planning for embedded systems: a realtime prospective. Annual Review in Automatic Programming, 16 (part 1):59–68, 1992. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413891900109.

Bar:1985:MCL

[BFH85] M. Bär, K. Fischer, and G. Hertel. A method of calculation for loss systems susceptible to failures with any number of repair channels and its application to delay systems. Annual Review in Automatic Programming, 12 (part 1):352–356, 1985. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413885900606.

Baecker:1964:CUS

[BG64] H. D. Baecker and B. J. Gibbens. A commercial use of stacks. Annual Review in Automatic Programming, 4:183–191, 1964. CO-

DEN ???? ISSN ???? URL https://www.sciencedirect.com/
science/article/pii/0066413864900187.

Berthaud:1973:ICC

[BG73] M. Berthaud and M. Griffiths. Incremental compilation and conversational interpretation. Annual Review in Automatic Programming, 7 (part 2):95–114, 1973. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413873900025.

Biewald:1980:RTF

[BGS80] J. Biewald, P. Goehner, and H. Schelling. Real-time features of EPOS: Formulation, evaluation and documentation. Annual Review in Automatic Programming, 10:95–101, 1980. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413882900131.

Babuska:1992:MFD

[BH92] R. Babuška and P. Horáček. Monitoring and fault diagnosis. Annual Review in Automatic Programming, 17:189–194, 1992. CO-

DEN ???? ISSN ???? URL https://www.sciencedirect.com/
science/article/pii/S0066413809910321.

Bergholz:1985:PMC

[BI85] G. Bergholz and K. Irmscher. Performance modeling of computer systems. Annual Review in Automatic Programming, 12 (part 1):469–473, 1985. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413885900886.

Bendtsen:1992:ITD

[BJ92] A. B. Bendtsen and Niels Jensen. The influence of training data selection on performance of neural networks for control of non-linear systems. Annual Review in Automatic Programming, 17:55–60, 1992. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/S0066413809910114.

Biro:1992:FID

[BJJ^{+}92] M. Biró, F. Jamrik, G. Janek, E. Knuth, and T. Remzsô. Factors influencing the design of a prototype decision support application for a distributed object management project. Annual Review in Automatic Programming, 167 (part 2):129–132, 1992.

CODEN ???? ISSN ???? URL https://www.sciencedirect.
com/science/article/pii/006641389290020P.

Bezivin:1977:MAD

[BJNR77] Jean Bézivin, Yvon Jégou, Jean-Louis Nebut, and Robert Rannou. A methodological approach to the design of real-time operating systems. Annual Review in Automatic Programming, 8: 153–162, 1977. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/0066413877900210.

Bejder:1994:CPW

[BJT^{+}94] H. Bejder, B. Holbæk Jensen, E. Thomsen, L. lundgaard Andresen, and S. Borchersen. Commercial perspective when applying ai techniques and traditional IT skills. Annual Review in Automatic Programming, 19:359–363, 1994. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413894900914.

Bronisz:1985:SMD

[BK85] P. Bronisz and L. Kru/ś. Supporting of multicriterial decision making in the case of cooperative game. Annual Review in Automatic Programming, 12 (part 1):185–189, 1985. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413885900242.

Boldyreff:1992:PRS

[BK92] C. Boldyreff and U. Krohn. The practitioner reuse support system (press): a consideration from the aspect of tool-to-organisation interconnection. Annual Review in Automatic Programming, 167 (part 2):55–62, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/006641389290009E.

Bekic:1973:IA

[BL73] H. Bekić and Olivier Lecarme. An introduction to ALGOL 68. Annual Review in Automatic Programming, 7 (Part 3):143–169, 1973. CODEN ???? ISBN 0-08-017049-8, 0-08-017128-1. ISSN

0066-4138. URL https://www.sciencedirect.com/science/
article/pii/0066413873900049.

Bossert:1985:KQP

[BL85] J. L. Bossert and J. A. LaBarr. The Kodak quality package. Annual Review in Automatic Programming, 13 (part 1):17–23, 1985.

CODEN ???? ISSN ???? URL https://www.sciencedirect.
com/science/article/pii/0066413885904380.

Beck:1989:AME

[BL89] Th Beck and R. J. Lauber. Asynchronous methods for expert systems in real-time applications. Annual Review in Automatic Programming, 15 (part 2):75–78, 1989. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413889900141.

Burgess:1994:BMK

[BLA94] P. Burgess, M. J. Livesey, and C. Allison. BED: a multithreaded kernel for embedded systems. Annual Review in Automatic Programming, 18:133–138, 1994. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/006641389490023X.

Brooker:1961:DMA

[BM61] R. A. Brooker and D. Morris. A description of Mercury Autocode in terms of a phrase structure language. Annual Review in Automatic Programming, 2:29–65, 1961. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/S0066413861800037.

Bull:1985:LBD

[BM85] G. M. Bull and R. J. Mitchell. A link between data abstraction and concurrency. Annual Review in Automatic Programming, 13 (part 1):49–53, 1985. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413885904434.

Brooker:1963:CC

[BMMR63] R. A. Brooker, I. R. MacCallum, D. Morris, and J. S. Rohl. The compiler compiler. Annual Review in Automatic Programming, 3:229–275, 1963. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/S0066413863800099.

Boettner:1985:FAS

[Boe85] R. Boettner. A fuzzy approach for the selection from a limited number of (nondominated) alternatives. Annual Review in Automatic Programming, 12 (part 1):259–262, 1985. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413885900382.

Booth:1960:OA

[Boo60] Andrew D. Booth. Opening address. Annual Review in Automatic Programming, 1:1–7, 1960. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413860900276.

Born:1985:OCS

[Bor85] Joachim Born. Optimizing complex systems by means of adaptively controlled random search. Annual Review in Automatic Programming, 12 (part 1):418–420, 1985. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413885900758.

Bosma:1977:HMC

[Bos77] H. Bosma. How to master complexity: Decomplexing. Annual Review in Automatic Programming, 8:xi–xiv, 1977. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413877900027.

Belschner:1994:MPH

[BP94] Ralf Belschner and Carlos E. Pereira. Mapping PEARL’s highlevel real-time constructs to a C run-time library under realtime UNIX. Annual Review in Automatic Programming, 18: 151–156, 1994. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/0066413894900264.

Brack:1985:SSD

[Bra85] Georg Brack. Systems with substantially distributed parameters. Annual Review in Automatic Programming, 12 (part 1): 421–424, 1985. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/006641388590076X.

Bristol:1985:HCL

[Bri85] E. H. Bristol. How control languages should be designed, with an integrated example design. Annual Review in Automatic Programming, 13 (part 1):93–96, 1985. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413885904501.

Brooker:1960:MAP

[Bro60] R. A. Brooker. MERCURY Autocode: Principles of the program library. Annual Review in Automatic Programming, 1: 93–110, 1960. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/0066413860900355.

Brown:1969:SMP

[Bro69] P. J. Brown. A survey of macro processors. Annual Review in Automatic Programming, 6 (part 2):37–88, 1969. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413869900019.

Brockmann:1994:CFR

[Bro94] W. Brockmann. Comparison of fuzzy, rule-based, and conventional process control. Annual Review in Automatic Programming, 19:91–96, 1994. CODEN ???? ISSN ???? URL https://

www.sciencedirect.com/science/article/pii/0066413894900485.

Bourdache-Siguerdidjane:1985:CON

[BSF85] H. Bourdache-Siguerdidjane and M. Fliess. On the computation of optimal nonlinear feedback controls. Annual Review in Automatic Programming, 13 (part 2):27–30, 1985. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413885904598.

Blanes:1994:DPC

[BSMS94] J. F. Blanes, J. Simó, M. Martínez, and J. S. Senent. Developing E.S. for process control using Unix based tools. Annual Review in Automatic Programming, 19:383–387, 1994. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413894900957.

Bousson:1992:CCM

[BTM92] K. Bousson and L. Travé-Massuyès. A computational causal model for process supervision. Annual Review in Automatic Programming, 17:133–139, 1992. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/S0066413809910230.

Buchner:1985:HGI

[Buc85] M. Buchner. A hierarchical graphics interface for control system programming. Annual Review in Automatic Programming, 13 (part 1):91–92, 1985. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413885904495.

Bradford:1961:MI

[BW61] Donald H. Bradford and Mark B. Wells. MADCAP II. Annual Review in Automatic Programming, 2:115–140, 1961. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/S0066413861800062.

Bull:1985:F

[BW85] Gordon M. Bull and Theodore J. Williams. Foreword. Annual Review in Automatic Programming, 13 (part 1):vii, 1985. CO-

DEN ???? ISSN ???? URL https://www.sciencedirect.com/
science/article/pii/0066413885904355.

Blesa:1990:IPP

[BZ90] P. Blesa and R. Zambardino. Incidence of parameters in the performance of a distributed database. Annual Review in Automatic Programming, 15 (part 1):41–46, 1990. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/006641389090025M.

Camacho:1994:NNB

[CA94] E. F. Camacho and M. R. Arahal. Neural network based adaptive control. Annual Review in Automatic Programming, 19: 13–24, 1994. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/006641389490037X.

Caro:1985:PLB

[Car85] R. H. Caro. A programming language for batch process control. Annual Review in Automatic Programming, 13 (part 1): 97–101, 1985. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/0066413885904513.

Caspi:1994:TRB

[Cas94] P. Caspi. Towards recursive block diagrams. Annual Review in Automatic Programming, 18:81–85, 1994. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413894900159.

Cauvin:1992:MBD

[CBGG92] S. Cauvin, B. Braunschweig, P. Galtier, and Y. Glaize. Modelbased diagnosis for continuous process supervision: The alexip experience. Annual Review in Automatic Programming, 17:415– 421, 1992. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/S0066413809910680.

Conroy:1985:ILS

[CC85] R. C. Conroy and G. Coleman. Implementing on-line statistical quality control techniques. Annual Review in Automatic Programming, 13 (part 1):25–27, 1985. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413885904392.

Cerrada:1990:UCI

[CCMM90] J. A. Cerrada, M. Collado, R. Morales, and J. J. Moreno. Using conversations to implement resilient objects in distributed systems. Annual Review in Automatic Programming, 15 (part 1):47–52, 1990. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/006641389090026N.

Clement:1985:SDA

[CH85] K. H. Clement and L. Hallager. Software for data acquisition and control. Annual Review in Automatic Programming, 13 (part 1):103–105, 1985. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413885904525.

Chen:1985:SAA

[Che85a] Han-Fu Chen. Stochastic approximation algorithm randomly truncated at time-varying bounds. Annual Review in Automatic Programming, 12 (part 1):380–383, 1985. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413885900679.

Chen:1985:STQ

[Che85b] Wen-De Chen. Systems theory over the quasi-field. Annual Review in Automatic Programming, 12 (part 1):143–146, 1985. CO-

DEN ???? ISSN ???? URL https://www.sciencedirect.com/
science/article/pii/0066413885900151.

Christensen:1985:GPL

[Chr85] J. H. Christensen. Graphical programming languages for programmable controllers. Annual Review in Automatic Programming, 13 (part 1):71–74, 1985. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/006641388590446X.

Clark:1963:LFP

[Cla63] K. W. Clark. ‘file processing’ in SEAL. Annual Review in Automatic Programming, 3:311–324, 1963. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/S0066413863800129.

Cleave:1960:AFT

[Cle60] J. P. Cleave. The application of formula translation to the automatic coding of ordinary differential equations. Annual Review in Automatic Programming, 1:81–92, 1960. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413860900343.

Clippinger:1961:FBC

[Cli61] R. F. Clippinger. FACT — a business compiler: Description and comparison with COBOL and commercial translator. Annual Review in Automatic Programming, 2:231–291, 1961. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/S0066413861800116.

Crespo:1992:RTE

[CNV^{+}92a] A. Crespo, J. L. Navarro, R. Vivó, A. García, and A. Espinosa. A real time expert system environment for process control. Annual Review in Automatic Programming, 16 (part 1):19–25, 1992.

CODEN ???? ISSN ???? URL https://www.sciencedirect.
com/science/article/pii/006641389190005V.

Crespo:1992:RES

[CNV^{+}92b] A. Crespo, J. L. Navarro, R. Vivó, A. García, and A. Espinosa. Rigas: an expert server task in real-time environments. Annual Review in Automatic Programming, 17:497–502, 1992. CO-

DEN ???? ISSN ???? URL https://www.sciencedirect.com/
science/article/pii/S0066413809910801.

Cord:1994:BAP

[Cor94] T. Cord. A blackboard application for process monitoring and supervision. Annual Review in Automatic Programming, 19:377– 382, 1994. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/0066413894900949.

Coulouris:1969:MIA

[Cou69] G. F. Coulouris. A machine-independent assembly language for systems programs. Annual Review in Automatic Programming, 6 (part 2):89–104, 1969. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413869900020.

Chen:1992:KRR

[CR92] J. Chen and M. G. Rodd. Knowledge representation for real-time on-line industrial inspection systems. Annual Review in Automatic Programming, 16 (part 1):115–120, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413891900198.

Crespo:1994:F

[Cre94] Alfons Crespo. Foreword. Annual Review in Automatic Programming, 19:vii, 1994. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/0066413894900353.

Cicin-Sain:1988:MMP

[ČŠ88] M. Čičin-Šain. Methods for monitoring productivity in applicative software development. Annual Review in Automatic Programming, 14 (part 2):59–62, 1988. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/006641389090011F.

Carvalho:1994:RTF

[CSDM94] M. T. Carvalho, J. M. Sousa, F. Durão, and P. M. Martins. Real time fuzzy control of column flotation process. Annual Review in Automatic Programming, 19:73–78, 1994. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413894900450.

Cassar:1992:MLV

[CSF92] J. P. Cassar, M. Staroswiecki, and R. Ferhati. Multivalued logic voting scheme for residual evaluation in failure detection and isolation systems. Annual Review in Automatic Programming, 17:267–272, 1992. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/S0066413809910448.

Cembrano:1994:NNR

[CTW94] G. Cembrano, C. Torras, and G. Wells. Neural networks for robot control. Annual Review in Automatic Programming, 19: 159–166, 1994. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/0066413894900590.

Cushing:1960:APB

[Cus60] Gordon Cushing. Automatic programming and business applications. Annual Review in Automatic Programming, 1:189– 195, 1960. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/0066413860900410.

Ceanga:1985:LSC

[CVB85] E. Ceangǎ, C. Vasiliu, and S. Bumbaru. Learning systems with control situations recognition. Annual Review in Automatic Programming, 12 (part 1):384–389, 1985. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413885900680.

Ciohocki:1985:ADI

[CW85] K. Ciohocki and W. Wojoiechowski. An attempt to dynamize an investment coefficient matrix in Leontieff models. Annual Review in Automatic Programming, 12 (part 1):151–157, 1985. CO-

DEN ???? ISSN ???? URL https://www.sciencedirect.com/
science/article/pii/0066413885900175.

Dabrowski:1985:MSR

[Dab̧85] Zdzisław Dab̧rowski. Modelling of the subject reality in the information system. Annual Review in Automatic Programming, 12 (part 1):328–331, 1985. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413885900552.

Dahme:1985:PMB

[Dah85] Christian Dahme. On the particularity of the model building of elementary social systems — a draft for a systems theory of elementary social systems. Annual Review in Automatic Programming, 12 (part 1):324–327, 1985. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413885900540.

DAgapeyeff:1963:PSC

[DBG63] A. D’Agapeyeff, H. D. Baecker, and B. J. Gibbens. Progress in some commercial source languages. Annual Review in Automatic Programming, 3:277–298, 1963. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/S0066413863800105.

Dominguez:1994:NNB

[DCA94] S. Dominguez, P. Campoy, and R. Aracil. A neural network based quality control system for steel strip manufacturing. Annual Review in Automatic Programming, 19:185–190, 1994. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413894900639.

Dekker:1985:MBP

[Dek85] L. Dekker. Methodology based parallel digital processors. Annual Review in Automatic Programming, 12 (part 1):38–50, 1985.

CODEN ???? ISSN ???? URL https://www.sciencedirect.
com/science/article/pii/0066413885900047.

Demidenko:1985:OCC

[Dem85] N. D. Demidenko. Optimal control of the complicated objects with distributed parameters. Annual Review in Automatic Programming, 12 (part 1):425–428, 1985. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413885900771.

Denison:1960:FDI

[Den60] S. J. M. Denison. Further DEUCE interpretative programs and some translating programs. Annual Review in Automatic Programming, 1:127–145, 1960. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413860900379.

Downes:1980:UAL

[DG80] V. A. Downes and S. J. Goldsack. The use of the Ada language for programming a distributed system. Annual Review in Automatic Programming, 10:39–44, ???? 1980. CO-DEN ARVAAM. ISSN 0066-4138 (print), 1878-545X (elec-

tronic).  URL https://www.sciencedirect.com/science/
article/pii/0066413882900064.

Davoli:1994:SAD

[DG94] R. Davoli and L. A. Giachini. A schedulability algorithm for data flow, hard-real-time distributed programs. Annual Review in Automatic Programming, 18:39–43, 1994. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413894900086.

Dao:1985:CCD

[DHD85] Dang Huu Dao, Hoang Minh Hai, and Phan Dinh Dieu. Controllability of a class of discrete bilinear systems. Annual Review in Automatic Programming, 12 (part 1):147–150, 1985. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413885900163.

DiPillo:1985:P

[Di 85] Gianni Di Pillo. Preface. Annual Review in Automatic Programming, 13 (part 2):vii, 1985. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413885904550.

DImperio:1969:DST

[D’I69] M. E. D’Imperio. Data structures and their representation in storage. Annual Review in Automatic Programming, 5 (part 1): 1–75, 1969. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/0066413869900366.

Deibert:1992:EFD

[DI92] R. Deibert and R. Isermann. Examples for fault detection in closed loops. Annual Review in Automatic Programming, 17: 235–240, 1992. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/S0066413809910394.

Dijkstra:1963:ATPa

[Dij63a] E. W. Dijkstra. Appendix: Two papers on an ALGOL translator for the X1. Annual Review in Automatic Programming, 3:327, 1963. CODEN ARVAAM. ISSN 0066-4138 (print), 1878-545X

(electronic). URL https://www.sciencedirect.com/science/
article/pii/B9780080097633500175;   https://www.sciencedirect.
com/science/article/pii/S0066413863800142.

Dijkstra:1963:DMI

[Dij63b] E. W. Dijkstra. On the design of machine independent programming languages. Annual Review in Automatic Programming, 3: 27–42, 1963. CODEN ARVAAM. ISSN 0066-4138 (print), 1878-

545X (electronic). URL https://www.sciencedirect.com/
science/article/pii/B9780080097633500072;   https://www.
sciencedirect.com/science/article/pii/S0066413863800038.

Originally issued as Report MR 34 of the Mathematics Centre, Amsterdam.

Dijkstra:1963:ATX

[Dij63c] Edsger W. Dijkstra. An ALGOL-60 translator for the X1. Annual Review in Automatic Programming, 3:329–345, 1963. CODEN ARVAAM. ISSN 0066-4138 (print), 1878-545X (electronic).

Dijkstra:1963:MTA

[Dij63d] Edsger W. Dijkstra. Making a translator for ALGOL-60. Annual Review in Automatic Programming, 3:347–356, 1963. CODEN ARVAAM. ISSN 0066-4138 (print), 1878-545X (electronic).

Dasai:1977:HLP

[DISN77] Takashi Dasai, Kunio Iharada, Takeshi Sunohara, and Toshiyuki Nakamura. High level process control language “ESPRIT” and its source level debugging system “SOLDA”. Annual Review in Automatic Programming, 8:19–28, 1977. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413877900052.

deJong:1985:ASQ

[dJM85] J. L. de Jong and K. C. P. Machielsen. On the application of sequential quadratic programming to state constrained optimal control problems. Annual Review in Automatic Programming, 13 (part 2):31–39, 1985. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413885904604.

Debou:1992:MSP

[DLP92] C. Debou, J. Lipták, and L. Pescoller. Managing software projects by applying AMI. Annual Review in Automatic Programming, 167 (part 2):103–108, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/006641389290016I.

Dovi:1985:RTN

[Dov85] V. G. Dovi. Restarting techniques in non linear programming algorithms for optimal chemical plant control. Annual Review in Automatic Programming, 13 (part 2):41–45, 1985. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413885904616.

Decotignie:1994:STC

[DR94] J.-D Decotignie and P. Raja. Spatio-temporal constraints in Fieldbus: Requirements and current solutions. Annual Review in Automatic Programming, 18:9–14, 1994. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413894900035.

Ditt:1988:PMS

[DS88] W. Ditt and U. Schlehofer. A project management support tool based on feedback loop technique. Annual Review in Automatic Programming, 14 (part 2):15–20, 1988. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/006641389090003A.

Debevc:1988:SDF

[DSD^{+}88] M. Debevc, R. Svečko, D. Donlagić, L. Bele-Tominc, and M. Venuti-Zorič. Software designing in the field of control engineering. Annual Review in Automatic Programming, 14 (part 2):41–44, 1988. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/006641389090008F.

Dubery:1985:MRR

[Dub85] J. M. Dubery. Modulas-1, -2, and -3(?) or a real real-time language. Annual Review in Automatic Programming, 13 (part 1): 81–89, ???? 1985. CODEN ARVAAM. ISSN 0066-4138 (print),

1878-545X (electronic). URL https://www.sciencedirect.
com/science/article/pii/0066413885904483.

Dubuisson:1992:GAL

[Dub92] B. Dubuisson. Genetic algorithms and learning. Annual Review in Automatic Programming, 17:299–302, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/S0066413809910497.

Duisburg:1985:GTA

[Dui85] F. Svaricek Duisburg. A graph-theoretic algorithm for computing the number of invariant zeros of large scale systems. Annual Review in Automatic Programming, 12 (part 1):276–279, 1985.

CODEN ???? ISSN ???? URL https://www.sciencedirect.
com/science/article/pii/0066413885900424.

Dummermuth:1994:RIF

[Dum94] E. Dummermuth. The real issues in fuzzy logic applications. Annual Review in Automatic Programming, 19:135–140, 1994. CO-

DEN ???? ISSN ???? URL https://www.sciencedirect.com/
science/article/pii/0066413894900558.

Espada:1994:RDF

[EB94] A. Espada and A. Barreiro. Robust design of fuzzy controllers based on small gain conditions. Annual Review in Automatic Programming, 19:55–60, 1994. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413894900426.

Efstathiou:1992:ATA

[Efs92] J. Efstathiou. Architectures and techniques of artificial intelligence in process control. Annual Review in Automatic Programming, 17:371–376, 1992. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/S0066413809910618.

Elzer:1992:P

[EH92] Peter F. Elzer and Volkmar Haase. Preface. Annual Review in Automatic Programming, 167 (part 2):vii, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413892900016.

E-hui:1989:AEF

[EhGhSl89] Xu E-hui, Xu Guo-hua, and Zhang Shi-liang. Application of expert fuzzy controller in the penicillin fermentation processes. Annual Review in Automatic Programming, 15 (part 2):97–100, 1989.

CODEN ???? ISSN ???? URL https://www.sciencedirect.
com/science/article/pii/0066413889900189.

Elzer:1980:RAM

[Elz80] P. F. Elzer. Resource allocation by means of access-rights, an alternative view on real-time programming? Annual Review in Automatic Programming, 10:73–79, 1980. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413882900106.

Engles:1972:TDB

[Eng72] Robert W. Engles. A tutorial on data-base organization. Annual Review in Automatic Programming, 7 (part 1):1, 3–64, 1972.

CODEN ???? ISSN ???? URL https://www.sciencedirect.
com/science/article/pii/0066413872900031.

Engelmann:1985:BCD

[Eng85] Bernd Engelmann. A basic concept for dual optimization methods for complex interconnected systems. Annual Review in Automatic Programming, 12 (part 1):288–290, 1985. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/006641388590045X.

Edwards:1977:UOC

[ES77] T. R. M. Edwards and P. E. Strickland. A user oriented control system for a magnetic field mapping device. Annual Review in Automatic Programming, 8:191–197, 1977. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413877900258.

Ester:1985:ASM

[Est85] Jochen Ester. About some methods and applications in multicriteria decision making. Annual Review in Automatic Programming, 12 (part 1):174–176, 1985. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413885900217.

Evans:1964:AC

[Eva64] Arthur Evans. An ALGOL 60 compiler. Annual Review in Automatic Programming, 4:87–124, January 1964. CO-DEN ARVAAM. ISSN 0066-4138 (print), 1878-545X (elec-

tronic).  URL https://www.sciencedirect.com/science/
article/pii/0066413864900151.

Fandel:1985:CAG

[Fan85] Günter Fandel. Cn the applicability of group decision making concepts to wage bargaining. Annual Review in Automatic Programming, 12 (part 1):62–68, 1985. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413885900060.

Fan:1994:MSR

[Fan94] Changpeng Fan. MMOSS: Soft real-time operating system support in a multimedia communication subsystem. Annual Review in Automatic Programming, 18:127–131, 1994. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413894900221.

French:1992:ITP

[FCWM92] I. French, C. Cox, M. J. Willis, and G. A. Montague. Intelligent tuning of P+I controllers for bioprocess application. Annual Review in Automatic Programming, 17:403–408, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/S0066413809910667.

Fedra:1985:IWR

[Fed85] Kurt Fedra. Interactive water resources planning and management — computer simulation with a friendly user-interface. Annual Review in Automatic Programming, 12 (part 1):490– 492, 1985. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/0066413885900941.

Felton:1960:AIC

[Fel60] G. E. Felton. Assembly, interpretive and conversion programs for PEGASUS. Annual Review in Automatic Programming, 1: 32–57, 1960. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/0066413860900318.

Fritz:1988:USS

[FHK88] W. Fritz, V. Haase, and R. Kalcher. The use of standard software in real time programming: An example demonstrating the integration of ADA, ORACLE and GKS. Annual Review in Automatic Programming, 14 (part 1):43–49, ???? 1988. CODEN ARVAAM. ISSN 0066-4138 (print), 1878-545X

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0066413888900079.

Findler:1992:DLR

[Fin92] N. V. Findler. Dimensions of learning in a real-time knowledgebased control system. Annual Review in Automatic Programming, 17:159–163, 1992. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/S0066413809910278.

Fischer:1985:CTS

[Fis85] A. M. Fischer. The computer — a tool for statistical process control. Annual Review in Automatic Programming, 13 (part 1):1–7, 1985. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/0066413885904367.

Foss:1992:PND

[FJ92] B. A. Foss and T. A. Johansen. Parallel nonlinear decoupling for process control — a NARMAX approach. Annual Review in Automatic Programming, 17:49–54, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/S0066413809910102.

Fukuoka:1981:PFB

[FKKK81] K. Fukuoka, M. Kondo, Y. Kubo, and T. Kawamura. Program fall-back method in real time systems. Annual Review in Automatic Programming, 11:57–62, 1981. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413881900094.

Fleming:1985:CAD

[Fle85] P. J. Fleming. Computer aided design of regulators using multiobjective optimization. Annual Review in Automatic Programming, 13 (part 2):47–52, 1985. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413885904628.

Fringuelli:1992:UAA

[FLMS92] B. Fringuelli, E. Lamma, P. Mello, and G. Santocchia. Using ATMS algorithms for real-time applications. Annual Review in Automatic Programming, 16 (part 1):81–86, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413891900132.

Fortier:1994:CRT

[FP94] Paul J. Fortier and Janet Prichard. Concept for a real-time structured database query language (RT-SQL). Annual Review in Automatic Programming, 18:167–174, 1994. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413894900299.

Fjellheim:1994:AMR

[FPCN94] Roar Arne Fjellheim, Thomas Bech Pettersen, Birger Christoffersen, and Andy Nicholls. Application methodology for REAKT1 systems. Annual Review in Automatic Programming, 19:325– 332, 1994. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/0066413894900868.

Fang:1990:DMD

[FQ90] Bin Fang and Peng Tian Qiang. A design model of distributed database in real-time control. Annual Review in Automatic Programming, 15 (part 1):53–56, 1990. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/006641389090027O.

Fernandez:1988:LBD

[FQVM88] A. Fernández, J. Quemada, L. Vidaller, and C. Miguel. LOTOS based derivation methodology in the PRODAT project. Annual Review in Automatic Programming, 14 (part 1):7–13, 1988. CO-

DEN ???? ISSN ???? URL https://www.sciencedirect.com/
science/article/pii/006641388890002X.

Frakes:1992:SRE

[Fra92a] William B. Frakes. Software reuse: an empirical approach. Annual Review in Automatic Programming, 167 (part 2):41–44, 1992.

CODEN ???? ISSN ???? URL https://www.sciencedirect.
com/science/article/pii/006641389290007C.

Frank:1992:PMB

[Fra92b] P. M. Frank. Principles of model-based fault detection. Annual Review in Automatic Programming, 17:213–220, 1992. CO-

DEN ???? ISSN ???? URL https://www.sciencedirect.com/
science/article/pii/S0066413809910369.

Fried:1964:ICP

[Fri64] David C. Fried. An ideal computer program support system and a specific IBM system. Annual Review in Automatic Programming, 4:193–215, 1964. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413864900199.

Fritz:1992:UPM

[Fri92] W. Fritz. Using the project management tool Microsoft Project for planning and controlling software development projects. Annual Review in Automatic Programming, 167 (part 2):35–40, 1992.

CODEN ???? ISSN ???? URL https://www.sciencedirect.
com/science/article/pii/006641389290006B.

Fatikow:1994:NBL

[FS94] S. Fatikow and K. Sundermann. Neural-based learning in grasp force control of a robot hand. Annual Review in Automatic Programming, 19:111–116, 1994. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413894900515.

Gafni:1988:MHR

[Gaf88] V. Gafni. A model for a hard real time system executive. Annual Review in Automatic Programming, 14 (part 1):69–74, 1988.

CODEN ???? ISSN ???? URL https://www.sciencedirect.
com/science/article/pii/0066413888900110.

Garcia-Cerezo:1992:SFC

[GCOA92] A. García-Cerezo, A. Ollero, and J. Aracil. Stability of fuzzy control systems by using nonlinear system theory. Annual Review in Automatic Programming, 17:121–126, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/S0066413809910217.

Garcia-Cerezo:1994:DAW

[GCOA94] A. García-Cerezo, A. Ollero, and J. Aracil. Dynamic analysis of weighted-output fuzzy control systems. Annual Review in Automatic Programming, 19:43–48, 1994. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/006641389490040X.

Gerhardt:1985:CNM

[Ger85] Hans-Detlef Gerhardt. Combi nets and mathematical modelling. Annual Review in Automatic Programming, 12 (part 1):332– 335, 1985. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/0066413885900564.

Gliss:1990:SIC

[GG90] B. Gliss and J. Gorski. Specification and implementation of concurrency in a simple database system. Annual Review in Automatic Programming, 15 (part 1):23–28, 1990. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/006641389090022J.

Gnevko:1985:AMS

[GGG85] S. V. Gnevko, V. S. Glushenkov, and A. V. Guminsky. Application of methods for solving special nonlinear programming problems to optimization of control systems. Annual Review in Automatic Programming, 13 (part 2):59–64, 1985. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413885904641.

Gill:1960:PP

[Gil60] Stanley Gill. The philosophy of programming. Annual Review in Automatic Programming, 1:178–188, 1960. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413860900409.

Gchonefeld:1985:SMM

[GJ85] R. Gchönefeld and H. Jacobi. Stochastic model of a multiprocessor system in the presence of memory contention. Annual Review in Automatic Programming, 12 (part 1):449–452, 1985.

CODEN ???? ISSN ???? URL https://www.sciencedirect.
com/science/article/pii/0066413885900837.

Gorski:1977:RTC

[GK77] Janusz Górski and Andrzej Kunicki. Real-time computer system for the data acquisition net control. Annual Review in Automatic Programming, 8:181–190, 1977. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413877900246.

Goldsack:1980:UIA

[GK80] S. J. Goldsack and J. Kramer. The use of invariants in the application — oriented specification of real time control systems. Annual Review in Automatic Programming, 10:63–71, 1980. CO-

DEN ???? ISSN ???? URL https://www.sciencedirect.com/
science/article/pii/006641388290009X.

Gabasov:1985:OCP

[GKKP85] R. Gabasov, F. M. Kirillova, O. I. Kostyukova, and A. V. Pokatayev. Optimal control problems treated with algorithms of linear and nonlinear programming. Annual Review in Automatic Programming, 13 (part 2):53–58, 1985. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/006641388590463X.

Gander:1980:RTP

[GL80] J.-G Gander and H.-U Liechti. Real-time process control based on a high level state language. Annual Review in Automatic Programming, 10:91–94, 1980. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/006641388290012X.

Glennie:1960:FTA

[Gle60] A. E. Glennie. Future trends in automatic programming. Annual Review in Automatic Programming, 1:8–15, 1960. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413860900288.

Garetti:1981:DDO

[GLR81] P. Garetti, P. Laface, and S. Rivoira. On the development of a distributed operating system kernel for real-time applications. Annual Review in Automatic Programming, 11:107–114, 1981. CO-

DEN ???? ISSN ???? URL https://www.sciencedirect.com/
science/article/pii/0066413881900161.

Garcia:1988:SDM

[GLTD88] D. Garcia, H. Lopez, J. Tuya, and A. Diez. Specification, design and Modula-2 implementation of a low cost industrial control system. Annual Review in Automatic Programming, 14 (part 1):51–56, 1988. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413888900080.

Gu:1985:OPA

[GM85] D. W. Gu and D. Q. Mayne. Optimization problems arising in control systems design. Annual Review in Automatic Programming, 13 (part 2):65–69, 1985. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413885904653.

Guasch:1994:RLL

[GMS94] A. Guasch, R. Milne, and R. Sarrate. Relay ladder logic diagnosis. Annual Review in Automatic Programming, 19:275–280, 1994.

CODEN ???? ISSN ???? URL https://www.sciencedirect.
com/science/article/pii/0066413894900787.

Glassey:1992:EFD

[GMWK92] J. Glassey, G. A. Montague, A. C. Ward, and B. Kara. Enhancing fermentation development procedures via artificial neural networks. Annual Review in Automatic Programming, 17: 435–440, 1992. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/S0066413809910710.

Goodman:1960:P

[Goo60] Richard Goodman. Preface. Annual Review in Automatic Programming, 1:ix–xi, 1960. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413860900264.

Goodman:1961:P

[Goo61] Richard Goodman. Preface. Annual Review in Automatic Programming, 2:vii–viii, 1961. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/S0066413861800013.

Goodman:1964:P

[Goo64] Richard Goodman. Preface. Annual Review in Automatic Programming, 4:v, 1964. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413864900126.

Gorski:1980:KKW

[Gór80] J. Górski. KRTM — a kernel which supports real time multiprograms. Annual Review in Automatic Programming, 10:9–18, 1980.

CODEN ???? ISSN ???? URL https://www.sciencedirect.
com/science/article/pii/0066413882900039.

Garcia:1994:LTA

[GP94] R. Ferreiro Garcia and J. Vidal Paz. Learning task applied to identification of a marine vehicle. Annual Review in Automatic Programming, 19:197–201, 1994. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413894900655.

Greenspan:1985:PMS

[Gre85] Donald Greenspan. Particle modelling of solids, liquids and gases with application to a natural self-reorganization. Annual Review in Automatic Programming, 12 (part 1):29–37, 1985. CO-

DEN ???? ISSN ???? URL https://www.sciencedirect.com/
science/article/pii/0066413885900035.

Gomez:1994:RTE

[GRSA94] P. Gomez, S. Romero, P. Serrahima, and I. Alarcon. A real time expert system for continuous assistance in process control: a successful approach. Annual Review in Automatic Programming, 19:371–375, 1994. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413894900930.

Gyorkos:1988:CEC

[GRW88] J. Györkös, I. Rozman, and T. Welzer. The concept of an efficient computer aided software engineering tool. Annual Review in Automatic Programming, 14 (part 2):89–94, 1988. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/006641389090016K.

Grzech:1985:FCV

[Grz85] Adam Grzech. Flow control in virtual circuit-oriented packetswitched computer communication network. Annual Review in Automatic Programming, 12 (part 1):437–440, 1985. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413885900801.

Giron-Sierra:1994:AOO

[GSGPAT94] J. M. Giron-Sierra, J. A. Gomez-Pulido, and B. Andres-Toro. Application of an object-oriented expert system shell to a fermentation process. Annual Review in Automatic Programming, 19:237–242, 1994. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413894900728.

Garbajosa:1994:NLF

[GTW94] J. Garbajosa, O. Tejedor, and M. Wolff. Natural language front end to test systems. Annual Review in Automatic Programming, 19:261–267, 1994. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413894900760.

Gutenbaum:1985:SPO

[Gut85] Jakub Gutenbaum. Smoothing of the Pareto optimal solutions set in linear multiobjective optimization problems. Annual Review in Automatic Programming, 12 (part 1):230–235, 1985. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413885900321.

Haase:1980:P

[Haa80] V. H. Haase. Preface. Annual Review in Automatic Programming, 10:vii, 1980. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/0066413882900015.

Halling:1977:STI

[Hal77] Horst Halling. Steps towards the implementation of a parallel code executor. Annual Review in Automatic Programming, 8:55–66, 1977. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/006641387790009X.

Halling:1981:RAF

[Hal81] H. Halling. Recent advances and future trends in distributed control systems. Annual Review in Automatic Programming, 11: 1–13, 1981. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/0066413881900021.

Halang:1994:P

[Hal94] Wolfgang A. Halang. Preface. Annual Review in Automatic Programming, 18:vii, 1994. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413894900019.

Hanford:1960:SOS

[Han60] K. V. Hanford. The share operating system for the IBM 709. Annual Review in Automatic Programming, 1:169–177, 1960. CO-

DEN ???? ISSN ???? URL https://www.sciencedirect.com/
science/article/pii/0066413860900392.

Hanzalek:1994:RTN

[Han94] Z. Hanzálek. Real-time neural controller implemented on parallel architecture. Annual Review in Automatic Programming, 19: 227–232, 1994. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/0066413894900701.

Harris:1989:DEI

[Har89] C. J. Harris. Distributed estimation, inferencing and multi-sensor data fusion for real time supervisory control. Annual Review in Automatic Programming, 15 (part 2):19–24, 1989. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413889900049.

Haramis:1992:IFS

[Har92] G. E. Haramis. Implementing a feasibility study “a procedural approach”. Annual Review in Automatic Programming, 167 (part 2):133–138, 1992. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/006641389290021G.

Hesse:1992:RIP

[HBS92] W. Hesse, U. Bittner, and J. Schnath. Results from the IPAS project: Influences of methods and tools, quality requirements and project management on the work situation of software developers. Annual Review in Automatic Programming, 167 (part 2):179–183, 1992. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/006641389290027M.

Hurteau:1985:WSC

[HD85] R. Hurteau and R. M. DeSantis. A work station for computer assisted design of real-time control systems. Annual Review in Automatic Programming, 13 (part 1):75–80, 1985. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413885904471.

Horvai:1985:SAL

[HGH^{+}85] M. Horvai, L. Gyimesi, A. Horváth, G. Juhász, K. Kovács, I. Sári, J. Szlankó, and E. Tóth. Software architecture of large distributed process control systems. Annual Review in Automatic Programming, 13 (part 1):43–48, 1985. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413885904422.

Herget:1992:AIT

[HGM92] C. J. Herget, E. L. Grasz, and R. D. Merrill. Applications of intelligent telerobotic control. Annual Review in Automatic Programming, 16 (part 1):149–152, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413891900257.

Herbstreith:1992:TDN

[HGP92] H. Herbstreith, L. Gmeiner, and P. Preuß. A targetdirected neurally controlled vehicle. Annual Review in Automatic Programming, 17:67–71, 1992. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/S0066413809910138.

Hawkins:1963:MPT

[HH63] E. N. Hawkins and D. H. R. Huxtable. A multi-pass translation scheme for ALGOL 60. Annual Review in Automatic Programming, 3:163–205, January 1963. CODEN ARVAAM. ISSN 0066-4138 (print), 1878-545X (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0066413863800075.

Halang:1988:APL

[HH88] W. A. Halang and R. Henn. Additional pearl language structures for the implementation of reliable and inherently safe real-time systems. Annual Review in Automatic Programming, 14 (part 1):35–42, 1988. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413888900067.

Heydeman:1992:ASI

[HH92a] J. Heydeman and G. Honderd. Adaptive and supplementary intelligent control of power system stabilizers. Annual Review in Automatic Programming, 17:177–182, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/S0066413809910308.

Hunt:1992:ICR

[HH92b] K. J. Hunt and Y. M. Han. Induction of control rules from human skill. Annual Review in Automatic Programming, 17:97–102, 1992.

CODEN ???? ISSN ???? URL https://www.sciencedirect.
com/science/article/pii/S0066413809910175.

Higman:1963:TAT

[Hig63] B. Higman. Towards an ALGOL translator. Annual Review in Automatic Programming, 3:121–162, January 1963. CO-DEN ARVAAM. ISSN 0066-4138 (print), 1878-545X (elec-

tronic).  URL https://www.sciencedirect.com/science/
article/pii/S0066413863800063.

Hinchey:1994:VMR

[Hin94] Michael G. Hinchey. Visual methods in real-time programming. Annual Review in Automatic Programming, 18:157–162, 1994.

CODEN ???? ISSN ???? URL https://www.sciencedirect.
com/science/article/pii/0066413894900272.

Hanford:1973:DSC

[HJ73] K. V. Hanford and C. B. Jones. Dynamic syntax: a concept for the definition of the syntax of programming languages. Annual Review in Automatic Programming, 7 (part 2):115–140, IN1–IN3, 141–142, 1973. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413873900037.

Hoyer:1994:AFC

[HJ94] R. Hoyer and U. Jumar. An advanced fuzzy controller for traffic lights. Annual Review in Automatic Programming, 19:67– 72, 1994. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/0066413894900442.

Horton:1977:RTD

[HJLT77] J. R. Horton, I. R. James, A. Langsford, and W. Teasdale. A real-time distributed processing system using GEC 4000 series computers. Annual Review in Automatic Programming, 8:91, 93–97, 1977. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413877900131.

Huzar:1985:SAS

[HK85] Z. Huzar and K. Koleśnik. A systematic approach to a simulation model construction. Annual Review in Automatic Programming, 12 (part 1):318–323, 1985. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413885900539.

Hazeyama:1992:PMS

[HK92] A. Hazeyama and S. Komiya. A process model for software project management. Annual Review in Automatic Programming, 167 (part 2):109–116, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/006641389290017J.

Haarman:1977:SAM

[HKM77] J. C. Haarman, H. U. Koller, and J. A. Muheim. Software approach for multi-processor systems. Annual Review in Automatic Programming, 8:107–117, 1977. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413877900155.

Hang:1992:IKB

[HLH92] C. C. Hang, T. H. Lee, and W. K. Ho. Implementation of a knowledge-based PID auto-tuner. Annual Review in Automatic Programming, 17:153–158, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/S0066413809910266.

Hooman:1992:TDE

[Hoo92] J. Hooman. Top-down design of embedded real-time AI systems. Annual Review in Automatic Programming, 17:453–458, 1992.

CODEN ???? ISSN ???? URL https://www.sciencedirect.
com/science/article/pii/S0066413809910746.

Howell:1992:KBD

[How92] J. Howell. Knowledge-based diagnosis in information poor plants: a materials accountancy application. Annual Review in Automatic Programming, 17:273–278, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/S006641380991045X.

Hanmandlu:1985:UNP

[HPP85] M. Hanmandlu, P. Purkayastha, and J. K. Pal. On the use of nonlinear programming in real time control in process industries. Annual Review in Automatic Programming, 13 (part 2):71–78, 1985.

CODEN ???? ISSN ???? URL https://www.sciencedirect.
com/science/article/pii/0066413885904665.

Hayes-Roth:1994:IRT

[HR94] Barbara Hayes-Roth. Integrating real-time AI techniques in adaptive intelligent agents. Annual Review in Automatic Programming, 19:1–11, 1994. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413894900361.

Hasiewicz:1985:GII

[HS85] Zygmunt Hasiewicz and Anna Stankiewicz. Global identification of interconnected steady-state systems by interaction balance method. Annual Review in Automatic Programming, 12 (part 1):267–271, 1985. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413885900400.

Hubner:1992:MBD

[Hüb92] T. Hübner. Model-based diagnosis of complex systems: Structuring and temporal reasoning. Annual Review in Automatic Programming, 16 (part 1):41–45, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/006641389190008Y.

Hughes:1992:UPP

[Hug92] R. T. Hughes. Using practitioners’ problems to shape the content of a course on software project management. Annual Review in Automatic Programming, 167 (part 2):143–152, 1992. CO-

DEN ???? ISSN ???? URL https://www.sciencedirect.com/
science/article/pii/006641389290023I.

Humby:1960:TCC

[Hum60] E. Humby. TIDE: a commercial compiler for the IBM 650. Annual Review in Automatic Programming, 1:207– 216, IN1–IN2, 217–219, 1960. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413860900434.

Humby:1963:R

[Hum63] E. Humby. Rapidwrite. Annual Review in Automatic Programming, 3:299–309, 1963. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/S0066413863800117.

Hiraiwa:1992:EPR

[HUS92] A. Hiraiwa, N. Uchida, and K. Shimohara. EMG pattern recognition by neural networks for prosthetic fingers control. Annual Review in Automatic Programming, 17:73–79, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/S006641380991014X.

Hilhorst:1992:SCM

[HvALT92] R. A. Hilhorst, J. van Amerongen, P. Löhnberg, and H. J. A. F. Tulleken. Supervisory control of mode-switch processes: Application to a flexible beam. Annual Review in Automatic Programming, 17:195–200, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/S0066413809910333.

Holden:1977:EPL

[HW77] J. Holden and I. C. Wand. Experience with the programming language MODULA. Annual Review in Automatic Programming, 8:3–11, 1977. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413877900039.

Ignova:1994:KII

[IGM^{+}94] M. Ignova, J. Glassey, G. A. Montague, A. C. Ward, and A. J. Morris. Knowledge integration for improved bioprocess supervision. Annual Review in Automatic Programming, 19:269– 273, 1994. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/0066413894900779.

Iliffe:1961:UGS

[Ili61] J. K. Iliffe. The use of the genie system in numerical calculation. Annual Review in Automatic Programming, 2:1–28, 1961.

CODEN ???? ISSN ???? URL https://www.sciencedirect.
com/science/article/pii/S0066413861800025.

Ichbiah:1972:GCS

[IM72] J. D. Ichbiah and S. P. Morse. General concepts of the Simula 67 programming language. Annual Review in Automatic Programming, 7 (part 1):65–93, 1972. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413872900043.

Ivachnenko:1985:CSO

[IM85] A. G. Ivachnenko and J. A. Müller. The contribution of selforganization to experimental system analysis. Annual Review in Automatic Programming, 12 (part 1):314–317, 1985. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413885900527.

Inoue:1981:CPS

[IMF^{+}81] K. Inoue, H. Mitsuoka, S. Fujii, J. Yamashita, and J. Matsuki. Centralised programming system for process computer at Kimitsu works of Nippon Steel Corporation. Annual Review in Automatic Programming, 11:77–84, 1981. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413881900124.

Irons:1963:SUS

[Iro63] Edgar T. Irons. The structure and use of the syntax directed compiler. Annual Review in Automatic Programming, 3:207–227, January 1963. CODEN ARVAAM. ISSN 0066-4138 (print),

1878-545X (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0066413863800087.

Johannsen:1994:HIP

[JA94] G. Johannsen and E. A. Averbukh. Human interaction for process supervision based on end-user knowledge and participation. Annual Review in Automatic Programming, 19:255–260, 1994. CO-

DEN ???? ISSN ???? URL https://www.sciencedirect.com/
science/article/pii/0066413894900752.

Jennings:1992:UGB

[Jen92] N. R. Jennings. Using grate to build cooperating agents for industrial control. Annual Review in Automatic Programming, 17: 397–402, 1992. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/S0066413809910655.

James:1992:STD

[JH92] J. R. James and C. J. Herget. Software tools for distributed intelligent control systems. Annual Review in Automatic Programming, 16 (part 1):87–90, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413891900143.

Jiahua:1990:RTD

[JHJ90] Wang Jiahua, Zheng Huaiyuan, and Zhu Jiakeng. A real-time distributed database system in cold tandem mill. Annual Review in Automatic Programming, 15 (part 1):83–86, 1990. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/006641389090032M.

Jansen:1985:PIC

[JK85] W. Jansen and E. U. Kriegel. Parameter identification in chaotic systems. Annual Review in Automatic Programming, 12 (part 1):162–169, 1985. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413885900199.

Jakubowski:1985:ISN

[JKW85] A. Jakubowski, R. Kulikowski, and D. Wagner. Interactive system for negotiations on allocation of research funds. Annual Review in Automatic Programming, 12 (part 1):202–207, 1985. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/006641388590028X.

Janning:1988:IPT

[JL88] T. Janning and C. Lewerentz. Integrated project team management in a software development environment. Annual Review in Automatic Programming, 14 (part 2):9–14, 1988. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413890900029.

Janiak:1985:GLT

[JS85] Adam Janiak and Anna Stankiewicz. Global and local timeoptimal resource allocation in a sequence of projects. Annual Review in Automatic Programming, 12 (part 1):408–413, 1985.

CODEN ???? ISSN ???? URL https://www.sciencedirect.
com/science/article/pii/0066413885900734.

Jalel:1992:AIC

[JST^{+}92] N. A. Jalel, F. Shui, D. Tsaptsinos, R. Tang, W. Vanichsriratana, and J. R. Leigh. Artificial intelligence in the control of a class of fermentation processes. Annual Review in Automatic Programming, 17:441–446, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/S0066413809910722.

Jager:1992:FIR

[JVB92] R. Jager, H. B. Verbruggen, and P. M. Bruijn. Fuzzy inference in rule-based real-time control. Annual Review in Automatic Programming, 17:183–188, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/S006641380991031X.

Jinfu:1992:PCS

[JX92] Sun Jinfu and Gao Xiaoqiang. Process control system for blast furnace based on artificial intelligence. Annual Review in Automatic Programming, 16 (part 1):159–163, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413891900279.

Kalashnikov:1985:PMT

[Kal85] V. V. Kalashnikov. Problems of methodology and technology for systems modelling and simulation. Annual Review in Automatic Programming, 12 (part 1):311–313, 1985. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413885900515.

Kappatsch:1977:FAP

[Kap77] Axel Kappatsch. A flexible approach to process communication. Annual Review in Automatic Programming, 8:13–18, 1977. CO-

DEN ???? ISSN ???? URL https://www.sciencedirect.com/
science/article/pii/0066413877900040.

Karr:1992:ASP

[Kar92] C. L. Karr. An adaptive system for process control using genetic algorithms. Annual Review in Automatic Programming, 17: 329–334, 1992. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/S0066413809910540.

Kindler:1985:UTP

[KCP85] Eugene Kindler, Stephen Chochol, and Charles Prokop. Universal transport processes. Annual Review in Automatic Programming, 12 (part 1):357–360, 1985. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413885900618.

Kellermayr:1985:SLA

[Kel85] Karl H. Kellermayr. Simulation of local area computer networks. Annual Review in Automatic Programming, 12 (part 1): 433–436, 1985. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/0066413885900795.

Kaiser:1994:IAE

[KGN94] M. Kaiser, A. Giordana, and M. Nuttin. Integrated acquisition, execution, evaluation, and tuning of elementary skills for intelligent robots. Annual Review in Automatic Programming, 19:117–122, 1994. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413894900523.

Kudoh:1981:EPC

[KHM81] T. Kudoh, M. Hirayama, and K. Mikami. Evaluation of pipelined common processor in distributed processing system. Annual Review in Automatic Programming, 11:121–127, 1981. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413881900185.

Kleinmann:1992:IRT

[KHP92] K. Kleinmann, M. Hormel, and W. Paetsch. Intelligent realtime control of a multifingered robot gripper by learning incremental actions. Annual Review in Automatic Programming, 17: 303–309, 1992. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/S0066413809910503.

Kienegger:1992:DWE

[Kie92] E. H. Kienegger. Development of a workstation environment for the assessment of a stormwater drainage fee. Annual Review in Automatic Programming, 167 (part 2):21–27, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413892900049.

Kiwiel:1985:ALC

[Kiw85] Krzysztof C. Kiwiel. An algorithm for linearly constrained nonsmooth convex multiobjective minimization. Annual Review in Automatic Programming, 12 (part 1):236–238, 1985. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413885900333.

Krijgsman:1992:DRT

[KJ92] A. J. Krijgsman and R. Jager. Development of real-time AI systems. Annual Review in Automatic Programming, 17:503– 507, 1992. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/S0066413809910813.

Krijgsman:1992:DFI

[KJVB92] A. J. Krijgsman, R. Jager, H. B. Verbruggen, and P. M. Bruijn. Dice: a framework for intelligent real-time control. Annual Review in Automatic Programming, 16 (part 1):13–18, 1992. CO-

DEN ???? ISSN ???? URL https://www.sciencedirect.com/
science/article/pii/006641389190004U.

Kapus-Kolar:1988:CLC

[KK88] M. Kapus-Kolar. Construction of logically correct communication protocols without over-specification. Annual Review in Automatic Programming, 14 (part 2):83–88, 1988. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/006641389090015J.

Karakas:1992:PMC

[KKK92] Gy Karakas, G. Kazai, and E. Knuth. Project management by configuration inheritance. Annual Review in Automatic Programming, 167 (part 2):29–33, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/006641389290005A.

Kouba:1992:SDC

[KLM^{+}92] Z. Kouba, J. Lažansky, V. Marik, G. Quirchmayr, W. Retschitzegger, T. Vlcek, and R. Wagner. Software development for CIM — a case study. Annual Review in Automatic Programming, 167 (part 2):11–19, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413892900038.

Kuwahara:1981:RAA

[KM81] H. Kuwahara and T. Miura. Recent advances of application and technology of industrial computers in Japan. Annual Review in Automatic Programming, 11:15–19, 1981. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413881900033.

Kersual:1994:RRT

[KM94] D. Kersual and A. Mensch. REAKT: a real time architecture for knowledge based systems. Annual Review in Automatic Programming, 19:319–324, 1994. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/006641389490085X.

Kupferschmid:1985:EAN

[KMEK85] M. Kupferschmid, K. Mohrmann, J. G. Ecker, and H. Kaufman. The ellipsoid algorithm: a new method for feedback gain optimization. Annual Review in Automatic Programming, 13 (part 2):85–94, 1985. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413885904689.

Kobayashi:1992:MIR

[KMO^{+}92] K. Kobayashi, R. Murai, H. Ogata, M. Takegaki, and T. Ishioka. A method of inducing rules for fuzzy controllers. Annual Review in Automatic Programming, 16 (part 1):133–138, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413891900224.

Kaldma:1994:SAT

[KMP^{+}94] T. Kaldma, L. Motus, R. Paluoja, J. Tekko, and T. Tommingas. Specification and analysis of timing properties in a case tool CON-RAD. Annual Review in Automatic Programming, 18:67–72, 1994.

CODEN ???? ISSN ???? URL https://www.sciencedirect.
com/science/article/pii/0066413894900132.

Knuth:1990:P

[Knu90] Elod Knuth. Preface. Annual Review in Automatic Programming, 15 (part 1):vii, 1990. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/006641389090018M.

Koch:1988:MME

[Koc88] W. H. Koch. Methodical and management experiences from an extensive software project. Annual Review in Automatic Programming, 14 (part 2):31–36, 1988. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/006641389090006D.

Kopf:1992:DRS

[Kop92] M. J. Kopf. Development of reusable software for business information systems. Annual Review in Automatic Programming, 167 (part 2):85–90, 1992. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/006641389290013F.

Kappatsch:1980:LUM

[KPR^{+}80] A. Kappatsch, F. Prunet, J. Robert, S. Savoysky, G. Verroust, and P. Wegmann. The LTPL-e universal I/O model for realtime programming. Annual Review in Automatic Programming, 10:19–28, 1980. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/0066413882900040.

Kraft:1985:OCH

[Kra85] D. Kraft. Optimal control of a high performance wind tunnel. Annual Review in Automatic Programming, 13 (part 2):79–83, 1985.

CODEN ???? ISSN ???? URL https://www.sciencedirect.
com/science/article/pii/0066413885904677.

Koivisto:1992:PNN

[KRK92] H. Koivisto, V. Ruoppila, and H. N. Koivo. Properties of the neural network internal model controller. Annual Review in Automatic Programming, 17:61–66, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/S0066413809910126.

Krystyna:1985:EOL

[Kry85] Koleśnik Krystyna. Events oriented language Pascalism. Annual Review in Automatic Programming, 12 (part 1):476–478, 1985.

CODEN ???? ISSN ???? URL https://www.sciencedirect.
com/science/article/pii/0066413885900904.

Kemp:1994:CCC

[KSK94] B. Kemp, P. Soper, and A. King. Continuation compilation for concurrent logic programming. Annual Review in Automatic Programming, 19:299–305, 1994. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413894900825.

Kokol:1992:MAI

[KSŽ92] P. Kokol, B. Stiglic, and V. Žumer. Metaparadigm’s adventures: Incorporating design situations in software process design. Annual Review in Automatic Programming, 167 (part 2):117–120, 1992.

CODEN ???? ISSN ???? URL https://www.sciencedirect.
com/science/article/pii/006641389290018K.

Kalaikov:1992:LCE

[KTF92] I. Kalaikov, I. Todorov, and N. Fidanov. Low-cost environment for analysis and design of knowledge-based control algorithms. Annual Review in Automatic Programming, 17:509–513, 1992. CO-

DEN ???? ISSN ???? URL https://www.sciencedirect.com/
science/article/pii/S0066413809910825.

Kunovsky:1985:NIP

[Kun85] Jiřĭ Kunovský. Numerical integration and parallel microprocessors systems. Annual Review in Automatic Programming, 12 (part 1):453–454, 1985. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413885900849.

Kuvaja:1992:ECI

[Kuv92] P. Kuvaja. Effects of case implementation on the productivity and maturity level of software development — a case study in a large industrial company. Annual Review in Automatic Programming, 167 (part 2):161–170, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/006641389290025K.

Kuzmuk:1985:CPS

[Kuz85] V. V. Kuzmuk. Concurrent process simulation by means op modified Petri nets. Annual Review in Automatic Programming, 12 (part 1):460–463, 1985. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413885900862.

Krose:1992:LAC

[KvD92] B. J. A. Kröse and J. W. M. van Dam. Learning to avoid collisions: a reinforcement learning paradigm for mobile robot navigation. Annual Review in Automatic Programming, 17:317–321, 1992.

CODEN ???? ISSN ???? URL https://www.sciencedirect.
com/science/article/pii/S0066413809910527.

Kokol:1988:CCC

[KZ88] P. Kokol and V. Zumer. Controlling complexity and cost of software projects with a spreadsheet software. Annual Review in Automatic Programming, 14 (part 2):75–82, 1988. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/006641389090014I.

Linkens:1992:RTS

[LA92] D. A. Linkens and M. F. Abbod. Real-time supervisory control for industrial processes. Annual Review in Automatic Programming, 17:377–382, 1992. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/S006641380991062X.

Larsson:1992:MBA

[Lar92] J. E. Larsson. Model-based alarm analysis using MFM. Annual Review in Automatic Programming, 16 (part 1):121–126, 1992.

CODEN ???? ISSN ???? URL https://www.sciencedirect.
com/science/article/pii/006641389190020C.

Lopes:1994:LDF

[LCM94] L. Seabra Lopes and L. M. Camarinha-Matos. Learning to diagnose failures of assembly tasks. Annual Review in Automatic Programming, 19:97–103, 1994. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413894900493.

leCalvez:1977:CGG

[lCMM77] F. le Calvez, F. Madaule, and H. G. Mendelbaum. Compiling GAELIC: a global real time language. Annual Review in Automatic Programming, 8:35, 37–45, 1977. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413877900076.

Litt:1985:NPI

[LD85] F.-X Litt and J. Delcommune. Numerical problems involved in finding optimal control strategies by nonlinear programming techniques. Annual Review in Automatic Programming, 13 (part 2):103–110, 1985. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413885904707.

Lecarme:1973:GPC

[Lec73] O. Lecarme. A general purpose conversational system for graphical programming. Annual Review in Automatic Programming, 7 (part 3):171–197, 1973. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413873900050.

Leitch:1992:KBC

[Lei92] R. Leitch. Knowledge based control: Selecting the right tool for the job. Annual Review in Automatic Programming, 17:1–9, 1992.

CODEN ???? ISSN ???? URL https://www.sciencedirect.
com/science/article/pii/S0066413809910035.

Lunderstaedt:1992:KBS

[LH92] R. A. Lunderstaedt and Th. Hillemann. Knowledge based sensor fault detection for gas turbines under consideration of model based methods. Annual Review in Automatic Programming, 17:287–292, 1992. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/S0066413809910473.

Lightbody:1992:NNM

[LI92] G. Lightbody and G. Irwin. Neural network modelling and control of a plant exhibiting the jump phenomena. Annual Review in Automatic Programming, 17:37–42, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/S0066413809910084.

Lions:1985:BCH

[Lio85] J-L Lions. Boundary control of hyperbolic systems and homogenization theory. Annual Review in Automatic Programming, 13 (part 2):95–101, 1985. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413885904690.

Lelashvili:1985:NAA

[LK85] Sh Lelashvili and K. Khundadze. The nonlinear algorithm of adaptive identification. Annual Review in Automatic Programming, 12 (part 1):340–343, 1985. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413885900588.

Laengle:1994:DCD

[LL94] Th Laengle and T. C. Lueth. Decentralized control of distributed intelligent robots and subsystems. Annual Review in Automatic Programming, 19:281–286, 1994. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413894900795.

Lun:1989:DIO

[LM89] V. Lun and I. M. MacLeod. Distributed intelligent objects in an architecture for real-time monitoring and control. Annual Review in Automatic Programming, 15 (part 2):25–30, 1989. CO-

DEN ???? ISSN ???? URL https://www.sciencedirect.com/
science/article/pii/0066413889900050.

Lord:1977:PES

[LMK77] R. Lord, J. Millar, and R. Kahane. A procedure for the estimation of software development costs. Annual Review in Automatic Programming, 8:211, 213–227, 1977. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413877900283.

Laplante:1994:RRT

[LMS94] P. Laplante, T. Marlowe, and A. Stoyenko. Requirements for a real-time image processing language. Annual Review in Automatic Programming, 18:175–180, 1994. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413894900302.

Linkens:1992:UAR

[LN92a] D. A. Linkens and Junhong Nie. A unifield approximate reasoning approach for real-time expert control system applications. Annual Review in Automatic Programming, 16 (part 1):69–74, 1992.

CODEN ???? ISSN ???? URL https://www.sciencedirect.
com/science/article/pii/006641389190011Y.

Linkens:1992:RTA

[LN92b] D. A. Linkens and H. O. Nyongesa. Real-time acquisition of fuzzy rules using genetic algorithms. Annual Review in Automatic Programming, 17:335–339, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/S0066413809910552.

Lan:1989:MCN

[LNYz89] Zhou Lan, Ye Nan, and Lu Yong-zai. Modelling and control for nonlinear time-delay system via pattern recognition approach. Annual Review in Automatic Programming, 15 (part 2):43–48, 1989. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413889900086.

Locke:1985:STS

[Loc85] M. Locke. Some trends in systems theory concerning the mathematical modeling of weekly structured systems. Annual Review in Automatic Programming, 12 (part 1):158–161, 1985. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413885900187.

Lukas:1989:EES

[LOKK89] M. P. Lukas, R. A. Oyen, M. A. Keyes, and A. Kaya. Evolution of expert systems for real-time process management: a case study on motor control. Annual Review in Automatic Programming, 15 (part 2):79–84, 1989. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413889900153.

LeParc:1994:TPM

[LQM94] P. Le Parc, B. Queguineur, and L. Marcé. Two proof methods for the GRAFCET language. Annual Review in Automatic Programming, 18:61–65, 1994. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413894900124.

Lunze:1992:LBP

[LS92] J. Lunze and F. Schiller. Logic-based process diagnosis utilising the causal structure of dynamical systems. Annual Review in Automatic Programming, 17:279–285, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/S0066413809910461.

Laube:1994:AIP

[LS94] S. J. P. Laube and E. F. Stark. Artificial intelligence in process control of pulsed laser deposition. Annual Review in Automatic Programming, 19:129–134, 1994. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/006641389490054X.

Lewoc:1990:DDB

[LŚT90] J. B. Lewoc, E. Ślusarska, and A. Tomczyk. A distributed data base for real-time control/monitoring of power plants. Annual Review in Automatic Programming, 15 (part 1):75–81, 1990. CO-

DEN ???? ISSN ???? URL https://www.sciencedirect.com/
science/article/pii/006641389090031L.

LeStrugeon:1992:KSR

[LTK92] E. Le Strugeon, M. Tendjaoui, and C. Kolski. Knowledge specification and representation for an “intelligent” interface devoted to process monitoring and supervision. Annual Review in Automatic Programming, 17:201–206, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/S0066413809910345.

Ludewig:1980:PCS

[Lud80] J. Ludewig. Process control software specification in PCSL. Annual Review in Automatic Programming, 10:103–109, 1980. CO-

DEN ???? ISSN ???? URL https://www.sciencedirect.com/
science/article/pii/0066413882900143.

Ludewig:1981:SSL

[Lud81] J. Ludewig. Specification of a specification language. Annual Review in Automatic Programming, 11:63–68, 1981. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413881900100.

Lukaszewicz:1961:SAC

[Łuk61] Leon Łukaszewicz. SAKO — an automatic coding system. Annual Review in Automatic Programming, 2:161–176, 1961. CO-

DEN ???? ISSN ???? URL https://www.sciencedirect.com/
science/article/pii/S0066413861800086.

Lunze:1985:SAI

[Lun85] J. Lunze. Stability analysis of interconnected systems by means of aggregate subsystem models. Annual Review in Automatic Programming, 12 (part 1):291–294, 1985. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413885900461.

Lucas:1969:FDP

[LW69] P. Lucas and K. Walk. On the formal description of PL/I. Annual Review in Automatic Programming, 6 (part 3):105–182, 1969.

CODEN ???? ISSN ???? URL https://www.sciencedirect.
com/science/article/pii/0066413869900056.

Li-wei:1989:NNB

[LwYz89] Bao Li-wei and Lu Yong-zai. Neural network based real-time production scheduling for industrial processes. Annual Review in Automatic Programming, 15 (part 2):85–90, 1989. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413889900165.

Marsh:1964:JC

[Mar64] Donald G. Marsh. JOVIAL in class. Annual Review in Automatic Programming, 4:167–181, January 1964. CO-DEN ARVAAM. ISSN 0066-4138 (print), 1878-545X (elec-

tronic).  URL https://www.sciencedirect.com/science/
article/pii/0066413864900175.

Martin:1992:OLM

[Mar92] David J. L. Martin. Organisation of a large multi-national systems development. Annual Review in Automatic Programming, 167 (part 2):1–9, 1992. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413892900027.

Miyoshi:1981:PSP

[MAU^{+}81] Y. Miyoshi, A. Asanuma, H. Ukaji, N. Nagata, and O. Sasaki. PROSUS: a software package for industrial control. Annual Review in Automatic Programming, 11:69–76, 1981. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413881900112.

Mazurkiewicz:1961:AFU

[Maz61] A. W. Mazurkiewicz. Arithmetic formulae and the use of subroutines in SAKO. Annual Review in Automatic Programming, 2:177–195, 1961. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/S0066413861800098.

Mills:1994:TNS

[MBH94] D. J. Mills, M. Brown, and C. J. Harris. Training neurofuzzy systems. Annual Review in Automatic Programming, 19:191– 196, 1994. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/0066413894900647.

Michal:1992:EAE

[MBK92] J. Michal, D. Burian, and P. Kmínek. Educational aspect of expert control of technological processes. Annual Review in Automatic Programming, 17:165–169, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/S006641380991028X.

Militky:1985:DIP

[MČ85] Jiří Militký and Jaroslav Čáp. Detection of influential points in general nonlinear problems. Annual Review in Automatic Programming, 12 (part 1):365–368, 1985. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413885900631.

Marchai:1992:DIS

[MC92] J. M. Marchai and E. F. Camacho. Design of an intelligent supervisor of a ship engine room. Annual Review in Automatic Programming, 17:141–146, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/S0066413809910242.

McGee:1969:GFP

[McG69] William C. McGee. Generalized file processing. Annual Review in Automatic Programming, 5 (part 2):77–149, January 1969. CODEN ARVAAM. ISSN 0066-4138 (print), 1878-545X

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0066413869900032.

McGettrick:1978:ARA

[McG78] Andrew D. McGettrick. An introduction to the formal definition of ALGOL 68. Annual Review in Automatic Programming, 9 (Part 1):1–84, January 1978. CODEN AR-VAAM. ISBN 0-08-023056-3. ISSN 0066-4138 (print), 1878-545x

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0066413878900010.

McKelvey:1992:NNA

[McK92] T. McKelvey. Neural networks applied to optimal flight control. Annual Review in Automatic Programming, 17:19–23, 1992. CO-

DEN ???? ISSN ???? URL https://www.sciencedirect.com/
science/article/pii/S0066413809910059.

Machado:1988:RTS

[MdCMC88] J. A. T. Machado, J. L. M. de Carvalho, J. A. S. Matos, and A. M. C. Costa. A real-time system for robot manipulator inverse dynamics computation. Annual Review in Automatic Programming, 14 (part 1):63–68, 1988. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413888900109.

Meggitt:1960:APN

[Meg60] J. E. Meggitt. Auto-programming for numerically controlled machine tools. Annual Review in Automatic Programming, 1: 220–229, 1960. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/0066413860900446.

Melli:1988:ASP

[Mel88] R. Melli. Automated software project planning and control. Annual Review in Automatic Programming, 14 (part 2):1–7, 1988.

CODEN ???? ISSN ???? URL https://www.sciencedirect.
com/science/article/pii/0066413890900018.

Mark:1977:EDI

[MEN77] Anthony Mark, Otto Eggenberger, and Jürgen Nehmer. Experiences in the design and implementation of a structured real-time operating system. Annual Review in Automatic Programming, 8:130, 133–137, 1977. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413877900180.

Meske:1994:CDA

[Mes94] Hans-Peter Meske. Conceptual design of an architecture for hard real time computing. Annual Review in Automatic Programming, 18:95–100, 1994. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413894900175.

Metcalfe:1964:PCB

[Met64] Howard H. Metcalfe. A parameterized compiler based on mechanical linguistics. Annual Review in Automatic Programming, 4:125–165, 1964. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/0066413864900163.

Milovanovic:1988:LSP

[MHS^{+}88] R. Milovanovic, M. Hajdukovic, Z. Salcic, E. Humo, and M. Kantardzic. A large software project — lessons learned. Annual Review in Automatic Programming, 14 (part 2):37–40, 1988. CO-

DEN ???? ISSN ???? URL https://www.sciencedirect.com/
science/article/pii/006641389090007E.

Mitchell:1988:APS

[Mit88] R. J. Mitchell. Applying the principle of separation of concerns in software development. Annual Review in Automatic Programming, 14 (part 1):21–27, 1988. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413888900043.

Morgan:1985:ASP

[ML85] J. Morgan and P. Loridan. Approximation of the Stackelberg problem and applications in control theory. Annual Review in Automatic Programming, 13 (part 2):121–124, 1985. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413885904720.

MacLeod:1992:TDR

[ML92] I. M. MacLeod and V. Lun. Towards distributed real-time intelligence. Annual Review in Automatic Programming, 16 (part 1):1–7, 1992. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/006641389190002S.

Magni:1992:GTO

[MMA92] J.-F. Magni, P. Mouyon, and M. I. Arsan. Geometric tools for an observer-based approach to residual generation. Annual Review in Automatic Programming, 17:229–233, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/S0066413809910382.

Munoz:1994:NUP

[MMO94] V. Muñoz, J. L. Martínez, and A. Ollero. Navigation with uncertain position estimation in the RAM-1 mobile robot. Annual Review in Automatic Programming, 19:215–219, 1994. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/006641389490068X.

Montague:1992:NNM

[MMW92] G. A. Montague, A. J. Morris, and M. J. Willis. Neural networks (methodologies for process modelling and control). Annual Review in Automatic Programming, 17:43–48, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/S0066413809910096.

Molnar:1985:NAH

[MNV85] Łudovít Molnár, Pavol Návrat, and Vladimír Vojtek. A new algorithm for heuristic search which uses hints. Annual Review in Automatic Programming, 12 (part 1):177–181, 1985. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413885900229.

Moody:1990:DDD

[Moo90] J. K. M. Moody. Digital design with a distributed environment in mind. Annual Review in Automatic Programming, 15 (part 1):9–15, 1990. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/006641389090020R.

Moritz:1985:COL

[Mor85] K. Moritz. Computation of open-loop solutions for zero-sum differential games by parametrization. Annual Review in Automatic Programming, 13 (part 2):125–130, 1985. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413885904732.

Mori:1992:MCP

[Mor92] H. Mori. Monitoring and control of power system voltage stability using an artificial neural network. Annual Review in Automatic Programming, 17:81–86, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/S0066413809910151.

Mostert:1994:ASR

[Mos94] S. Mostert. An architecture supporting real-time programming and implementations. Annual Review in Automatic Programming, 18:145–150, 1994. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413894900256.

Motus:1992:AIH

[Mot92] L. Motus. Artificial intelligence in hard real-time — a new paradigm needed? Annual Review in Automatic Programming, 16 (part 1):9–12, 1992. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/006641389190003T.

Motus:1994:TAI

[Mot94] L. Motus. Trends in artificial intelligence applications for real-time control (a speculative study). Annual Review in Automatic Programming, 19:31–41, 1994. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413894900396.

Massacci:1985:OCS

[MP85] P. Massacci and G. Patrizi. Optimal control and stochastic control policies in mineral and metallurgical processing. Annual Review in Automatic Programming, 13 (part 2):111–120, 1985. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413885904719.

Maechtel:1994:ROS

[MR94] M. Maechtel and H. Rzehak. On realtime operating systems: How to compare performance? Annual Review in Automatic Programming, 18:139–144, 1994. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413894900248.

Mazo:1994:ECW

[MRL^{+}94] M. Mazo, F. J. Rodriguez, J. L. Lázaro, J. Ureña, J. C. García, E. Santiso, P. Revenga, and J. J. García. Electronic control for a wheel-chair guided by oral commands and ultrasonic and infrared sensors. Annual Review in Automatic Programming, 19:249–254, 1994. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413894900744.

Mazo:1994:RFA

[MRSS94] M. Mazo, F. J. Rodriguez, E. Santiso, and M. A. Sotelo. Road following by artificial vision using neural network. Annual Review in Automatic Programming, 19:209–214, 1994. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413894900671.

Meijer:1992:ICE

[MS92a] B. R. Meijer and J. Stigter. Integration of control and error management for a flexible assembly cell, using a cost function. Annual Review in Automatic Programming, 17:207–212, 1992. CO-

DEN ???? ISSN ???? URL https://www.sciencedirect.com/
science/article/pii/S0066413809910357.

Mittermeir:1992:SIS

[MS92b] R. T. Mittermeir and R. A. Schlemmer. Stepwise improvement of the software-process in a multidimensional framework. Annual Review in Automatic Programming, 167 (part 2):63–70, 1992. CO-

DEN ???? ISSN ???? URL https://www.sciencedirect.com/
science/article/pii/006641389290010M.

Macleod:1994:SSD

[MS94] I. M. Macleod and A. Stothert. A simulation study of distributed intelligent control for a deep shaft mine winder. Annual Review in Automatic Programming, 19:293–298, 1994. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413894900817.

Marlowe:1994:MGO

[MSL^{+}94] T. J. Marlowe, A. D. Stoyenko, P. A. Laplante, R. S. Daita, C. C. Amaro, C. M. Nguyen, and S. L. Howell. Multiple-goal objective functions for optimization of task assignment in complex computer systems. Annual Review in Automatic Programming, 18:55–60, 1994. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413894900116.

Morasso:1977:ICB

[MSTZ77] P. Morasso, G. Sandini, V. Tagliasco, and R. Zaccaria. Interactive control of a biomedical manipulator. Annual Review in Automatic Programming, 8:199–205, 1977. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/006641387790026X.

Miki:1994:PPR

[MTK^{+}94] M. Miki, Y. Tamura, T. Kakumu, T. Hirai, and K. Itoh. Performance prototyping for a real time system with qualitative parameter tuning methods. Annual Review in Automatic Programming, 18:27–32, 1994. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/006641389490006X.

Nilsson:1992:MBD

[NAP92] A. Nilsson, K.-E. Arzén, and T. F. Petti. Model-based diagnosisstate transition events and constraint equations. Annual Review in Automatic Programming, 17:359–364, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/S006641380991059X.

Naur:1964:ARR

[Nau64a] Peter Naur. Appendix: Revised report on the algorithmic language ALGOL 60. Annual Review in Automatic Programming, 4:217–258, January 1964. CODEN ARVAAM. ISSN 0066-4138 (print), 1878-545X (electronic). URL https://www.

sciencedirect.com/science/article/pii/0066413864900205.

Naur:1964:DGA

[Nau64b] Peter Naur. The design of the GIER ALGOL compiler. Annual Review in Automatic Programming, 4:49–85, January 1964. CODEN ARVAAM. ISSN 0066-4138 (print), 1878-545X

(electronic). URL https://www.sciencedirect.com/science/
article/pii/006641386490014X.

Nevriva:1985:DBO

[Nev85] Pavel Nevřiva. The data-based optimal control algorithms. Annual Review in Automatic Programming, 12 (part 1):390– 393, 1985. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/0066413885900692.

Nakanishi:1977:DLS

[NHIK77] Hiroaki Nakanishi, Koji Harai, Jushi Ide, and Toshiro Kamiuchi. Distributed load sharing functions in multi-computer system. Annual Review in Automatic Programming, 8:119–124, 1977. CO-

DEN ???? ISSN ???? URL https://www.sciencedirect.com/
science/article/pii/0066413877900167.

Nohmi:1981:MBB

[NIYY81] M. Nohmi, H. Ihara, M. Yabushita, and M. Yasunami. Microcomputer based on-board real-time control system for automated guideway rapid transit system. Annual Review in Automatic Programming, 11:43–48, 1981. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413881900070.

Nowinski:1992:KAP

[NK92] G. Nowinski and V. Krebs. A knowledge acquisition and processing strategy based on formal concept analysis. Annual Review in Automatic Programming, 17:103–108, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/S0066413809910187.

Nacsa:1994:CPE

[NK94] J. Nacsa and G. L. Kovács. Communication problems of expert systems in manufacturing environment. Annual Review in Automatic Programming, 19:243–247, 1994. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413894900736.

Naito:1981:MPP

[NKIT81] K. Naito, H. Kato, K. Ikenouchi, and T. Tagami. The multi-processor process computer system employing a data way used at the yawata no. 3 steelmaking plant. Annual Review in Automatic Programming, 11:35–42, 1981. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413881900069.

Nicolas:1994:DFH

[NLÁR94] C. F. Nicolás, E. Lombraña, A. Álvarez, and R. Reyero. Derivation of fuzzy hybrid models for real-time fuzzy control design: Application to a furnace. Annual Review in Automatic Programming, 19:85–89, 1994. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413894900477.

Nakamori:1992:IHF

[NS92] Y. Nakamori and Y. Sawaragi. Implementation of a humanfriendly systems methodology for intelligent control system modelling and simulation. Annual Review in Automatic Programming, 17:109–114, 1992. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/S0066413809910199.

Nakagawa:1992:PKA

[NSY92] T. Nakagawa, Y. Sawaragi, and Y. Yagihara. Process knowledge acquisition and control by quantitative and qualitative complementarity. Annual Review in Automatic Programming, 17: 353–358, 1992. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/S0066413809910588.

Ng:1994:TPR

[NY94] Jim M. Ng and Norris T. C. Yu. Transport protocol for real-time multimedia communication. Annual Review in Automatic Programming, 18:15–20, 1994. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413894900043.

Oberle:1985:NTM

[Obe85] H. J. Oberle. Numerical treatment of minimax control problems by multiple shooting. Annual Review in Automatic Programming, 13 (part 2):131–138, 1985. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413885904744.

Orasanu:1985:CSP

[OGF85] L. Orăşanu, R. Gaspar, and F. G. Filip. A CAD system of the physical control structure for large scale systems. Annual Review in Automatic Programming, 12 (part 1):445–448, 1985. CO-

DEN ???? ISSN ???? URL https://www.sciencedirect.com/
science/article/pii/0066413885900825.

Ozaki:1992:RBI

[OHSH92] N. Ozaki, K. Harayama, W. Shinohara, and S. Hayashi. Recovery boiler intelligent control. Annual Review in Automatic Programming, 16 (part 1):139–143, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413891900235.

Ono:1989:EFS

[Ono89] O. Ono. An extended feedback structure of intelligent computeraided control systems design based on object-oriented language. Annual Review in Automatic Programming, 15 (part 2):59– 64, 1989. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/0066413889900116.

Ortner:1988:DSP

[OR88] E. Ortner and J. Röhrle. Dictionary-supported prototyping of database applications. Annual Review in Automatic Programming, 14 (part 2):21–25, 1988. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/006641389090004B.

Ord-Smith:1960:SZS

[OS60] R. J. Ord-Smith. The STANTEC-ZEBRA simple code and its interpretation. Annual Review in Automatic Programming, 1: 146–168, 1960. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/0066413860900380.

Osyczka:1985:IMD

[Osy85]  Andrzej Osyczka.  An interactive multicriterion decision

making system. Annual Review in Automatic Programming, 12 (part 1):208–211, 1985. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413885900291.

Ohba:1981:FHL

[OTTK81] M. Ohba, Y. Tanitsu, N. Takimoto, and H. Kadota. Functor: a higher-level co-operating program model. Annual Review in Automatic Programming, 11:21–28, 1981. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413881900045.

Palmgren:1988:MBD

[PA88] B. Palmgren and F. Aasma. Microprocessor based development tool — design and realization of SISO controllers. Annual Review in Automatic Programming, 14 (part 1):57–62, 1988. CO-

DEN ???? ISSN ???? URL https://www.sciencedirect.com/
science/article/pii/0066413888900092.

Pao:1992:NNC

[Pao92a] Yoh-Han Pao. Neural net control. Annual Review in Automatic Programming, 17:11–17, 1992. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/S0066413809910047.

Pao:1992:RTE

[Pao92b] Yoh-Han Pao. Real-time environments for intelligent control. Annual Review in Automatic Programming, 17:475–482, 1992. CO-

DEN ???? ISSN ???? URL https://www.sciencedirect.com/
science/article/pii/S0066413809910771.

Pao:1994:CIP

[Pao94] Y. H. Pao. A computational intelligence perspective on process monitoring and optimization. Annual Review in Automatic Programming, 19:25–30, 1994. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413894900388.

Parizek:1985:ERS

[Par85] Peter Parízek. Experience and results in simulation modeling of computer networks. Annual Review in Automatic Programming, 12 (part 1):441–444, 1985. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413885900813.

Payne:1960:OEP

[Pay60] W. F. M. Payne. Operational experience with the PEGASUS Autocode. Annual Review in Automatic Programming, 1:58– 63, 1960. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/006641386090032X.

Pessi:1994:MRT

[PB94] E. Pessi and E. Bullo. MORSAF: a real-time assistant for the management of a petrochemical plant. Annual Review in Automatic Programming, 19:341–346, 1994. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413894900884.

Peters:1994:LDN

[PBMR94] J. F. Peters, L. Baumela, D. Maravall, and S. Ramanna. Logical design of neural controllers. Annual Review in Automatic Programming, 19:179–184, 1994. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413894900620.

Patton:1992:LRC

[PC92] R. J. Patton and J. Chen. On-line residual compensation in robust fault diagnosis of dynamic systems. Annual Review in Automatic Programming, 17:221–227, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/S0066413809910370.

Pao:1992:NNA

[PCY92] Yoh-Han Pao, H. J. Chizeck, and P. P. C. Yip. Neural-net augmentation of adaptive predictive control. Annual Review in Automatic Programming, 16 (part 1):109–113, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413891900187.

Pyle:1980:APC

[PEGR80] I. C. Pyle, P. Elzer, S. J. Goldsack, and J. Robert. ADA in process control. Annual Review in Automatic Programming, 10: 57–62, ???? 1980. CODEN ARVAAM. ISSN 0066-4138 (print),

1878-545X (electronic). URL https://www.sciencedirect.
com/science/article/pii/0066413882900088.

Perne:1992:SCE

[Per92] R. Perne. Supervision and control of an exothermic batch process. Annual Review in Automatic Programming, 17:247–251, 1992.

CODEN ???? ISSN ???? URL https://www.sciencedirect.
com/science/article/pii/S0066413809910412.

Pym:1961:EAM

[PF61] J. Pym and G. K. Findlay. The Elliott 803 Autocode Mark 2. Annual Review in Automatic Programming, 2:77–113, 1961. CO-

DEN ???? ISSN ???? URL https://www.sciencedirect.com/
science/article/pii/S0066413861800050.

Precigout:1992:ESS

[PF92] J.-P Précigout and J. Favaro. The European space software development environment. Annual Review in Automatic Programming, 167 (part 2):95–102, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/006641389290015H.

Pfeiffer:1994:IFR

[Pfe94] Bernd-Markus Pfeiffer. Identification of fuzzy rules from learning data. Annual Review in Automatic Programming, 19:49–54, 1994.

CODEN ???? ISSN ???? URL https://www.sciencedirect.
com/science/article/pii/0066413894900418.

Peter:1985:IIP

[PFTD85] Lorenz Peter, Preuss Fritz, Schulze Thomas, and Ziems Dietrich. IPS — interactive programming environment for simulation. Annual Review in Automatic Programming, 12 (part 1): 487–489, 1985. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/006641388590093X.

Pearcey:1960:MSA

[PHW60] T. Pearcey, S. N. Higgins, and P. M. Woodward. The mark 5 system of automatic coding for TREAC. Annual Review in Automatic Programming, 1:23–31, 1960. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413860900306.

Pierson:1985:OAL

[Pie85] B. L. Pierson. Optimal aircraft landing-approach trajectories: a comparison of two dynamic models. Annual Review in Automatic Programming, 13 (part 2):139–145, 1985. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413885904756.

Pirotte:1973:ATP

[Pir73] Alain Pirotte. Automatic theorem proving based on resolution. Annual Review in Automatic Programming, 7 (part 4): 201–266, 1973. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/0066413873900013.

Pendic:1988:AEQ

[PKS88] Z. R. Pendić, Lj Kovačević, and J. Stupar. An approach to evaluation of quality of integrated information systems. Annual Review in Automatic Programming, 14 (part 2):63–68, 1988. CO-

DEN ???? ISSN ???? URL https://www.sciencedirect.com/
science/article/pii/006641389090012G.

Peschel:1985:ECM

[PMH85] M. Peschel, W. Mende, and J. Hofbauer. Exponential chains as models for growth processes. Annual Review in Automatic Programming, 12 (part 1):82–90, 1985. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413885900084.

Pohjolainen:1985:STR

[Poh85] S. Pohjolainen. A self-tuning robust controller for semilinear parabolic systems. Annual Review in Automatic Programming, 13 (part 2):147–156, 1985. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413885904768.

Popescu:1985:ECA

[Pop85] Th. D. Popescu. Experiences with a computer aided procedure for time series analysis and forecasting using Box-Jenkins philosophy. Annual Review in Automatic Programming, 12 (part 1):361–364, 1985. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/006641388590062X.

Park:1994:NNB

[PPE^{+}94] G. H. Park, Y. H. Pao, K. G. Eyink, S. R. Leclair, and M. S. Soclof. Neural-net based optical ellipsometry for monitoring growth of semiconductor films. Annual Review in Automatic Programming, 19:123–128, 1994. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413894900531.

Perez:1994:IKR

[PPMM94] J. A. Perez, J. Pujol, and P. R. Muro-Medrano. Increasing a knowledge representation schema for FMS control with fault detection and error recovery capabilities. Annual Review in Automatic Programming, 19:313–317, 1994. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413894900841.

Pleyber:1977:SSL

[PS77] Joël Pleyber and Manuel Silva. Software specification language for sequential processes. Annual Review in Automatic Programming, 8:67–73, 1977. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413877900106.

Plustwik:1985:TRT

[PS85] Hermann W. Plustwik and Peter P. Schreiber. A trilogy for real time information. Annual Review in Automatic Programming, 12 (part 1):212–225, 1985. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413885900308.

Puchol:1994:AHP

[Puc94] Carlos Puchol. Achieving high performance sonar-based wallfollowing. Annual Review in Automatic Programming, 19:233– 236, 1994. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/006641389490071X.

Puta:1985:TDU

[Put85] H. Puta. Trajectory decomposition used for control design in water systems. Annual Review in Automatic Programming, 12 (part 1):303–306, 1985. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413885900497.

Prenevitchius:1985:ACN

[PV85] H. Prenevitchius and E. Valekevitchius. Automatic construction of numerical models for systems represented by Markovian processes. Annual Review in Automatic Programming, 12 (part 1):336–339, 1985. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413885900576.

Papp:1992:EER

[PvWD92] Z. Papp, J. A. van Woerden, and B. J. F. Driessen. An execution environment for real-time model-based supervisory control and diagnostic systems. Annual Review in Automatic Programming, 17:521–526, 1992. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/S0066413809910849.

Rauber:1994:LSP

[RBSG94] T. W. Rauber, M. M. Barata, and A. S. Steiger-Garção. Learning of specific process monitors in machine tool supervision. Annual Review in Automatic Programming, 19:105–110, 1994. CO-

DEN ???? ISSN ???? URL https://www.sciencedirect.com/
science/article/pii/0066413894900507.

Raju:1992:TR

[RC92] G. K. Raju and C. L. Cooney. Temporal reasoning. Annual Review in Automatic Programming, 17:447–452, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/S0066413809910734.

Rastrigin:1985:APM

[RE85] L. A. Rastrigin and J. Y. Eiduk. Adaptation in problems of multi-criteria optimization. Annual Review in Automatic Programming, 12 (part 1):226–229, 1985. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/006641388590031X.

Redish:1960:SPU

[Red60] K. A. Redish. Some problems of a universal autocode. Annual Review in Automatic Programming, 1:16–22, 1960. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/006641386090029X.

Reinisch:1985:HLC

[Rei85a] K. Reinisch. Hierarchical on-line control with applications to regional water distribution and agricultural systems. Annual Review in Automatic Programming, 12 (part 1):91–104, 1985. CO-

DEN ???? ISSN ???? URL https://www.sciencedirect.com/
science/article/pii/0066413885900096.

Reinschke:1985:SPL

[Rei85b] K. J. Reinschke. Structural properties of linear control systems should be investigated by means of graphs. Annual Review in Automatic Programming, 12 (part 1):280–283, 1985. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413885900436.

Rodd:1994:NPD

[RHJ^{+}94] M. G. Rodd, J. D. Holt, A. V. Jones, M. Zhao, and M. B. Zungu. A new paradigm for distributed, integrated, real-time control systems. Annual Review in Automatic Programming, 19:287–292, 1994. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413894900809.

Rich:1961:ACC

[Ric61] R. P. Rich. APT, a common computer language. Annual Review in Automatic Programming, 2:141–159, 1961. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/S0066413861800074.

Rudzinska:1985:SOC

[RK85] K. Rudzinska and H. Kormanski. Software for optimal control of dynamic systems. Annual Review in Automatic Programming, 13 (part 2):171–176, 1985. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413885904781.

Respondek:1989:HSS

[RK89] T. Respondek and V. Krebs. Hardware and software structure of a real-time expert system for control of chemical plants. Annual Review in Automatic Programming, 15 (part 2):37–42, 1989.

CODEN ???? ISSN ???? URL https://www.sciencedirect.
com/science/article/pii/0066413889900074.

Rew:1985:CSA

[RKC85] Dongwon Rew, H. J. Kelley, and E. M. Cliff. A computational study of active-set logics for nonlinear programming. Annual Review in Automatic Programming, 13 (part 2):157–170, 1985. CO-

DEN ???? ISSN ???? URL https://www.sciencedirect.com/
science/article/pii/006641388590477X.

Rutherford:1994:CSP

[RLM94] P. Rutherford, B. Lennox, and G. A. Montague. Case studies in process modelling and condition monitoring using artificial neural networks. Annual Review in Automatic Programming, 19:141–146, 1994. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413894900566.

Renders:1992:GAP

[RNB92] J. M. Renders, J. P. Nordvik, and H. Bersini. Genetic algorithms for process control: a survey. Annual Review in Automatic Programming, 17:323–328, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/S0066413809910539.

Schafer:1985:OCC

[rnhWS85] rer. nat. habil. Wolfgang Schäfer. Optimal control of complex systems. Annual Review in Automatic Programming, 12 (part 1): 284–287, 1985. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/0066413885900448.

Robinson:1960:APD

[Rob60] C. Robinson. Automatic programming on DEUCE. Annual Review in Automatic Programming, 1:111–126, 1960. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413860900367.

Rodd:1990:RTI

[Rod90] M. G. Rodd. Real-time issues in distributed data bases for real-time control. Annual Review in Automatic Programming, 15 (part 1):1–7, 1990. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/006641389090019N.

Ronaldson:1960:PEA

[Ron60] P. M. Ronaldson. PEGASUS: an example of an autocoded program for sales analysis and forecasting. Annual Review in Automatic Programming, 1:64–80, 1960. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413860900331.

Ramanna:1992:ECT

[RP92] S. Ramanna and J. F. Peters. Explicit clock temporal logic in constraint checking for real-time systems. Annual Review in Automatic Programming, 16 (part 1):47–58, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/006641389190009Z.

Robinson:1992:DIM

[RPF^{+}92] G. Robinson, S. Pallett, R. Fripp, J. Mullhi, and A. Spence. The development of an intelligent monitoring and control system for a solvent extraction process. Annual Review in Automatic Programming, 17:383–389, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/S0066413809910631.

Rauterberg:1992:WOS

[RS92a] M. Rauterberg and O. Strohm. Work organization and software development. Annual Review in Automatic Programming, 167 (part 2):121–128, 1992. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/006641389290019L.

Rodd:1992:P

[RS92b] M. G. Rodd and G. J. Suski. Preface. Annual Review in Automatic Programming, 16 (part 1):vii, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/006641389190001R.

Romano:1992:NLI

[RTC92] J. M. G. Romano, J. A. Ternero, and E. F. Camacho. Natural language interface for process control centers. Annual Review in Automatic Programming, 16 (part 1):103–108, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413891900176.

Rushing:1985:DLS

[Rus85] J. R. Rushing. A description of an on-line statistical quality control package. Annual Review in Automatic Programming, 13 (part 1):9–16, 1985. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413885904379.

Russell:1992:FSA

[Rus92] D. W. Russell. Further studies in A.I. augmented process control using the BOXES methodology. Annual Review in Automatic Programming, 16 (part 1):75–79, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/006641389190012Z.

Rutishauser:1961:IAP

[Rut61] H. Rutishauser. Interference with an ALGOL-procedure. Annual Review in Automatic Programming, 2:67–75, January 1961. CODEN ARAPDX. ISSN 0066-4138 (print), 1878-545X

(electronic). URL https://www.sciencedirect.com/science/
article/pii/S0066413861800049.

Rutishauser:1963:URPa

[Rut63] H. Rutishauser. The use of recursive procedures in ALGOL 60. Annual Review in Automatic Programming, 3:43–51, January 1963. CODEN ARVAAM. ISSN 0066-4138 (print), 1878-545X

(electronic). URL https://www.sciencedirect.com/science/
article/pii/S006641386380004X.

Rodd:1989:KBV

[RW89] M. G. Rodd and Q. M. Wu. Knowledge-based vision systems in real-time control. Annual Review in Automatic Programming, 15 (part 2):13–18, 1989. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413889900037.

Rada:1992:CSR

[RWMC92] R. Rada, W. Wang, A. Michailidis, and C. Chen. Collaborative software reuse integrated into the work flow. Annual Review in Automatic Programming, 167 (part 2):79–83, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/006641389290012E.

Sabathe:1977:CIA

[Sab77] J. P. Sabathe. Computer implementation of automata for electricity distribution control. Annual Review in Automatic Programming, 8:169, 171–180, 1977. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413877900234.

Sammet:1961:DDC

[Sam61a] Jean E. Sammet. Detailed description of COBOL. Annual Review in Automatic Programming, 2:197–230, 1961. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/S0066413861800104.

Sammet:1961:GVC

[Sam61b] Jean E. Sammet. General views on COBOL. Annual Review in Automatic Programming, 2:345–349, 1961. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/S0066413861800153.

Samuel:1969:SSM

[Sam69] A. L. Samuel. Some studies in machine learning using the game of checkers. II — recent progress. Annual Review in Automatic Programming, 6 (part 1):1–36, 1969. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413869900044.

Singh:1985:PSM

[SAM85] M. G. Singh, A. Y. Allidina, and K. Malinowski. Parallel simulation methods for dynamical systems: issues and challenges. Annual Review in Automatic Programming, 12 (part 1):15–28, 1985.

CODEN ???? ISSN ???? URL https://www.sciencedirect.
com/science/article/pii/0066413885900023.

Sarancha:1985:MMT

[Sar85] D. A. Sarancha. Mathematical models of tundra communities and populations. Annual Review in Automatic Programming, 12 (part 1):377–379, 1985. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413885900667.

Saukkonen:1992:SPM

[Sau92] S. Saukkonen. Software project management education — experiences of training full-time professionals. Annual Review in Automatic Programming, 167 (part 2):139–142, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/006641389290022H.

Savoysky:1980:ADA

[Sav80] S. Savoysky. Analysis and description of automatic control system. Annual Review in Automatic Programming, 10:45–55, ???? 1980. CODEN ARVAAM. ISSN 0066-4138 (print), 1878-545X

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0066413882900076.

Savoysky:1981:UAS

[Sav81] S. Savoysky. The use of Ada for the specification of automata in civil engineering. Annual Review in Automatic Programming, 11: 129–138, ???? 1981. CODEN ARVAAM. ISSN 0066-4138 (print),

1878-545X (electronic). URL https://www.sciencedirect.
com/science/article/pii/0066413881900197.

Simo:1994:RTP

[SC94] José Simó and Alfons Crespo. Real time planning in n-dim state space. Annual Review in Automatic Programming, 19: 307–312, 1994. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/0066413894900833.

Schrott:1977:GDR

[Sch77] Gerhard Schrott. Generation of dedicated realtime operating systems by dialogue. Annual Review in Automatic Programming, 8:145–151, 1977. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413877900209.

Schoitsch:1980:ICC

[Sch80] E. Schoitsch. Implementation of conditional critical regions in systems using different process communication techniques. Annual Review in Automatic Programming, 10:29–38, 1980. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413882900052.

Schmandt:1985:SET

[Sch85a] Jurgen Schmandt. Scientific evidence and technically complex policy issues. Annual Review in Automatic Programming, 12 (part 1):170–173, 1985. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413885900205.

Schmidt:1985:GFV

[Sch85b] B. Schmidt. GPSS-FORTRAN version 3. Annual Review in Automatic Programming, 12 (part 1):139–142, 1985. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/006641388590014X.

Schwarze:1985:UMM

[Sch85c] Gunter Schwarze. Use of models and model experiments solving problems in informatics. Annual Review in Automatic Programming, 12 (part 1):125–132, 1985. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413885900126.

Scherr:1992:IRS

[Sch92] W. Scherr. Introducing reuse of software designs into a company. Annual Review in Automatic Programming, 167 (part 2): 45–53, 1992. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/006641389290008D.

Schmid:1994:SUD

[Sch94a] U. Schmid. Synchronized UTC for distributed real-time systems. Annual Review in Automatic Programming, 18:101–107, 1994.

CODEN ???? ISSN ???? URL https://www.sciencedirect.
com/science/article/pii/0066413894900183.

Schrott:1994:DTO

[Sch94b] G. Schrott. A distributed task-oriented real-time programming system. Annual Review in Automatic Programming, 18:163– 166, 1994. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/0066413894900280.

Sedillot:1990:PCH

[Séd90] S. Sédillot. Process control: How to process distributed data in less than a second? Annual Review in Automatic Programming, 15 (part 1):17–22, 1990. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/006641389090021I.

Stroeve:1992:ANN

[SFVvL92] S. H. Stroeve, H. M. Franken, P. H. Veltink, and W. T. C. van Luenen. Adaptive neural network control of fes-induced cyclical lower leg movements. Annual Review in Automatic Programming, 17:25–30, 1992. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/S0066413809910060.

Schy:1985:MIC

[SG85] A. A. Schy and D. P. Giesy. Multiobjective insensitive computeraided design of aerospace control systems. Annual Review in Automatic Programming, 13 (part 2):177–188, 1985. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413885904793.

Sproesser:1992:MFD

[SG92] T. Sproesser and G. L. Gissinger. A method for fault detection using parameter and state estimation. Annual Review in Automatic Programming, 17:241–246, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/S0066413809910400.

Song:1989:RBD

[SGZ89] Z. Song, J. C. Gao, and C. H. Zhou. The rule-based distillation decision and control system. Annual Review in Automatic Programming, 15 (part 2):101–106, 1989. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413889900190.

Shaw:1963:JPL

[Sha63] C. J. Shaw. Jovial — a programming language for real-time command systems. Annual Review in Automatic Programming, 3:53– 119, January 1963. CODEN ARVAAM. ISSN 0066-4138 (print),

1878-545X (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0066413863800051.

Shi:1989:ESS

[Shi89] Zhongke Shi. Expert system for sensor failure detection of aircraft. Annual Review in Automatic Programming, 15 (part 2): 91–96, 1989. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/0066413889900177.

Sima:1985:CML

[Sim85] V. Sima. Convergence of multirecursive least squares adaptive algorithms. Annual Review in Automatic Programming, 12 (part 1):414–417, 1985. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413885900746.

Sassen:1992:DRT

[SJ92] J. M. A. Sassen and R. B. M. Jaspers. Designing realtime knowledge based systems with perfect. Annual Review in Automatic Programming, 17:491–496, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/S0066413809910795.

Sjoberg:1992:RSP

[Sjö92] J. Sjöberg. Regularization as a substitute for pre-processing of data in neural network training. Annual Review in Automatic Programming, 17:31–35, 1992. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/S0066413809910072.

Schram:1994:OAC

[SKKG94] G. Schram, L. Karsten, B. J. A. Kröse, and F. C. A. Groen. Optimal attitude control of satellites by artificial neural networks: a pilot study. Annual Review in Automatic Programming, 19:173–178, 1994. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413894900612.

Svireshev:1985:MEF

[SKL85] Yu. M. Svireshev, V. F. Krapivin, and D. O. Logofet. Mathematical ecology: Forecast, assessment and optimization of impacts on natural ecosystems. Annual Review in Automatic Programming, 12 (part 1):114–124, 1985. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413885900114.

Shen:1992:PC

[SL92a] Q. Shen and R. R. Leitch. Process control. Annual Review in Automatic Programming, 17:365–370, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/S0066413809910606.

Sonnleitner:1992:PRB

[SL92b] B. Sonnleitner and G. Locher. Pattern recognition for bioprocess control. Annual Review in Automatic Programming, 17: 431–433, 1992. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/S0066413809910709.

Smuts:1989:UDA

[SM89] W. B. Smuts and I. M. MacLeod. Using discrete AI techniques for designing a real-world control supervisor. Annual Review in Automatic Programming, 15 (part 2):31–35, 1989. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413889900062.

Schweinzer:1994:CSF

[SM94] Herbert Schweinzer and Christoph Mittermayer. Combined sequential and function-oriented programming for real-time systems. Annual Review in Automatic Programming, 18:181– 185, 1994. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/0066413894900310.

Smedema:1977:P

[Sme77] C. H. Smedema. Preface. Annual Review in Automatic Programming, 8:ix, 1977. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/0066413877900015.

Smith:1970:JID

[Smi70] Joseph W. Smith. JOSS — II: Design philosophy. Annual Review in Automatic Programming, 6 (part 4):183–256, 1970. CO-

DEN ???? ISSN ???? URL https://www.sciencedirect.com/
science/article/pii/0066413870900017.

Sundin:1994:DUI

[SNG94] Ulf Sundin, Peter Nordin, and Shahab Ghavami. Designing user interfaces for applications based on the hint architecture. Annual Review in Automatic Programming, 19:365–370, 1994. CO-

DEN ???? ISSN ???? URL https://www.sciencedirect.com/
science/article/pii/0066413894900922.

Sassen:1992:PIR

[SOJ92] J. M. A. Sassen, A. Ollongren, and R. B. M. Jaspers. Predicting and improving response-times of perfect-models. Annual Review in Automatic Programming, 17:515–520, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/S0066413809910837.

Son:1990:PBS

[Son90] Sang H. Son. On priority-based synchronization protocols for distributed real-time database systems. Annual Review in Automatic Programming, 15 (part 1):29–33, 1990. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/006641389090023K.

Sowa:1981:RTM

[Sow81] M. Sowa. Real time multi-microprocessor system introduced concept of data-driven. Annual Review in Automatic Programming, 11:115–120, 1981. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413881900173.

Sala:1994:ARC

[SP94] A. Sala and J. Picó. Analysis of rulebase coherence in fuzzy control systems. Annual Review in Automatic Programming, 19: 79–84, 1994. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/0066413894900469.

Spenke:1985:DPL

[Spe85] Volker Spenke. On the development of the programming language CSS-PASCAL for continuous system simulation. Annual Review in Automatic Programming, 12 (part 1):483–486, 1985.

CODEN ???? ISSN ???? URL https://www.sciencedirect.
com/science/article/pii/0066413885900928.

Spiegel:1985:IPE

[Spi85] Bernhard Spiegel. An interactive procedure for experimental multi-criterion optimization. Annual Review in Automatic Programming, 12 (part 1):239–245, 1985. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413885900345.

Schmidt:1985:MBO

[SR85] M. Schmidt and K. Reinisch. Model building and optimization of the plant growth process of the greenhouse cucumber. Annual Review in Automatic Programming, 12 (part 1):394–397, 1985.

CODEN ???? ISSN ???? URL https://www.sciencedirect.
com/science/article/pii/0066413885900709.

Sanchez:1994:DDA

[SRRP^{+}94] M. Sànchez, I. R-Roda, M. Poch, U. Cortés, and J. Lafuente. DAI-DEPUR architecture: Distributed agents for real-time wwtp supervision and control. Annual Review in Automatic Programming, 19:147–152, 1994. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413894900574.

Shapiro:1970:NAO

[SS70] R. M. Shapiro and H. Saint. A new approach to optimization of sequencing decisions. Annual Review in Automatic Programming, 6 (part 5):257–288, 1970. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413870900030.

Szmuc:1994:HCV

[SSSS94] T. Szmuc, P. Szwed, J-J Schwarz, and J. Skubich. Hierarchical correctness verification in multiphase real-time software design. Annual Review in Automatic Programming, 18:87–94, 1994. CO-

DEN ???? ISSN ???? URL https://www.sciencedirect.com/
science/article/pii/0066413894900167.

Stanciulescu:1985:PMS

[Stă85a] F. Stănciulescu. Principles of modelling and simulation, large scale and complex systems. application in ecology. Annual Review in Automatic Programming, 12 (part 1):133–138, 1985. CO-

DEN ???? ISSN ???? URL https://www.sciencedirect.com/
science/article/pii/0066413885900138.

Stantchev:1985:SAD

[Sta85b] I. Stantchev. System analysis and decision making with interdependent goals-a linguistic approach. Annual Review in Automatic Programming, 12 (part 1):194–197, 1985. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413885900266.

Starke:1985:PNM

[Sta85c] Peter H. Starke. Petri-Netz-Maschine: a software tool for analysis and validation of Petri nets. Annual Review in Automatic Programming, 12 (part 1):474–475, 1985. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413885900898.

Steel:1961:UMF

[Ste61] Thomas B. Steel. UNCOL: The myth and the fact. Annual Review in Automatic Programming, 2:325–344, 1961. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/S0066413861800141.

Stegbauer:1990:ICI

[Ste90] W. Stegbauer. Intertask-communication inside a real-time database. Annual Review in Automatic Programming, 15 (part 1):63–69, 1990. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/006641389090029Q.

Stockman:1988:RSP

[Sto88] S. G. Stockman. The recovery of software projects. Annual Review in Automatic Programming, 14 (part 2):53–58, 1988. CO-

DEN ???? ISSN ???? URL https://www.sciencedirect.com/
science/article/pii/006641389090010O.

Styer:1992:PIA

[SV92] D. L. Styer and V. Vemuri. Preprocessing inputs for adaptive critic control. Annual Review in Automatic Programming, 16 (part 1):27–31, 1992. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/006641389190006W.

Straubel:1985:IOB

[SWR85] R. Straubel, A. Wittmüß, and R. Rosenmüller. An interactive optimization-based software system for multi-criteria (non concave) control problems. Annual Review in Automatic Programming, 12 (part 1):190–193, 1985. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413885900254.

Solntseff:1974:SEP

[SY74] N. Solntseff and A. Yezerski. A survey of extensible programming languages. Annual Review in Automatic Programming, 7 (part 5):267–307, 1974. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413874900019.

Sydow:1985:OOD

[Syd85a] Achim Sydow. Objective-oriented decomposition of complex decision problems. Annual Review in Automatic Programming, 12 (part 1):295–298, 1985. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413885900473.

Sydow:1985:P

[Syd85b] Achim Sydow. Preface. Annual Review in Automatic Programming, 12 (part 1):5–7, 1985. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413885900011.

Szlanko:1977:PNP

[Szl77] János Szlankó. Petri nets for proving some correctness properties of parallel programs. Annual Review in Automatic Programming, 8:75–83, 1977. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413877900118.

Szmuc:1988:CVR

[Szm88] T. Szmuc. Correctness verification of real-time programs. Annual Review in Automatic Programming, 14 (part 1):1–6, 1988.

CODEN ???? ISSN ???? URL https://www.sciencedirect.
com/science/article/pii/0066413888900018.

Tamura:1985:MDM

[Tam85] Hiroyuki Tamura. Multiobjective decision making — utility theoretic approach —. Annual Review in Automatic Programming, 12 (part 1):51–61, 1985. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413885900059.

Toro:1992:QMS

[TAR92] M. Toro, J. J. Arrabal, and L. Romero. Qualitative modelling and simulation by piecewise linear analysis. Annual Review in Automatic Programming, 17:147–152, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/S0066413809910254.

Taoka:1981:MSP

[TAT81] H. Taoka, S. Abe, and S. Takeda. Multiprocessor system for power system analysis. Annual Review in Automatic Programming, 11:101–106, 1981. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/006641388190015X.

Taylor:1960:FMM

[Tay60] Alan Taylor. The FLOW-MATIC and MATH-MATIC automatic programming systems. Annual Review in Automatic Programming, 1:196–206, 1960. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413860900422.

Taylor:1980:MEF

[Tay80] B. Taylor. A method for expressing the functional requirements of real-time systems. Annual Review in Automatic Programming, 10:111–120, 1980. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413882900155.

Thiele:1994:TCP

[TBR^{+}94] G. Thiele, H. J. Beestermöller, L. Renner, M. Dorno, and D. Popovic. Task-configuration of a pearl-based programmable controller for process-automation. Annual Review in Automatic Programming, 18:115–120, 1994. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413894900205.

Tempelmeier:1988:CSD

[Tem88] T. Tempelmeier. The creative step in designing real-time software. Annual Review in Automatic Programming, 14 (part 1): 15–20, 1988. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/0066413888900031.

Tempelmeier:1994:NHS

[Tem94] T. Tempelmeier. A note on hardware–software codesign. Annual Review in Automatic Programming, 18:121–126, ???? 1994. CODEN ARVAAM. ISSN 0066-4138 (print), 1878-545X

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0066413894900213.

Tsuruta:1981:DTR

[TFMM81] S. Tsuruta, K. Fukuoka, S. Miyamoto, and S. Mitsumori. A debugging tool for real time software: The non repeated runs approach. Annual Review in Automatic Programming, 11:49– 56, 1981. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/0066413881900082.

Thomesse:1977:NSS

[Tho77] J. P. Thomesse. A new set of software tools for designing and realizing distributed systems in process control. Annual Review in Automatic Programming, 8:47–53, 1977. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413877900088.

Tien:1985:SAA

[Tie85] Hoang Tien. Some application aspects of system modelling and simulation for large scale system control. Annual Review in Automatic Programming, 12 (part 1):299–302, 1985. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413885900485.

Troch:1985:MRS

[TKD85] I. Troch, P. Kopacek, and K. Desoyer. Models for robots — a simulation approach. Annual Review in Automatic Programming, 12 (part 1):307–310, 1985. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413885900503.

Tits:1985:ISR

[TM85] A. L. Tits and Z. Ma. Interaction, specification refinement, and tradeoff exploration in optimization-based design of engineering systems. Annual Review in Automatic Programming, 13 (part 2):189–194, 1985. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/006641388590480X.

Toshkov:1985:DCL

[Tos85] A. Toshkov. Distributed control linear systems design with pole assignment. Annual Review in Automatic Programming, 12 (part 1):429–432, 1985. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413885900783.

Tapolcai:1990:CRT

[TS90] L. Tapolcai and Z. Sopronfalvi. CRDB — a real time data base for distributed system with high reliability. Annual Review in Automatic Programming, 15 (part 1):71–73, 1990. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/006641389090030U.

Tol:1994:SSE

[TW94] Ronald M. Tol and Gert Te Winkel. Some scheduling experiments with ARTIE. Annual Review in Automatic Programming, 18:51–54, 1994. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/0066413894900108.

Tong:1989:ESS

[TZ89] Wei-Guang Tong and Lin-Liang Zhao. An expert system shell for analysis of real time signals. Annual Review in Automatic Programming, 15 (part 2):65–67, 1989. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413889900128.

Tzafestas:1985:APC

[Tza85] Spyros G. Tzafestas. Analysis and parameter computation of a delay coupled core nuclear reactor using Walsh series expansions. Annual Review in Automatic Programming, 12 (part 1):105–113, 1985. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413885900102.

Urbancic:1992:QR

[UJFB92] T. Urbančič, D. Juričić, B. Filipič, and I. Bratko. Qualitative reasoning. Annual Review in Automatic Programming, 17:341– 346, 1992. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/S0066413809910564.

Ulery:1985:SRS

[Ule85] D. L. Ulery. Software requirements for statistical quality control. Annual Review in Automatic Programming, 13 (part 1): 39–42, 1985. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/0066413885904410.

Unbehauen:1985:SIM

[Unb85] H. Unbehauen. System identification methods using parameter estimation/a survey. Annual Review in Automatic Programming, 12 (part 1):69–81, 1985. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413885900072.

Umeno:1981:LSC

[UOH^{+}81] H. Umeno, K. Ohmachi, A. Hino, J. Imura, and S. Obuki. Online system changeover using a virtual machine system. Annual Review in Automatic Programming, 11:29–34, 1981. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413881900057.

Ushiroda:1981:FDS

[Ush81] T. Ushiroda. File design of small real-time computer system. Annual Review in Automatic Programming, 11:85–92, 1981. CO-

DEN ???? ISSN ???? URL https://www.sciencedirect.com/
science/article/pii/0066413881900136.

Ulm:1985:GTM

[UW85] M. Ulm and H.-D Wend. A graph theoretic method for the characterization of decentrally stabilizable control systems. Annual Review in Automatic Programming, 12 (part 1):272–275, 1985.

CODEN ???? ISSN ???? URL https://www.sciencedirect.
com/science/article/pii/0066413885900412.

Verbruggen:1989:AIF

[VÅ89] H. B. Verbruggen and K. J. Åström. Artificial intelligence and feedback control. Annual Review in Automatic Programming, 15 (part 2):1–11, 1989. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413889900025.

Valyi:1985:ASV

[Vál85] István Vályi. Approximate solutions of vector optimization problems. Annual Review in Automatic Programming, 12 (part 1): 246–250, 1985. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/0066413885900357.

Vijayananda:1994:AMS

[VBRD94] K. Vijayananda, G. Berthet, P. Raja, and J. D. Decotignie. An architecture of a map stack for real time applications. Annual Review in Automatic Programming, 18:1–7, 1994. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413894900027.

Vina:1992:B

[VD92] A. Viña and J. M. Domínguez. Biotechnology. Annual Review in Automatic Programming, 17:423–430, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/S0066413809910692.

vanderLaan:1994:DAR

[vdL94] Marten D. van der Laan. Data acquisition for real-time process control systems. Annual Review in Automatic Programming, 18: 109–114, 1994. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/0066413894900191.

Vazquez:1994:LGM

[VG94] F. Vázquez and E. Garcia. A local guidance method for low-cost mobile robots using fuzzy logic. Annual Review in Automatic Programming, 19:203–207, 1994. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413894900663.

Vega:1994:DAI

[VGQ94] Juan J. Galán Vega and Agustín González-Quel. The development of an artificial intelligence real-time toolkit: REAKT. Annual Review in Automatic Programming, 19:333–339, 1994. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413894900876.

Vina:1992:KBR

[VHR92] A. Vina and B. Hayes-Roth. Knowledge-based real time control: The use of abstraction to satisfy deadlines. Annual Review in Automatic Programming, 16 (part 1):33–40, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/006641389190007X.

Vasilyev:1985:AMS

[VK85] V. V. Vasilyev and V. V. Kuzmuk. Application of multiprocessor systems for complex asynchronous process simulation. Annual Review in Automatic Programming, 12 (part 1):455–459, 1985.

CODEN ???? ISSN ???? URL https://www.sciencedirect.
com/science/article/pii/0066413885900850.

Vingerhoeds:1994:NPR

[VK94] R. A. Vingerhoeds and A. J. Krijgsman. The NECTAR-project research into the application of neural networks for flight control. Annual Review in Automatic Programming, 19:167–172, 1994.

CODEN ???? ISSN ???? URL https://www.sciencedirect.
com/science/article/pii/0066413894900604.

Vale:1992:FDE

[VM92] Z. A. Vale and A. Machado e Moura. Fault detection and emergency control in power systems. Annual Review in Automatic Programming, 17:409–414, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/S0066413809910679.

Voigt:1985:SFV

[Voi85] Hans-Michael Voigt. Structure formation, vector optimization and evolution. Annual Review in Automatic Programming, 12 (part 1):251–254, 1985. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413885900369.

Vojnovic:1977:KRT

[Voj77] D. Vojnovic. Kernel real-time system. Annual Review in Automatic Programming, 8:139–143, 1977. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413877900192.

Vorobiev:1985:AMS

[Vor85] O. Yu Vorobiev. Average measure simulation and cell analysis. Annual Review in Automatic Programming, 12 (part 1): 369–372, 1985. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/0066413885900643.

Vazquez:1994:RTV

[VPM94] F. Vázquez, E. Paz, and R. Marín. Real-time visionbased navigation and 3D depth estimation for an indoor autonomous mobile robot. Annual Review in Automatic Programming, 19:221–226, 1994. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413894900698.

Verbruggen:1992:PP

[VR92] H. B. Verbruggen and M. G. Rodd. Plenary papers. Annual Review in Automatic Programming, 17:vii–viii, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/S0066413809910023.

Velilla:1988:SMS

[VS88] S. Velilla and M. Silva. The spy: a mechanism for safe implementation of highly concurrent systems. Annual Review in Automatic Programming, 14 (part 1):75–81, 1988. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413888900122.

vanSchrick:1992:MRA

[vS92] D. van Schrick. Markovian reliability analysis of state-estimator-based instrument fault detection schemes. Annual Review in Automatic Programming, 17:253–258, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/S0066413809910424.

vanWijngaarden:1963:GA

[vW63] A. van Wijngaarden. Generalized ALGOL. Annual Review in Automatic Programming, 3:17–26, 1963. CODEN ARVAAM. ISSN 0066-4138 (print), 1878-545X (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0066413863800026.

Walter:1980:SLC

[Wal80] C. Walter. A structuring language for computer controlled multilevel systems. Annual Review in Automatic Programming, 10: 81–89, 1980. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/0066413882900118.

Weinreb:1985:MTC

[WB85] A. Weinreb and A. E. Bryson. Minimum-time control of a two-link robot arm. Annual Review in Automatic Programming, 13 (part 2):195–199, 1985. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413885904811.

Woods:1992:SEH

[WB92] E. A. Woods and J. G. Balchen. Structural estimation with the hybrid phenomena theory. Annual Review in Automatic Programming, 16 (part 1):127–132, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413891900213.

Wang:1992:FC

[WDM92] G. Wang, O. Dubant, and J. Magnier. Fuzzy control. Annual Review in Automatic Programming, 17:171–176, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/S0066413809910291.

Weisman:1985:LSP

[Wei85] R. A. Weisman. On-line statistical process control for a glass tank ingredient scale. Annual Review in Automatic Programming, 13 (part 1):29–38, 1985. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413885904409.

Weis:1988:EMF

[Wei88] E. Weis. Experience with modern, flexible methods in software project management. Annual Review in Automatic Programming, 14 (part 2):45–52, 1988. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/006641389090009G.

Weller:1985:CAD

[Wel85] W. Weller. Computer aided decision aids for manufacturing processes with flexible applications. Annual Review in Automatic Programming, 12 (part 1):198–201, 1985. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413885900278.

Wild:1992:FFL

[WFR92] D. Wild, J. Foldenauer, and U. Rembold. Flanex — a flow line analyzing expert system. Annual Review in Automatic Programming, 16 (part 1):145–148, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413891900246.

Willey:1977:IHL

[WHS77] M. S. Willey, J. F. Hoffmaster, and James D. Schoeffler. The integration of high level interpretive software with microprocessor based distributed control systems. Annual Review in Automatic Programming, 8:99–105, 1977. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413877900143.

Watanabe:1981:RTP

[WI81] T. Watanabe and S. Iwai. Real time programming of computer numerical control of a machine tool in CAM. Annual Review in Automatic Programming, 11:93–100, 1981. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413881900148.

Wilkes:1964:ESC

[Wil64] M. V. Wilkes. An experiment with a self-compiling compiler for a simple list-processing language. Annual Review in Automatic Programming, 4:1–48, January 1964. CO-DEN ARVAAM. ISSN 0066-4138 (print), 1878-545X (elec-

tronic).  URL https://www.sciencedirect.com/science/
article/pii/0066413864900138.

Wittmuss:1985:SMO

[Wit85] A. Wittmüß. Scalarizing multiobjective optimization problems and an interactive approach for multiobjective decision making. Annual Review in Automatic Programming, 12 (part 1): 255–258, 1985. CODEN ???? ISSN ???? URL https://www.

sciencedirect.com/science/article/pii/0066413885900370.

Witt:1992:DOO

[Wit92] Jan Witt. On drinking one’s own medicine — reusing software to build tools for software reuse. Annual Review in Automatic Programming, 167 (part 2):71–77, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/006641389290011D.

Wolczowski:1985:FMM

[Woł85] Andrzej Wołczowski. Formal model of movable object control with application of parametric automaton. Annual Review in Automatic Programming, 12 (part 1):398–403, 1985. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413885900710.

Woodger:1963:DCP

[Woo63] M. Woodger. The description of computing processes: Some observations on automatic programming and ALGOL 60. Annual Review in Automatic Programming, 3:1–15, January 1963. CODEN ARVAAM. ISSN 0066-4138 (print), 1878-545X (elec-

tronic).  URL https://www.sciencedirect.com/science/
article/pii/S0066413863800014.

Woods:1992:RCD

[Woo92] E. A. Woods. On representations for continuous dynamic systems. Annual Review in Automatic Programming, 17:347–352, 1992.

CODEN ???? ISSN ???? URL https://www.sciencedirect.
com/science/article/pii/S0066413809910576.

Woschni:1985:SPM

[Wos85] E.-G Woschni. Some problems of modelling of digital measuring systems. Annual Review in Automatic Programming, 12 (part 1):373–376, 1985. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/0066413885900655.

Wakamiya:1992:SRF

[WTI92] Y. Wakamiya, M. Tsuruda, and M. Isobe. Slab reheating furnace temperature control using AI. Annual Review in Automatic Programming, 16 (part 1):153–157, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413891900268.

Xinping:1985:OCX

[XL85] Yang Xinping and Fu Longzhou. Optimal control of xenon oscillations in load follow of a large nuclear reactor. Annual Review in Automatic Programming, 13 (part 2):201–207, 1985. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413885904823.

Yashkov:1985:DNJ

[Yas85] S. F. Yashkov. Distribution of number of jobs in a feedback processor-sharing queue. Annual Review in Automatic Programming, 12 (part 1):464–468, 1985. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413885900874.

Yu-Ji:1989:CMK

[YJ89] Huang Yu-Ji. A clustering method of knowledge acquisition in a real-time control system. Annual Review in Automatic Programming, 15 (part 2):49–51, 1989. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413889900098.

Yokoyama:1992:SOC

[YKI92] Y. Yokoyama, T. Kohda, and K. Inoue. Synthesis of optimal control using neural network with mixed structure. Annual Review in Automatic Programming, 17:311–316, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/S0066413809910515.

Yu-Lan:1989:ESL

[YLZ89] Zhou Yu-Lan and Qiu Zhengyue. An expert self-learning fuzzy controller. Annual Review in Automatic Programming, 15 (part 2):69–73, 1989. CODEN ???? ISSN

????  URL https://www.sciencedirect.com/science/
article/pii/006641388990013X.

Ypsilantis:1992:MLU

[YY92] J. Ypsilantis and H. Yee. Machine learning using version spaces for a power distribution network fault diagnostician. Annual Review in Automatic Programming, 17:115–120, 1992. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/S0066413809910205.

Zhukovin:1985:MDM

[ZBZ85] V. E. Zhukovin, F. V. Burshtein, and E. V. Zhukovin. Multicriterial decision making model with the empty Pareto set. Annual Review in Automatic Programming, 12 (part 1):182–184, 1985.

CODEN ???? ISSN ???? URL https://www.sciencedirect.
com/science/article/pii/0066413885900230.

Zhang:1989:AIC

[ZEZ89] Huiping Zhang, Yutian En, and Jinliang Zhang. Application of intelligent control of time-delay processes to the mix moisture control system in sinterplants. Annual Review in Automatic Programming, 15 (part 2):107–109, 1989. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413889900207.

Zhao:1992:PSN

[Zha92] Feng Zhao. Phase space navigator: Towards automating control synthesis in phase spaces for nonlinear control systems. Annual Review in Automatic Programming, 16 (part 1):97–102, 1992. CO-

DEN ???? ISSN ???? URL https://www.sciencedirect.com/
science/article/pii/0066413891900165.

Zhang:1994:FDC

[ZMM94] J. Zhang, A. J. Morris, and G. A. Montague. Fault diagnosis of a CSTR using fuzzy neural networks. Annual Review in Automatic Programming, 19:153–158, 1994. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/0066413894900582.

Zhang:1990:ILO

[ZZY90] Xia Zhang, Huaiyuan Zheng, and Ge Yu. Integrating logic and object paradigm in HDDBMS. Annual Review in Automatic Programming, 15 (part 1):57–61, 1990. CODEN ????

ISSN ???? URL https://www.sciencedirect.com/science/
article/pii/006641389090028P.